### v1

# On Poincaré–Sobolev level involving fractional GJMS operators on hyperbolic space

Abstract: This paper is devoted to a qualitative analysis of the Poincaré–Sobolev level associated with the fractional GJMS operators $\mathcal{P}_{s}$ $\bigl(s\in(0,\tfrac{n}{2})\setminus\mathbb{N}\bigr)$ on the hyperbolic space $\mathbb{H}^{n}$. In contrast to the integer-order case, when $s\notin\mathbb{N}$ the operator $\mathcal{P}_{s}$ does not enjoy the conformal covariance that allows one, in the upper half-space or ball model, to relate it to the Euclidean fractional Laplacian $(-\Delta)^{s}$; this link is crucial for importing Euclidean theory. We therefore introduce $\widetilde{\mathcal{P}}_{s}$ ($s>0$), which is conformally related to the $(-\Delta)^{s}$. Our purpose in the paper is to analyze the monotonicity, attainability, and strict-gap regions of the Poincaré–Sobolev levels associated with $\mathcal{P}_{s}$ and with $\widetilde{\mathcal{P}}_{s}$ with $s\in(0,\tfrac{n}{2})\setminus\mathbb{N}$.
First, we reinterpret the Brezis–Nirenberg problem through the lens of Poincaré–Sobolev levels, connecting earlier results for the Euclidean Laplacian and for operators $\mathcal{P}_{k}$ on $\mathbb{H}^{n}$ with integer $k\in(0,\tfrac{n}{2})$. We then establish new, explicit lower bounds for the Hardy term in fractional Hardy–Sobolev–Maz’ya inequalities involving both $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$. By applying the concentration–compactness principle together with a detailed analysis of the strict-gap regions for the Poincaré–Sobolev levels, we prove the existence of solutions to the Brezis–Nirenberg problem on $\mathbb{H}^{n}$ for both operators. Finally, combining the Hardy lower bounds with criteria for attainability, we obtain a complete characterization of the Poincaré–Sobolev levels $H_{n,s}$ and $\widetilde{H}_{n,s}$.
Keywords: Fractional GJMS Operators, Hyperbolic Space, Poincaré–Sobolev Constant, Brezis-Nirenberg Problem
1  Introduction and Main Results
The aim of this paper is to investigate the quantitative behavior of the Poincaré–Sobolev level on the hyperbolic space $\mathbb{H}^{n}$, defined by
$H_{n,s}(\lambda):=\inf_{u\in C_{c}^{\infty}(\mathbb{H}^{n})\setminus\{0\}}\frac{\displaystyle\int_{\mathbb{H}^{n}}(\mathcal{P}_{s}u)\,u\,dV_{\mathbb{H}^{n}}-\lambda\int_{\mathbb{H}^{n}}|u|^{2}\,dV_{\mathbb{H}^{n}}}{\Bigl(\displaystyle\int_{\mathbb{H}^{n}}|u|^{2_{s}^{*}}\,dV_{\mathbb{H}^{n}}\Bigr)^{\!2/2_{s}^{*}}},\quad\lambda\in\mathbb{R},$ (1.1)
where $n\geq 2$, $s\in\big(0,\frac{n}{2}\big)\setminus\mathbb{N},\ 2_{s}^{*}:=\frac{2n}{n-2s}$, and $\mathcal{P}_{s}$ denotes the $s$-order GJMS operator on $\mathbb{H}^{n}$. It admits the explicit spectral representation
$\mathcal{P}_{s}=2^{2s}\,\frac{\Bigl|\Gamma\!\Bigl(\frac{3+2s}{4}+\frac{{\bf i}}{2}\,\mathcal{A}\Bigr)\Bigr|^{2}}{\Bigl|\Gamma\!\Bigl(\frac{3-2s}{4}+\frac{{\bf i}}{2}\,\mathcal{A}\Bigr)\Bigr|^{2}}\qquad\text{with}\quad\mathcal{A}:=\sqrt{-\Delta_{\mathbb{H}^{n}}-\rho^{2}},\ \ \rho:=\frac{n-1}{2}.$ (1.2)
Here ${\bf i}=\sqrt{-1}$, $\Gamma$ denotes the Gamma function, and the functional calculus is taken on $L^{2}(\mathbb{H}^{n})$; see [24]. We emphasize that the right-hand side of (1.2) is well defined for all $s>0$, and we adopt (1.2) as the definition of $\mathcal{P}_{s}$ for $s\in(0,\infty)$.
For integer orders $s=k\in\mathbb{N}$, the GJMS operator $\mathcal{P}_{k}$ is conformally intertwined with the Euclidean fractional Laplacian in the standard models of $\mathbb{H}^{n}$. More precisely, in the upper half–space model $(\mathbb{H}^{n},g_{\mathbb{H}^{n}})$,
$x_{1}^{\,k+\frac{n}{2}}\,(-\Delta)^{k}\!\bigl(x_{1}^{\,k-\frac{n}{2}}u\bigr)\;=\;{\mathcal{P}}_{k}\,u\quad{\rm for}\ \,u\in C^{\infty}({\mathbb{R}^{n}_{+}}),$ (1.3)
while in the Poincaré ball model $(\mathbb{B}^{n},g_{\mathbb{B}^{n}})$,
$\Bigl(\frac{1-|x|^{2}}{2}\Bigr)^{\!k+\frac{n}{2}}(-\Delta)^{k}\!\Bigl[\Bigl(\frac{1-|x|^{2}}{2}\Bigr)^{\!k-\frac{n}{2}}u\Bigr]\;=\;{\mathcal{P}}_{k}\,u\quad{\rm for}\ \,u\in C^{\infty}({\mathbb{B}^{n}}),$ (1.4)
where $\Delta$ is the Laplacian in Euclidean space. In contrast, for non-integer orders $s\in(0,\frac{n}{2})\setminus\mathbb{N}$, the operator $\mathcal{P}_{s}$ is not conformally equivalent to the Euclidean fractional Laplacian on $\mathbb{R}^{n}_{+}$ or $\mathbb{B}^{n}$. To recover a usable intertwining structure, an auxiliary operator is involved
$\widetilde{\mathcal{P}}_{s}:=\frac{\bigl|\Gamma\!\bigl(s+\tfrac{1}{2}+i\mathcal{A}\bigr)\bigr|^{2}}{\bigl|\Gamma\!\bigl(\tfrac{1}{2}+i\mathcal{A}\bigr)\bigr|^{2}},$ (1.5)
which satisfies the conformal intertwining identities (1.3)–(1.4) for all $s>0$; see [24, Theorem 1.7]. This operator provides a convenient bridge between analysis on $\mathbb{H}^{n}$ and the Euclidean setting. Moreover, the precise relation between $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$ is given by [24, Corollary 5.3]:
$\mathcal{P}_{s}\;=\;\widetilde{\mathcal{P}}_{s}\;+\;\frac{\sin(\pi s)}{\pi}\,\Bigl|\Gamma\!\Bigl(s+\tfrac{1}{2}+i\mathcal{A}\Bigr)\Bigr|^{2},\quad s\in(0,\infty).$ (1.6)
In particular, for integer orders $s\in(0,\tfrac{n}{2})\cap\mathbb{N}$, one has $\mathcal{P}_{s}\;=\;\widetilde{\mathcal{P}}_{s}.$
We also introduce the hyperbolic Poincaré–Sobolev level associated with $\widetilde{\mathcal{P}}_{s}$. This auxiliary level will serve as a key tool in our analysis of $H_{n,s}(\lambda)$, allowing us to circumvent the fact that $\mathcal{P}_{s}$ does not admit a direct reduction to the Euclidean fractional Laplacian in the non-integer regime:
$\widetilde{H}_{n,s}(\lambda):=\inf_{u\in C_{c}^{\infty}(\mathbb{H}^{n})\setminus\{0\}}\frac{\displaystyle\int_{\mathbb{H}^{n}}(\widetilde{\mathcal{P}}_{s}u)\,u\,dV_{\mathbb{H}^{n}}-\lambda\int_{\mathbb{H}^{n}}|u|^{2}\,dV_{\mathbb{H}^{n}}}{\Bigl(\displaystyle\int_{\mathbb{H}^{n}}|u|^{2_{s}^{*}}\,dV_{\mathbb{H}^{n}}\Bigr)^{\!2/2_{s}^{*}}},\quad\lambda\in\mathbb{R}.$ (1.7)
In this paper, we will mainly analyze qualitative properties of the two level functions $\lambda\mapsto H_{n,s}(\lambda)$ and $\lambda\mapsto\widetilde{H}_{n,s}(\lambda)$, including monotonicity, attainability, as well as the associated threshold phenomena.
The motivation for introducing $H_{n,s}(\lambda)$ and $\widetilde{H}_{n,s}(\lambda)$ stems from its tight connection with Brezis–Nirenberg problem
$\mathcal{P}_{s}u=\lambda u+|u|^{p-1}u\quad\ \text{in }\,\mathbb{H}^{n}$ (1.8)
and
$\widetilde{\mathcal{P}}_{s}u=\lambda u+|u|^{p-1}u\quad\ \text{in }\,\mathbb{H}^{n},$ (1.9)
where
$1<p\leq 2_{s}^{*}-1,\qquad 2_{s}^{*}:=\frac{2n}{n-2s}.$
In fact, to establish the existence of nontrivial solution to (1.8), a key step is to determine for which values of $\lambda$ the following strict inequality holds:
$H_{n,s}(\lambda)\;<\;H_{n,s}(0),\quad\widetilde{H}_{n,s}(\lambda)\;<\;\widetilde{H}_{n,s}(0).$ (1.10)
Precisely, when inequality (1.10) holds, positive Poincaré–Sobolev levels $H_{n,s}(\lambda)$ and $\widetilde{H}_{n,s}(\lambda)$ attains their infimum, and the corresponding minimizer constitutes a nontrivial solution to the Brezis–Nirenberg problem (1.8) and (1.9). Consequently, a precise characterization of the strict-gap region, along with the attainability and monotonicity properties of Poincaré–Sobolev level—is not only analytically fundamental but also indispensable for establishing existence results for the underlying nonlinear equation. Our primary objective is therefore to characterize both the attainability of the infimum and the strict-gap region $\mathcal{G}_{n,s}[H_{n,s}]$ and $\mathcal{G}_{n,s}[\widetilde{H}_{n,s}]$associated with the function $H_{n,s}(\cdot)$ and $\widetilde{H}_{n,s}(\cdot)$, where
$\mathcal{G}_{n,s}[f]\;:=\;\big\{\lambda\in\mathbb{R}:\ f(\lambda)<f(0)\big\}.$ (1.11)
1.1 Euclidean (Fractional) Laplacian and Integer-Order GJMS Operators
To clarify how the strict-gap region (1.10) governs solvability of the Brezis–Nirenberg problem (1.8) and (1.9) on $\mathbb{H}^{n}$, let’s review the known results in a differential viewpoint on the settings: the classical and fractional Laplacians on bounded Euclidean domains and then local (integer-order) conformal Laplacian on hyperbolic space. We firstly recall
$S_{n,s}:=\inf_{v\in C_{c}^{\infty}(\mathbb{R}^{n})\setminus\{0\}}\frac{\displaystyle\int_{\mathbb{R}^{n}}v\,(-\Delta)^{s}v\,dx}{\Bigl(\displaystyle\int_{\mathbb{R}^{n}}|v|^{2_{s}^{*}}\,dx\Bigr)^{\!2/2_{s}^{*}}}$ (1.12)
with $s\in\big(0,\frac{n}{2}\big)$ and $(-\Delta)^{s}$ denoting the fractional laplacian on the Euclidean space $\mathbb{R}^{n}$ by
$\widehat{\,(-\Delta)^{s}v\,}(\xi)=|\xi|^{2s}\,\hat{v}(\xi),\quad\xi\in\mathbb{R}^{n},\ v\in C_{c}^{\infty}(\mathbb{R}^{n}).$
For $s\in(0,1]$, denote by $S_{n,s,\Omega}(\lambda)$ the Poincaré–Sobolev level associated with $(-\Delta)^{s}$ and $\lambda\in\mathbb{R}$:
$S_{n,s,\Omega}(\lambda):=\inf_{\begin{subarray}{c}v\in C_{c}^{\infty}(\Omega)\setminus\{0\}\end{subarray}}\frac{\displaystyle\int_{\Omega}v\,(-\Delta)^{s}v\,dx-\lambda\int_{\Omega}v^{2}\,dx}{\Bigl(\displaystyle\int_{\Omega}|v|^{2_{s}^{*}}\,dx\Bigr)^{2/2_{s}^{*}}},$ (1.13)
where $2_{s}^{*}=\frac{2n}{n-2s}$ and $\Omega\subset\mathbb{R}^{n}$ is either a bounded Lipschitz domain or $\Omega=\mathbb{R}^{n}$. When $\Omega=\mathbb{R}^{n}$, we simply write $S_{n,s}(\lambda):=S_{n,s,\mathbb{R}^{n}}(\lambda)$. In fact, for bounded Lipschitz domain $\Omega,$ the infimum in (1.13) is unchanged if one replaces $C_{c}^{\infty}(\Omega)$ by $H_{0}^{s}(\Omega)$, where
$H_{0}^{s}(\Omega):=\bigl\{u\in H^{s}(\mathbb{R}^{n}):\ u=0\ \text{in }\Omega^{c}\bigr\},$
since $C_{c}^{\infty}(\Omega)$ is dense in $H_{0}^{s}(\Omega)$ with respect to the $H^{s}$-norm (see [3]). Remark that
$S_{n,s,\Omega}(0)=S_{n,s}\quad{\rm and}\quad S_{n,s,\Omega}(\lambda)>0\ \ {\rm if}\ \ \lambda<\lambda_{1,s}(\Omega),$
where $S_{n,s}$ is given in (1.12) and $\lambda_{1,s}(\Omega)$ is the first eigenvalue of $(-\Delta)^{s}$ on $\Omega$ subject to boundary condition that $u=0$ in $\partial\Omega$ for $s=1$ or $u=0$ in $\mathbb{R}^{n}\setminus\Omega$ for $s\in(0,1)$. Obviously, the function $\lambda\in\mathbb{R}\mapsto S_{n,s,\Omega}(\lambda)$ is non-increasing. By (1.11), the strict gap set defined by
$\mathcal{G}_{n,s,\Omega}[S_{n,s,\Omega}]\;=\;\{\lambda\in\mathbb{R}:\ S_{n,s,\Omega}(\lambda)<S_{n,s}\}.$
When $s=1$, We start from the seminal paper of Haïm Brezis and Louis Nirenberg [4] in 1983, which concerns the existence of positive solutions to the critical semilinear Dirichlet problem
$\left\{\begin{aligned} -\Delta u&=\lambda u+|u|^{2^{*}-2}u&&\text{in }\Omega,\\[2.84526pt] u&=0&&\text{on }\partial\Omega,\end{aligned}\right$. (1.14)
where $\lambda\in\mathbb{R}$ is a real parameter, and $2^{*}=\frac{2n}{n-2}$ is the critical Sobolev exponent for the embedding $H_{0}^{1}(\Omega)\hookrightarrow L^{2^{*}}(\Omega)$. Later on, problem (1.14) is named as the Brezis–Nirenberg problem.
Note that if $\Omega$ is star-shaped, then (1.14) admits no positive solution for $\lambda\leq 0$ by Pohozaev’s identity and one rules out positive solutions when $\lambda\geq\lambda_{1,1}(\Omega)$ by testing (1.14) against the first eigenfunction. Brezis and Nirenberg built the crucial inequalities $0<S_{n,1,\Omega}(\lambda)<S_{n,1}$ to guarantee the existence for $\lambda\in\big(0,\lambda_{1,1}(\Omega)\big)$ when $n\geq 4$ and for $\lambda\in\big(\tfrac{1}{4}\lambda_{1,1}(B_{1}),\lambda_{1,1}(B_{1})\big)$ when $N=3$. In these cases, $S_{n,1,\Omega}(\lambda)$ is achieved in $H_{0}^{1}(\Omega)$ and by a Lagrange multiplier argument, the minimizer solves (1.14). Moreover, for $n=3$ and $\lambda\leq\tfrac{1}{4}\lambda_{1,1}(B_{1})$, they also showed that problem (1.14) admits no positive solution.
When $s\in(0,1)$, if the infimum in (1.13) is achieved by some nontrivial $v\in H_{0}^{s}(\Omega)$ and $S_{n,s,\Omega}(\lambda)>0$, then a suitable scaling of $v$ yields a positive solution to the fractional Brezis–Nirenberg problem.
$\left\{\begin{aligned} (-\Delta)^{s}u&=\lambda u+|u|^{2_{s}^{*}-2}u&&\text{in }\Omega,\\ u&=0&&\text{in }\mathbb{R}^{n}\setminus\Omega.\end{aligned}\right$. (1.15)
A remarkable feature of the nonlocal regime is that the existence theory depends on the interplay between the dimension $n$ and the order $s$. In the range $n>2s$, one has the following picture:
  • High dimensions relative to $s\in(0,1)$: $n\geq 4s$. One has ([3, Claim 14.1])
    $(0,\infty)\subset\mathcal{G}_{n,s,\Omega}[S_{n,s,\Omega}]$
    and consequently (1.15) admits a nontrivial weak solution for $\lambda\in(0,\lambda_{1,s}(\Omega)).$ ([3, Theorem 14.1])
  • Low dimensions relative to $s\in(0,1)$: $2s<n<4s$. There exists a constant $\lambda_{s}^{*}>0$ such that ([3, Proposition 16.4])
    $(\lambda_{s}^{*},\infty)\subset\mathcal{G}_{n,s,\Omega}[S_{n,s,\Omega}].$
    In this regime, the relation between $\lambda_{s}^{*}$ and $\lambda_{1,s}(\Omega)$ is unknown—even when $\Omega$ is the unit ball; consequently, the standard mountain–pass scheme alone does not guarantee existence. In [3, Theorem 16.1], the authors combine mountain–pass and linking arguments, treating separately the cases $\lambda_{s}^{*}<\lambda_{1,s}(\Omega)$ and $\lambda_{s}^{*}\geq\lambda_{1,s}(\Omega)$, and thereby obtain nontrivial solutions of (1.15) for every $\lambda>\lambda_{s}^{*}$ that is not a Dirichlet eigenvalue of $(-\Delta)^{s}$.
Building upon the aforementioned findings, we derive the following conclusion.
Proposition 1.1.
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded Lipschitz domain, $s\in(0,1]$ and $n>2s$, and set $2_{s}^{*}=\frac{2n}{n-2s}$. Then the following statements hold:
  1. (i)
    For $n\geq 4s$, $\mathcal{G}_{n,s,\Omega}[S_{n,s,\Omega}]=(0,+\infty)$ and $S_{n,s,,\Omega}(\lambda)$ is achieved in $H_{0}^{s}\left(\Omega\right)$ if and only if $\lambda\in\mathcal{G}_{n,s,\Omega}[S_{n,s}]$. Moreover, $S_{n,s,\Omega}(\cdot)$ is strictly decreasing in $\mathcal{G}_{n,s,\Omega}[S_{n,s}]$,
    $\mu|\Omega|^{\frac{2s}{n}}\leq S_{n,s,\Omega}(\mu)<S_{n,s,\Omega}(\lambda_{1,s}\left(\Omega\right))=0\ \,{\rm for}\ \mu>\lambda_{1,s}(\Omega)\quad{\rm and}\quad S_{n,s,\Omega}(\lambda)=S_{n,s}\ \,{\rm for}\ \lambda\leq 0.$
  2. (ii)
    For $s\in(0,1)$ and $2s<n<4s$, $(\lambda_{s}^{*},+\infty)\subset\mathcal{G}_{n,s,\Omega}[S_{n,s,\Omega}]$ and $S_{n,s,\Omega}(\lambda)$ is achieved in $H_{0}^{s}\left(\Omega\right)$ if $\lambda\in(\lambda_{s}^{*},+\infty)$. Moreover, $S_{n,s,\Omega}(\cdot)$ is strictly decreasing in $(\lambda_{s}^{*},+\infty)$ and for any $\lambda\leq 0$,
    $S_{n,s,\Omega}(\lambda)=S_{n,s}.$
  3. (iii)
    For $\Omega=B_{1}$, $\mathcal{G}_{3,1,B_{1}}[S_{3,1,B_{1}}]=(\tfrac{1}{4}\lambda_{1,1}(B_{1}),+\infty)$, $S_{3,1,B_{1}}(\lambda)$ is achieved in $H_{0}^{s}\left(\Omega\right)$ if and only if $\lambda\in\mathcal{G}_{3,1,B_{1}}[S_{3,1,B_{1}}]$. Moreover, $S_{3,1,B_{1}}(\cdot)$ is strictly decreasing in $\mathcal{G}_{3,1,B_{1}}[S_{3,1,B_{1}}]$, and for $\lambda\leq\tfrac{1}{4}\lambda_{1,1}(B_{1})<\lambda_{1,1}(B_{1})<\mu$
    $\mu|B_{1}|^{\frac{2}{3}}\leq S_{3,1,B_{1}}(\mu)<S_{3,1,B_{1}}(\lambda_{1,1}\left(B_{1}\right))=0\quad{\rm and}\quad S_{3,1,B_{1}}(\lambda)=S_{3,1}.$
Remark 1.1.
Let $\mathcal{G}_{n,s,\Omega}[S_{n,s,\Omega}]$ denote the strict–gap region, i.e., the set of parameters for which $S_{n,s,\Omega}(\lambda)<S_{n,s}$. When $n\geq 4s$, this region is completely understood: one has the full classification
$\mathcal{G}_{n,s,\Omega}[S_{n,s,\Omega}]\;=\;(0,+\infty).$
In contrast, in the intermediate regime $2s<n<4s$, a sharp description of $\mathcal{G}_{n,s,\Omega}[S_{n,s,\Omega}]$ remains open. In the local case $s=1$, the situation is settled for the unit ball $\Omega=B_{1}$. However, for $s\in(0,1)$, it is still open to identify a critical threshold $\Lambda\leq\lambda_{s}^{*}$ such that
[…]
even when $\Omega=B_{1}$. Furthermore, for $s\in(1,\tfrac{n}{2})$, the strict–gap region poses significant analytical challenges and the Brezis–Nirenberg problem remains comparatively underexplored.
On the hyperbolic space $\mathbb{H}^{n}$, let’s introduce the conformal GJMS operators (see [12, 7, 10, 18])
$\mathcal{P}_{1}=-\Delta_{\mathbb{H}^{n}}-\tfrac{n(n-2)}{4}=\mathcal{A}^{2}+\tfrac{1}{4}$
and for integer order $k\in\big[2,\frac{n}{2}\big)$
$\mathcal{P}_{k}\;=\;\mathcal{P}_{1}\,(\mathcal{P}_{1}+2)\cdots\bigl(\mathcal{P}_{1}+k(k-1)\bigr)\;=\;\prod_{j=1}^{k}\Bigl(\mathcal{A}^{2}+\bigl(j-\tfrac{1}{2}\bigr)^{2}\Bigr),$
where $\Delta_{\mathbb{H}^{n}}$ is the Laplace–Beltrami and $\mathcal{A},\rho$ are given in (1.2). Let $2_{k}^{*}:=\frac{2n}{n-2k}$, which is the critical exponent for order $2k$. In particular, the bottom of $\mathcal{P}_{k}$ satisfies
$\lambda_{0,k}^{\mathrm{conf}}:=\inf\sigma(\mathcal{P}_{k})\;=\;\prod_{j=1}^{k}\Bigl(j-\tfrac{1}{2}\Bigr)^{\!2}.$
Recently, Brezis–Nirenberg type problems on the hyperbolic space have been studied in two principal settings: (i) on bounded domains, and (ii) on the whole space.
In what follows, we focus on the whole space problem; for results on bounded domains we refer to [29, 2, 20]. Specifically, we consider
$\Delta_{\mathbb{H}^{n}}u=\lambda u+|u|^{2^{*}-2}u\qquad\text{in }\mathbb{H}^{n},$ (1.16)
where $2^{*}=\frac{2n}{n-2}$, and the bottom of the $L^{2}$–spectrum of $\Delta_{\mathbb{H}^{n}}$ equals $\frac{(n-1)^{2}}{4}$. In [25], authors established the following existence result for positive solutions:
  • If $n\geq 4$, then for $\frac{n(n-2)}{4}<\lambda\leq\frac{(n-1)^{2}}{4},$ problem (1.16) admits a positive entire solution ([25, Theorem 1.5]); for $\lambda\leq\frac{n(n-1)}{4}$, problem (1.16) does not have any positive entire solution.
  • If $n=3$, then for $\lambda\;\leq\;1$, problem (1.16) has no positive entire solution.
The appearance of the quantity $\frac{n(n-2)}{4}$ in (1.16) is in fact natural from the conformal viewpoint. In fact, (1.16) rewrites as
$\mathcal{P}_{1}u=\mu u+|u|^{2^{*}-2}u\quad\text{in }\mathbb{H}^{n},\qquad\mu=\lambda-\frac{n(n-2)}{4}.$ (1.17)
Since the bottom of the $L^{2}$–spectrum of $\Delta_{\mathbb{H}^{n}}$ is $\frac{(n-1)^{2}}{4}$, the bottom of the spectrum of $\mathcal{P}_{1}$ is $\frac{1}{4}$, so the admissible window for $\lambda$ is exactly a shift of size $\frac{n(n-2)}{4}$ from the spectral bottom. With this normalization, the results of [25] can be restated as follows:
  • If $n\geq 4$, then $0<\mu\leq\frac{1}{4}$ guarantees a positive entire solution to (1.17) [25, Theorem 1.5]; whereas for $\mu\;\leq\;0,$ problem (1.17) admits no positive entire solution [25, Theorem 1.6].
  • If $n=3$, then for $\mu\;\leq\;\frac{1}{4},$ problem (1.17) has no positive entire solution [25, Theorem 1.7].
When the integer $k\geq 2$, the higher integer order Brezis–Nirenberg problems on hyperbolic spaces have been studied in [20, 23]. Based upon these results, we derive the following conclusion.
Proposition 1.2.
Assume that integers $k\geq 1,\,n>2k$ and $H_{n,k}(\lambda),S_{n,k},\mathcal{G}_{n,k}$ be given in (1.1),(1.12) and (1.11) respectively.
  1. (i)
    When $n\geq 4k$, $\mathcal{G}_{n,k}[H_{n,k}]=(0,+\infty)$. $H_{n,k}(\lambda)$ is achieved and strictly decreasing for every $\lambda\in(0,\lambda_{0,k}^{\mathrm{conf}})$ when $k\geq 2$, while for $k=1,$ it is achieved if and only if $\lambda\in(0,\lambda_{0,1}^{\mathrm{conf}}]$ and strictly decreasing on $(0,\lambda_{0,k}^{\mathrm{conf}}]$, -∞=H_n,k(μ)¡0¡H_n,k(λ_0,k^conf)   for μ¿λ_0,k^conf and H_n,k(λ)=S_n,k for $\lambda\leq 0$.
  2. (ii)
    When $2k+2\leq n\leq 4k-1$, there exists $\lambda_{k}^{\mathrm{conf}}\in(0,\lambda_{0,k}^{\mathrm{conf}})$ such that $(\lambda_{k}^{\mathrm{conf}},\infty)\subset\mathcal{G}_{n,k}[H_{n,k}]$. Moreover, $H_{n,k}(\lambda)$ is achieved and strictly decreasing for $\lambda\in(\lambda_{k}^{\mathrm{conf}},\lambda_{0,k}^{\mathrm{conf}})$, -∞=H_n,k(μ)¡0¡H_n,k(λ_0,k^conf)   for μ¿λ_0,k^conf and H_n,k(λ)=S_n,k for $\lambda\leq 0$.
  3. (iii)
    When $n=2k+1$, $\mathcal{G}_{n,k}[H_{n,k}]=(\lambda_{0,k}^{\mathrm{conf}},+\infty)$. Moreover,
    $\infty=H_{n,k}(\mu)<0<S_{n,k}=H_{n,k}(\lambda),\quad\mu>\lambda_{0,k}^{\mathrm{conf}}\geq\lambda.$
    In particular, $H_{n,k}(\lambda)$ is never achieved for any $\lambda\in\mathbb{R}$ when $k=1$.
Remark 1.2.
(a) There is a jump discontinuity of $H_{n,k}(\cdot)$ at $\lambda=\lambda_{0,k}^{\mathrm{conf}}$. This phenomenon stems from the fact that $\tfrac{1}{4}$ is not an eigenvalue, but rather the bottom of the continuous spectrum of $\mathcal{P}_{k}$. This is in sharp contrast with the Euclidean quantity $S_{n,s,\Omega}(\lambda)$ at $\lambda=\lambda_{1,s}(\Omega)$, since $\lambda_{1,s}(\Omega)$ is the first Dirichlet eigenvalue of $(-\Delta)^{s}$ on $\Omega$.
$(b)$ Note the strict–gap region $\mathcal{G}_{n,k}[H_{n,k}]$ can be given explicitly for $n\geq 4k$ and $n=2k+1$. In contrast, for $2k+2\leq n\leq 4k-1$, it remains open to give $\mathcal{G}_{n,k}[H_{n,k}]$ with an explicit interval.
(c) We conjecture that, in the regime $n\geq 4k$ with $k\geq 2$, the level $H_{n,k}(\lambda)$ is achieved if and only if $\lambda\in(0,\lambda_{0,k}^{\mathrm{conf}}]$, and when $n=2k+1$, $H_{n,k}(\lambda)$ is never achieved for any $\lambda\in\mathbb{R}$ as soon as $k\geq 2$.
The proofs of Propositions 1.1 and 1.2 mainly rely on the classical analysis of Poincaré–Sobolev levels, together with several auxiliary lemmas, and on combining these tools with previously known results in the literature. For the convenience of the reader, we provide detailed proofs in the Appendix 5.
1.2 Fractional GJMS Operators
Fractional conformally covariant operators on the conformal infinity of a Poincaré–Einstein manifold were introduced by Graham and Zworski through scattering theory in their seminal work in [13], building on the foundational analytic framework of Mazzeo and Melrose on meromorphic continuation of the resolvent in the [26]. More precisely, let $(X^{n+1},g_{+})$ be a conformally compact Einstein manifold with conformal infinity $(M^{n},[g])$, and fix $\gamma\in(0,\tfrac{n}{2})\setminus\mathbb{N}$ with $s=\tfrac{n}{2}+\gamma$. Given boundary data $f$ on $M$, consider the generalized eigenvalue problem
$\Delta_{g_{+}}u-s(n-s)u=0,$
whose solutions admit an expansion $u=r^{n-s}(f+\cdots)+r^{s}(h+\cdots)$ near $M$. The scattering operator $S(s)$ is defined by $S(s)f=h$, and the fractional GJMS operator $\mathcal{P}_{\gamma}$ is obtained (up to a normalization) from $S(\tfrac{n}{2}+\gamma)$. The resulting $\mathcal{P}_{\gamma}$ is an elliptic nonlocal pseudodifferential operator of order $2\gamma$ and obeys the expected conformal covariance law. We refer to Chang–González in [5] for the extension-type characterization and further analytic developments, and to related works on fractional Yamabe-type problems (e.g., in [9]), as well as to the broader scattering literature on asymptotically hyperbolic geometry (e.g., in [17]).
We now return to the general fractional conformal operators $\mathcal{P}_{s}$ with $s\in(0,\tfrac{n}{2})\setminus\mathbb{N}$ and $\widetilde{\mathcal{P}}_{s}$ with $s\in(0,\tfrac{n}{2})$. Recalling the spectral representation (1.2) and (1.5), we obatin (see section 2.3)
$\lambda_{0,s}^{\mathrm{conf}}\;:=\;\inf\sigma(\mathcal{P}_{s})\;=\;2^{2s}\,\frac{\Gamma\!\bigl(\tfrac{3+2s}{4}\bigr)^{2}}{\Gamma\!\bigl(\tfrac{3-2s}{4}\bigr)^{2}}\geq 0$ (1.18)
and
$\widetilde{\lambda}_{0,s}^{\mathrm{conf}}=\inf\sigma(\widetilde{\mathcal{P}}_{s})=\frac{\Gamma\!\bigl(s+\frac{1}{2}\bigr)^{2}}{\Gamma\!\bigl(\frac{1}{2}\bigr)^{2}}>0.$ (1.19)
We first establish explicit lower bounds for the Hardy-term coefficient in the fractional Hardy–Sobolev–Maz’ya inequalities involving $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$. These bounds, in turn, yield partial characterizations of the Poincaré–Sobolev levels $H_{n,s}$ and $\widetilde{H}_{n,s}$.
Theorem 1.3.
Assume that $n\geq 2,s\in(0,\tfrac{n}{2})$, $S_{n,s},\,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$ are defined in (1.12)(1.19) respectively. Consider the optimal lower shift for the inequality
$\widetilde{\Lambda}_{n,s}^{\mathrm{HS}}\;:=\;\inf\Bigl\{\lambda\in\mathbb{R}:\ \int_{\mathbb{H}^{n}}(\widetilde{\mathcal{P}}_{s}u)\,u\,dV_{\mathbb{H}^{n}}+\lambda\int_{\mathbb{H}^{n}}|u|^{2}\,dV_{\mathbb{H}^{n}}\;\geq\;S_{n,s}\Bigl(\int_{\mathbb{H}^{n}}|u|^{2_{s}^{*}}\,dV_{\mathbb{H}^{n}}\Bigr)^{\!\frac{2}{2_{s}^{*}}}\ \ \forall\,u\in C_{c}^{\infty}(\mathbb{H}^{n})\Bigr\}.$
Then:
  1. (i)
    If $s\in(0,\tfrac{n}{4}]$, one has $\widetilde{\Lambda}_{n,s}^{\mathrm{HS}}\geq 0$.
  2. (ii)
    If $s\in(\tfrac{n}{4},\tfrac{n}{2})$, then there exists $\widetilde{\lambda}_{s}^{\mathrm{conf}}\in(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}]$ such that $\widetilde{\Lambda}_{n,s}^{\mathrm{HS}}\;\geq\;-\widetilde{\lambda}_{s}^{\mathrm{conf}}.$
For $s\in(0,\,n/4]\cap\mathbb{N}$, the above result recovers [23, Theorem 1.7] and yields an alternative proof thereof; whereas for $s\in(n/4,\,n/2)$, they established that $\lambda_{s}^{\mathrm{conf}}\in\bigl(0,\,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}\bigr).$
Next, from the relation identity (1.6) between $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$ we introduce an important parameter
$b_{s}:=\max\Bigl\{0,\ \frac{\sin(\pi s)}{\pi}\Bigr\}\,\bigl|\Gamma\!\bigl(s+\tfrac{1}{2}\bigr)\bigr|^{2}\quad{\rm for}\ \,s>0.$ (1.20)
And we have following result:
Theorem 1.4.
Let $n\geq 2$, $s\in\left(0,\frac{n}{2}\right)\setminus\mathbb{N}$, $b_{s},\lambda_{0,s}^{\mathrm{conf}}$ be defined in (1.20), (1.18) respectively and the optimal lower shift for the inequality
$\Lambda_{n,s}^{\mathrm{HS}}\;:=\;\inf\Bigl\{\lambda\in\mathbb{R}:\ \int_{\mathbb{H}^{n}}(\mathcal{P}_{s}u)\,u\,dV_{\mathbb{H}^{n}}+\lambda\int_{\mathbb{H}^{n}}|u|^{2}\,dV_{\mathbb{H}^{n}}\;\geq\;S_{n,s}\Bigl(\int_{\mathbb{H}^{n}}|u|^{2_{s}^{*}}\,dV_{\mathbb{H}^{n}}\Bigr)^{\!\frac{2}{2_{s}^{*}}}\ \ \forall\,u\in C_{c}^{\infty}(\mathbb{H}^{n})\Bigr\}.$
$(a)$ One has that
$\lambda_{0,s}^{\mathrm{conf}}-b_{s}=\begin{cases}\displaystyle\frac{\Gamma\bigl(s+\tfrac{1}{2}\bigr)^{2}}{\pi},&\sin(\pi s)>0,\\[5.16663pt] \displaystyle\frac{1+\sin(\pi s)}{\pi}\,\Gamma\bigl(s+\tfrac{1}{2}\bigr)^{2},&\sin(\pi s)\leq 0.\end{cases}$ (1.21)
$(b)$ One has that
$\Lambda_{n,s}^{\mathrm{HS}}\geq-b_{s}\quad\text{for $s\in(0,\tfrac{n}{4}]$}$
and
$\Lambda_{n,s}^{\mathrm{HS}}\;\geq\;-\min\left\{b_{s}+\widetilde{\lambda}_{s}^{\mathrm{conf}},\lambda_{0,s}^{\mathrm{conf}}\right\}\quad\text{for $s\in(\tfrac{n}{4},\tfrac{n}{2})$. }$
We next provide a characterization of the Poincaré–Sobolev level $\widetilde{H}_{n,s}$. The proof combines Theorem 1.3 with the attainability analysis for $\widetilde{H}_{n,s}$.
Theorem 1.5.
Let $n\geq 2$ and $s\in(0,\frac{n}{2})$. Then:
  1. (i)
    For $s\in\left(0,\frac{n}{4}\right]$, $\mathcal{G}_{n,s}[\widetilde{H}_{n,s}]=(0,+\infty)$. $\widetilde{H}_{n,s}(\lambda)$ is achieved and strictly decreasing for every $\lambda\in(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}})$, -∞=~H_n,s(μ)¡0¡~H_n,s(~λ_0,s^conf)   for μ¿~λ_0,s^conf and ~H_n,s(λ)=S_n,s for $\lambda\leq 0$.
  2. (ii)
    For $s\in(\frac{n}{4},\frac{n-1}{2})$ when $n\geq 3$ and $s\in(\frac{1}{2},1)$ when $n=2$, then there exists $\widetilde{\lambda}_{s}^{\mathrm{conf}}\in(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}]$ such that $\bigl(\widetilde{\lambda}_{s}^{\mathrm{conf}},\infty\bigr)\;\subset\;\mathcal{G}_{n,s}\bigl[\widetilde{H}_{n,s}\bigr]$ and for any $\lambda\leq 0<\widetilde{\lambda}_{0,s}^{\mathrm{conf}}<\mu$ one has
    $\infty=\widetilde{H}_{n,s}(\mu)\;<\;\widetilde{H}_{n,s}(\lambda)\;=\;S_{n,s}.$
  3. (iii)
    When $n\geq 3$ and $s\in[\frac{n-1}{2},\frac{n}{2}),$ $\mathcal{G}_{n,s}\bigl[\widetilde{H}_{n,s}\bigr]\;=\;\bigl(\widetilde{\lambda}_{0,s}^{\mathrm{conf}},\infty\bigr)$. Moreover,
[…]
Remark 1.3.
We conjecture that, in the regime $s\in(0,\frac{n}{4}]$, the level $\widetilde{H}_{n,s}(\lambda)$ is achieved if and only if $\lambda\in(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}]$, and when $n\geq 3$, $s\in[\frac{n-1}{2},\frac{n}{2}),$ $\widetilde{H}_{n,s}(\lambda)$ is never achieved for any $\lambda\in\mathbb{R}$.
We introduce the following notation
$\mathcal{B}_{+}:=\Bigl\{s\in(0,\infty)\,:\,\sin(\pi s)>0\Bigr\},\qquad\mathcal{B}_{0}:=\Bigl\{s\in(0,\infty)\,:\,\sin(\pi s)\leq 0\Bigr\}.$
Accordingly, we have $\lambda_{0,s}^{\mathrm{conf}}>b_{s}>0$ for $s\in\mathcal{B}_{+}$, while $\lambda_{0,s}^{\mathrm{conf}}\geq b_{s}=0$ for $s\in\mathcal{B}_{0}$.
Theorem 1.6.
Let $n\geq 2$ and $s\in(0,\tfrac{n}{2})\setminus\mathbb{N}$.
  1. (i)
    For $s\in\left(0,\frac{n}{4}\right]\cap\mathcal{B}_{0}$, one has that H_n,s(λ)¡S_n,s   for λ¿0 and H_n,s(λ)≤S_n,s for $\lambda\leq 0$.
  2. (ii)
    For $s\in\left(0,\frac{n}{4}\right]\cap\mathcal{B}_{+}$, $(b_{s},\infty)\subset\mathcal{G}_{n,s}\bigl[H_{n,s}\bigr]$ and $H_{n,s}(\lambda)$ is achieved and strictly decreasing for every $\lambda\in(b_{s},\lambda_{0,s}^{\mathrm{conf}})$. Moreover, H_n,s(λ)=S_n,s  for $\lambda\leq 0$.
  3. (iii)
    For $s\in(\frac{n}{4},\frac{n-1}{2})$ when $n\geq 3$ and $s\in(\frac{1}{2},1)$ when $n=2$, then there exists
    $\lambda_{s}^{\mathrm{conf}}:=\min\left\{{b_{s}+\widetilde{\lambda}_{s}^{\mathrm{conf}},\lambda_{0,s}^{\mathrm{conf}}}\right\}\in[0,\lambda_{0,s}^{\mathrm{conf}}]$
    such that for every $\lambda\leq 0<\lambda_{s}^{\mathrm{conf}}<\mu$,
    $H_{n,s}(\mu)\;<\;S_{n,s},\quad H_{n,s}(\lambda)\leq S_{n,s}.$
  4. (iv)
    When $n\geq 3$ and $s\in[\frac{n-1}{2},\frac{n}{2})$, $\mathcal{G}_{n,s}\bigl[H_{n,s}\bigr]=\bigl({\lambda}_{0,s}^{\mathrm{conf}},\infty\bigr).$ Moreover,
[…]
Below we systematically collect several open questions concerning the operators considered above and, in particular, the still largely unresolved issue of attainability for the associated Poincaré–Sobolev levels.
Unsolved problems:
(i) Let $k\geq 2$ and $n=2k+1$, we conjecture that problem (1.8) admits no positive solution for any $\lambda\in\mathbb{R}$, thus $H_{n,k}$ is never achieved; when $n\geq 4k$ with $k\geq 2$, the level $H_{n,k}(\lambda)$ is achieved if and only if $\lambda\in(0,\lambda_{0,k}^{\mathrm{conf}}]$
(ii) Let $k\geq 2$ and $2k+2\leq n\leq 4k-1$. Find an explicit optimal threshold $\Lambda_{k}^{\mathrm{conf}}$ satisfying
$0\leq\Lambda_{k}^{\mathrm{conf}}\leq\lambda_{k}^{\mathrm{conf}}<\lambda_{0,k}^{\mathrm{conf}}$
such that the strict–gap region for the hyperbolic level $H_{n,k}$ satisfies
$\mathcal{G}_{n,k}[H_{n,k}]=\bigl(\Lambda_{k}^{\mathrm{conf}},\,\infty\bigr).$
Equivalently, find the sharp value $\Lambda_{k}^{\mathrm{conf}}$ for which $H_{n,k}(\lambda)<S_{n,k}$ holds if and only if $\lambda>\Lambda_{k}^{\mathrm{conf}}$.
(iii) When $n\geq 3$ and $s\in[\frac{n-1}{2},\frac{n}{2}),$ we conjecture that problem (1.9) admits no positive solution for any $\lambda\in\mathbb{R}$, thus $\widetilde{H}_{n,s}$ is never achieved; for $s\in(0,\frac{n}{4}]$, we conjecture that $\widetilde{H}_{n,s}(\lambda)$ is achieved if and only if $\lambda\in(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}]$.
(iv) Assume $s\in(\frac{n}{4},\frac{n-1}{2})$ when $n\geq 3$ and $s\in(\frac{1}{2},1)$ when $n=2$ with $s\notin\mathbb{N}$. Find an explicit optimal threshold $\Lambda_{s}^{\mathrm{conf}}$ satisfying
$0\leq\Lambda_{s}^{\mathrm{conf}}\leq\widetilde{\lambda}_{s}^{\mathrm{conf}}\leq\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$
such that the strict–gap region for the fractional hyperbolic level $\widetilde{H}_{n,s}$ satisfies
$\mathcal{G}_{n,s}\!\bigl[\widetilde{H}_{n,s}\bigr]=\bigl(\Lambda_{s}^{\mathrm{conf}},\,\infty\bigr).$
Equivalently, find the sharp value $\Lambda_{s}^{\mathrm{conf}}$ for which $\widetilde{H}_{n,s}(\lambda)<S_{n,s}$ holds if and only if $\lambda>\Lambda_{s}^{\mathrm{conf}}$.
(v) For $\mathcal{P}_{s}$, we can consider analogous problems to (iii) (iv).
Finally, we address the Brezis–Nirenberg problem driven by $\mathcal{P}_{s}$ and by $\widetilde{\mathcal{P}}_{s}$, and describe the existence of nontrivial solutions in terms of the parameter $\lambda$. In fact, these existence results follow as a consequence of the attainability of the corresponding Poincaré–Sobolev levels $H_{n,s}$ and $\widetilde{H}_{n,s}$.
Theorem 1.7.
Let $s\in(0,\frac{n}{2})$, then problem (1.9) admits a nontrivial solution under assumptions:
  • (i)
    (Subcritical case) If $1<p<2_{s}^{*}-1$, assume that $\lambda<\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$.
  • (ii)
    (Critical case) If $p=2_{s}^{*}-1$, assume that $s$ and $\lambda$ satisfy
    $\mathcal{G}_{n,s}\!\big[\widetilde{H}_{n,s}\big]\cap(-\infty,\widetilde{\lambda}_{0,s}^{\mathrm{conf}})\neq\varnothing\qquad\text{and}\qquad\lambda\in\mathcal{G}_{n,s}\!\big[\widetilde{H}_{n,s}\big]\cap(-\infty,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}).$
    In particular, when $s\in(0,\frac{n}{4}]$, one has
    $\mathcal{G}_{n,s}\!\big[\widetilde{H}_{n,s}\big]\cap(-\infty,\widetilde{\lambda}_{0,s}^{\mathrm{conf}})=(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}).$
Theorem 1.8.
Let $s\in(0,\frac{n}{2})\setminus\mathbb{N}$. Problem (1.8) admits a nontrivial solution under assumptions:
  • (i)
    (Subcritical case) If $1<p<2_{s}^{*}-1$, assume that $\lambda<\lambda_{0,s}^{\mathrm{conf}}$.
  • (ii)
    (Critical case) If $p=2_{s}^{*}-1$, assume that $s$ and $\lambda$ satisfy
    $\mathcal{G}_{n,s}\!\big[H_{n,s}\big]\cap(-\infty,\lambda_{0,s}^{\mathrm{conf}})\neq\varnothing\qquad\text{and}\qquad\lambda\in\mathcal{G}_{n,s}\!\big[H_{n,s}\big]\cap(-\infty,\lambda_{0,s}^{\mathrm{conf}}).$
    In particular, for $s\in\left(0,\frac{n}{4}\right]\cap\mathcal{B}_{+}$, one has
    $(b_{s},\lambda_{0,s}^{\mathrm{conf}})\subset\mathcal{G}_{n,s}\!\big[H_{n,s}\big]\cap(-\infty,\lambda_{0,s}^{\mathrm{conf}}).$
Comments on the main difficulties. We conclude the introduction with several remarks on the analytic challenges inherent in our approach. The operators studied in this paper—namely, $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$—are genuine nonlocal fractional GJMS operators. Their analysis is significantly more delicate than that of their integer-order counterparts, owing to the absence of local structure and the associated loss of classical elliptic analysis tools. For instance, the lack of locality prevents us from reducing the problem to purely compactly supported computations: even when $u$ is compactly supported, $\mathcal{P}_{s}u$ (or $\widetilde{\mathcal{P}}_{s}u$) need not be compactly supported. This phenomenon is quantified by the off-diagonal behavior of the kernels (see Proposition 4.10).
Moreover, thanks to the nonlocality, classical integration-by-parts techniques fails in nonlocal cases, while such identities are usually crucial for energy estimates and for localization procedures in variational methods. To overcome this issue, we make systematic use of the pseudodifferential calculus on manifolds: this framework is well adapted to operators given by spectral multipliers, since it requires precisely the Fourier-side symbol estimates available for $\mathcal{P}_{s}$, $\widetilde{\mathcal{P}}_{s}$, and this allows us to establish the relevant boundedness and commutator properties and to justify the key localization estimate (see Section 4.2).
Finally, $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$ are defined via spectral functional calculus. As a consequence, the most accessible information is encoded in the frequency side through explicit multipliers, whereas direct control in the physical variable is not immediate. An explicit kernel representation of $\widetilde{\mathcal{P}}_{s}$ must be involved for certain priori estimates, see Proposition 4.10.
The remainder of this paper is organized as follows. In Section 2, we recall basic facts on the Helgason–Fourier transform and on fractional GJMS operators on the hyperbolic space. Section 3 is devoted to the Poincaré–Sobolev levels associated with $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$, and contains the proofs of Theorems 1.3 and 1.4. In Section 4, we analyze the attainability of these Poincaré–Sobolev levels and, as a consequence, establish the existence of positive solutions to the corresponding Brezis–Nirenberg problems, proving Theorems 1.7 and 1.8); we also provide the proofs of Theorems 1.5 and 1.6 for $H_{n,s}$ and $\widetilde{H}_{n,s}$. Finally, in the Appendix we present the proofs of Propositions 1.1 and 1.2.
2  Geometric and Analytic Preliminaries
2.1 Half-Space Model and Poincaré Ball Model
Throughout the paper, we set
$\rho:=\frac{n-1}{2},\quad\phi(x):=\frac{2}{1-|x|^{2}},\ \ n\geq 2,$
then the bottom of the $L^{2}$–spectrum equals $\rho^{2}$.
We first recall the two most classical models of the hyperbolic space. Let $\mathbb{H}^{n}$ denote the nn–dimensional hyperbolic space, realized in the upper half–space model
$\mathbb{H}^{n}:=\{x=(r,z)\in\mathbb{R}\times\mathbb{R}^{n-1}:r>0,\ z\in\mathbb{R}^{n-1}\}.$
It is endowed with the hyperbolic metric
$g_{\mathbb{H}^{n}}(r,z)=\frac{dr^{2}+|dz|^{2}}{r^{2}},$
and the corresponding volume element
$dV_{\mathbb{H}^{n}}(r,z)=r^{-n}\,dr\,dz.$
We write $\nabla_{\mathbb{H}^{n}}$ and $\Delta_{\mathbb{H}^{n}}$ for the hyperbolic gradient and Laplace–Beltrami operator, respectively. In the coordinates $(r,z)$, for any smooth function $u$, one has
$\nabla_{\mathbb{H}^{n}}u(r,z)=r^{2}\big(\partial_{r}u(r,z)\,\partial_{r}+\nabla_{z}u(r,z)\big),$
and therefore
$|\nabla_{\mathbb{H}^{n}}u(r,z)|_{g_{\mathbb{H}^{n}}}^{2}=r^{2}\Big(|\partial_{r}u(r,z)|^{2}+|\nabla_{z}u(r,z)|^{2}\Big),$
the Laplace–Beltrami operator is given by
$\Delta_{\mathbb{H}^{n}}=r^{2}\big(\partial_{r}^{2}+\Delta_{z}\big)-(n-2)\,r\,\partial_{r},$
where $\nabla_{z}$ and $\Delta_{z}$ denote the Euclidean gradient and Laplacian in the zz–variables.
Moreover, the hyperbolic space can also be identified with the unit ball
$\mathbb{B}^{n}:=\big\{x\in\mathbb{R}^{n}:\ |x|<1\big\}$
endowed with the metric
$g_{\mathbb{B}^{n}}=\frac{4\,(dx_{1}^{2}+\cdots+dx_{n}^{2})}{(1-|x|^{2})^{2}}=\phi(x)^{2}\,dx^{2}.$
The corresponding volume element is
$dV_{\mathbb{B}^{n}}(x)=\phi(x)^{n}\,dx.$
With this normalization, the Laplace–Beltrami operator in ball coordinates reads
$\Delta_{\mathbb{B}^{n}}=\frac{1-|x|^{2}}{4}\Big\{(1-|x|^{2})\sum_{i=1}^{n}\partial_{x_{i}x_{i}}+2(n-2)\sum_{i=1}^{n}x_{i}\partial_{x_{i}}\Big\}.$
The isometry group of the Poincaré ball model $(\mathbb{B}^{n},g_{\mathbb{B}^{n}})$ consists precisely of those Möbius transformations that preserve $\mathbb{B}^{n}$. Moreover, the hyperbolic volume measure $dV_{\mathbb{B}^{n}}$ is invariant under these transformations. For any $y\in\mathbb{B}^{n}$, define the Möbius transformation $T_{y}:\mathbb{B}^{n}\to\mathbb{B}^{n}$ by
$T_{y}(x):=\frac{|x-y|^{2}\,y-(1-|y|^{2})(x-y)}{1-2x\cdot y+|x|^{2}|y|^{2}}\quad{\rm for}\ \,x\in\mathbb{B}^{n}.$
A direct computation shows that
$|T_{y}(x)|^{2}=\frac{|x-y|^{2}}{1-2x\cdot y+|x|^{2}|y|^{2}},$
and consequently the hyperbolic distance between $x$ and $y$ admits the Euclidean representation
$\cosh d(x,y)=\frac{1+|T_{y}(x)|^{2}}{1-|T_{y}(x)|^{2}}=1+\frac{2|x-y|^{2}}{\left(1-|x|^{2}\right)\left(1-|y|^{2}\right)}.$
Equivalently, one has the distance formula
$d(x,y)=\log\frac{1+|T_{y}(x)|}{1-|T_{y}(x)|},\qquad x,y\in\mathbb{B}^{n}.$
Using Möbius transformations, we may define the convolution of measurable functions $f$ and $g$ on $\mathbb{B}^{n}$ by (see, e.g., [21])
$(f\ast g)(x):=\int_{\mathbb{B}^{n}}f(y)\,g\bigl(T_{x}(y)\bigr)\,dV_{\mathbb{B}^{n}}(y),$ (2.1)
whenever the integral is well defined.
2.2 Helgason Fourier Transform
In this subsection, we briefly recall the Helgason–Fourier analysis on the hyperbolic space, working in the Poincaré ball model $\bigl(\mathbb{B}^{n},g_{\mathbb{B}^{n}}\bigr)$. Since $(\mathbb{B}^{n},g_{\mathbb{B}^{n}})$ is a complete Riemannian manifold, the Laplace–Beltrami operator $\Delta_{\mathbb{B}^{n}}$ with initial domain $C_{c}^{\infty}(\mathbb{B}^{n})\subset L^{2}(\mathbb{B}^{n})$ is essentially self–adjoint on $C_{c}^{\infty}(\mathbb{B}^{n})$. We denote its unique self–adjoint extension again by $\Delta_{\mathbb{B}^{n}}$. The quadratic form associated with $\Delta_{\mathbb{B}^{n}}$ is given by
$a(u,v):=\int_{\mathbb{B}^{n}}\langle\nabla u,\nabla v\rangle_{g}\,dV_{\mathbb{B}^{n}},\quad u,v\in C_{c}^{\infty}(\mathbb{B}^{n}),$
and it extends by closure to a densely defined, closed, nonnegative form on $L^{2}(\mathbb{B}^{n})$ with form domain
$\mathcal{D}(a)=H^{1}(\mathbb{B}^{n}),$
where the integer-order Sobolev space on the hyperbolic space
$H^{k}(\mathbb{B}^{n}):=\Bigl\{u\in L^{2}(\mathbb{B}^{n})\;:\;\nabla_{\mathbb{B}^{n}}^{\,\alpha}u\in L^{2}(\mathbb{B}^{n})\ \text{for all multi-indices }\alpha\ \text{with }|\alpha|\leq k\Bigr\},\ k\in\mathbb{N}$
equipped with the norm
$\|u\|_{H^{k}(\mathbb{B}^{n})}^{2}:=\sum_{|\alpha|\leq k}\bigl\|\nabla_{\mathbb{B}^{n}}^{\,\alpha}u\bigr\|_{L^{2}(\mathbb{B}^{n})}^{2}.$
Equivalently, $H^{k}(\mathbb{B}^{n})$ is the completion of $C_{c}^{\infty}(\mathbb{B}^{n})$ with respect to the above norm. Thus, $H^{1}(\mathbb{B}^{n})$ is the natural energy space for $\Delta_{\mathbb{B}^{n}}$. The operator domain of $\Delta_{\mathbb{B}^{n}}$ is given by
$\mathcal{D}(-\Delta_{\mathbb{B}^{n}})=\bigl\{u\in H^{1}(\mathbb{B}^{n}):-\Delta_{\mathbb{B}^{n}}u\in L^{2}(\mathbb{B}^{n})\bigr\},$
which coincides with Sobolev space $H^{2}(\mathbb{B}^{n})$, see [19, Appendix B].
We set
$\lambda(\beta):=\beta^{2}+\rho^{2},\>\>\beta\in\mathbb{R}.$
The basic facts about the Fourier transform on the hyperbolic space and in the Poincaré ball model could see the references [21, 8, 15].
For $(\beta,\theta)\in\mathbb{R}\times\mathbb{S}^{n-1}$, set
$h_{\beta,\theta}(x):=\Bigl(\frac{\sqrt{1-|x|^{2}}}{|x-\theta|}\Bigr)^{\,n-1-2\beta{\bf i}},\quad\forall\,x\in\mathbb{B}^{n},$
it satisfies
$\Delta_{\mathbb{B}^{n}}h_{\beta,\theta}=\bigl(\beta^{2}+\rho^{2}\bigr)\,h_{\beta,\theta}.$ (2.2)
For $f\in C_{c}^{\infty}(\mathbb{B}^{n})$, the Helgason–Fourier transform is defined by
$\hat{f}(\beta,\theta):=\int_{\mathbb{B}^{n}}f(x)\,h_{\beta,\theta}(x)\,dV_{\mathbb{B}^{n}}(x),\quad(\beta,\theta)\in\mathbb{R}\times\mathbb{S}^{n-1}.$ (2.3)
This transform extends uniquely to a unitary operator
$\mathcal{F}:\ L^{2}(\mathbb{B}^{n})\longrightarrow L^{2}\!\Bigl(\mathbb{R}\times\mathbb{S}^{n-1},\frac{d\beta\,d\sigma(\theta)}{|c(\beta)|^{2}}\Bigr)$
where $d\sigma(\theta)$ is the normalized surface measure on $\mathbb{S}^{n-1}$ and the corresponding Plancherel identity reads
$\int_{\mathbb{B}^{n}}|f(x)|^{2}\,dV_{\mathbb{B}^{n}}(x)=\int_{\mathbb{R}}\int_{\mathbb{S}^{n-1}}|\hat{f}(\beta,\theta)|^{2}\,\frac{d\sigma(\theta)\,d\beta}{|c(\beta)|^{2}}\quad{\rm for\ all}\ \,f\in L^{2}(\mathbb{B}^{n}).$
Here $c(\beta)$ is the Harish–Chandra cc–function (see [21]):
$c(\beta):=2^{\frac{n-1}{2}}\,\pi^{\frac{n}{4}}\sqrt{\Gamma\!\left(\frac{n}{2}\right)}\,\frac{\Gamma({\bf i}\beta)}{\Gamma\!\left(\frac{n-1}{2}+{\bf i}\beta\right)}$
satisfying
$\left|c(\beta)\right|^{-2}=\frac{2^{1-n}}{\Gamma\left(\frac{n}{2}\right)\pi^{\frac{n}{2}}}\frac{\left|\Gamma\!\bigl({\bf i}\beta+\frac{n-1}{2}\bigr)\right|^{2}}{\left|\Gamma({\bf i}\beta)\right|^{2}}.$
Moreover, for $f\in C_{c}^{\infty}(\mathbb{B}^{n}),$ there is the inversion formula
$f(x)=\int_{\mathbb{R}}\int_{\mathbb{S}^{n-1}}\overline{h_{\beta,\theta}(x)}\,\hat{f}(\beta,\theta)\,\frac{d\sigma(\theta)\,d\beta}{|c(\beta)|^{2}},$ (2.4)
with convergence in $L^{2}(\mathbb{B}^{n})$.
A key feature of the Helgason transform is that it diagonalizes the Laplace Beltrami operator. A direct computation using (2.2) and (2.3) shows that,
$\widehat{\Delta_{\mathbb{B}^{n}}f}(\beta,\theta)=-\bigl(\beta^{2}+\rho^{2}\bigr)\,\hat{f}(\beta,\theta),\quad f\in C_{c}^{\infty}(\mathbb{B}^{n}).$
We next clarify the connection between the Laplace–Beltrami operator and multiplication operator under the Helgason transform, which will serve as the basis for identifying Fourier symbols of the nonlocal operators considered later.
Lemma 2.1.
Let $M_{\lambda}$ be the multiplication operator on $L^{2}\!\bigl(\mathbb{R}\times\mathbb{S}^{n-1},|c(\beta)|^{-2}d\beta d\sigma(\theta)\bigr):$
$(M_{\lambda}f)(\beta,\theta):=\lambda(\beta)\,f(\beta,\theta).$
Then, $\Delta_{\mathbb{B}^{n}}$ on $L^{2}(\mathbb{B}^{n})$ is unitarily equivalent to $M_{\lambda}$, that is,
$\Delta_{\mathbb{B}^{n}}=\mathcal{F}^{-1}M_{\lambda}\mathcal{F},$ (2.5)
with equality of domains. Equivalently, for every $f\in H^{2}(\mathbb{B}^{n})$, one has
$\widehat{-\Delta_{\mathbb{B}^{n}}f}(\beta,\theta)=\lambda(\beta)\,\hat{f}(\beta,\theta)\quad\text{for a.e. }(\beta,\theta)\in\mathbb{R}\times\mathbb{S}^{n-1}.$
Proof: Recall that the Helgason transform
$\mathcal{F}:L^{2}(\mathbb{B}^{n})\longrightarrow L^{2}\!\Bigl(\mathbb{R}\times\mathbb{S}^{n-1},\frac{d\beta\,d\sigma(\theta)}{|c(\beta)|^{2}}\Bigr)$
is unitary, with inverse given by the inversion formula, and for $f\in C_{c}^{\infty}(\mathbb{B}^{n}),$ we have
$\widehat{\Delta_{\mathbb{B}^{n}}f}(\beta,\theta)=-\bigl(\beta^{2}+\rho^{2}\bigr)\,\hat{f}(\beta,\theta)=-\lambda(\beta)\,\hat{f}(\beta,\theta).$ (2.6)
Define the operator
$\mathcal{T}:=\mathcal{F}^{-1}M_{\lambda}\mathcal{F}$
with domain
$\mathcal{D}(\mathcal{T}):=\Bigl\{f\in L^{2}(\mathbb{B}^{n}):\lambda(\beta)\,\hat{f}(\beta,\theta)\in L^{2}\!\Bigl(\mathbb{R}\times\mathbb{S}^{n-1},|c(\beta)|^{-2}d\beta d\sigma(\theta)\Bigr)\Bigr\}.$
Since $M_{\lambda}$ is a self–adjoint multiplication operator on the Helgason side and $\mathcal{F}$ is unitary, $\mathcal{T}$ is self–adjoint.
From (2.6), we see that
$\mathcal{T}f=-\Delta_{\mathbb{B}^{n}}f,\qquad f\in C_{c}^{\infty}(\mathbb{B}^{n}).$
On the other hand, on a complete Riemannian manifold $\mathbb{B}^{n}$, the operator $\Delta_{\mathbb{B}^{n}}$ with initial domain $C_{c}^{\infty}(\mathbb{B}^{n})$ admits a unique self–adjoint Friedrichs extension, we thus have two self–adjoint operators, $T$ and $\Delta_{\mathbb{B}^{n}}$, which coincide on the dense subspace $C_{c}^{\infty}(\mathbb{B}^{n})$:
$\mathcal{T}=-\Delta_{\mathbb{B}^{n}},$
and their domains agree: $\mathcal{D}(\mathcal{T})=\mathcal{D}(\Delta_{\mathbb{B}^{n}})$, which yields (2.5). ∎
For every Borel measurable function $\Phi:[\rho^{2},\infty)\to\mathbb{R},$ the operator $\Phi(-\Delta_{\mathbb{B}^{n}})$ defined by the functional calculus satisfies
$\Phi(-\Delta_{\mathbb{B}^{n}})=\mathcal{H}^{-1}M_{\Phi\circ\lambda}\,\mathcal{H}$,
where $M_{\Phi\circ\lambda}$ is the multiplication operator
$(M_{\Phi\circ\lambda}F)(\beta,\theta):=\Phi\bigl(\lambda(\beta)\bigr)\,F(\beta,\theta).$
In particular,
$\mathcal{D}\bigl(\Phi(-\Delta_{\mathbb{B}^{n}})\bigr)=\Bigl\{f\in L^{2}(\mathbb{B}^{n}):\Phi\bigl(\lambda(\beta)\bigr)\,\hat{f}(\beta,\theta)\in L^{2}\!\Bigl(\mathbb{R}\times\mathbb{S}^{n-1},|c(\beta)|^{-2}d\beta d\sigma(\theta)\Bigr)\Bigr\},$
and for $f\in\mathcal{D}(\Phi(-\Delta_{\mathbb{B}^{n}}))$ one has
$\widehat{\Phi(-\Delta_{\mathbb{B}^{n}})f}(\beta,\theta)=\Phi\bigl(\lambda(\beta)\bigr)\,\hat{f}(\beta,\theta),\quad f\in\mathcal{D}\bigl(\Phi(-\Delta_{\mathbb{B}^{n}})\bigr).$
Remark 2.2.
The spectral theorem for multiplication operators on $L^{2}$ implies that the spectrum of $M_{\varphi}$ is the essential range of its symbol $\varphi$, see [19, Appendix A]. Since $\mathcal{H}$ is unitary, the spectrum is invariant under this unitary equivalence,
$\sigma(-\Delta_{\mathbb{B}^{n}})=\sigma(M_{\lambda})=\operatorname{ess\,ran}\lambda=\operatorname{ess\,ran}\bigl(\beta^{2}+\rho^{2}\bigr).$
Similarly,
$\sigma\bigl(\Phi(-\Delta_{\mathbb{B}^{n}})\bigr)=\sigma\bigl(M_{\Phi\circ\lambda}\bigr)=\operatorname{ess\,ran}\bigl(\Phi\circ\lambda\bigr)=\operatorname{ess\,ran}\bigl(\Phi(\beta^{2}+\rho^{2})\bigr).$
In particular, if $\Phi$ is continuous, then $\sigma\bigl(\Phi(-\Delta_{\mathbb{B}^{n}})\bigr)=\overline{\Phi\bigl(\sigma(-\Delta_{\mathbb{B}^{n}})\bigr)}$.
We now specialize to the choice $\Phi(\lambda)=\lambda^{s},\ s>0$, see [6]. For $s>0,$ we define the spectral fractional Sobolev space
$H^{2s}(\mathbb{B}^{n}):=\Bigl\{f\in L^{2}(\mathbb{B}^{n}):(\beta^{2}+\rho^{2})^{s}\,\hat{f}(\beta,\theta)\in L^{2}\!\Bigl(\mathbb{R}\times\mathbb{S}^{n-1},|c(\beta)|^{-2}d\beta d\sigma(\theta)\Bigr)\Bigr\},$ (2.7)
endowed with the norm
$\|f\|_{H^{2s}(\mathbb{B}^{n})}^{2}:=\int_{\mathbb{R}}\int_{\mathbb{S}^{n-1}}\bigl(1+(\beta^{2}+\rho^{2})^{2s}\bigr)\,|\hat{f}(\beta,\theta)|^{2}\,\frac{d\sigma(\theta)\,d\beta}{|c(\beta)|^{2}}.$
In particular, for an integer $k\in\mathbb{N}$, the fractional Sobolev space $H^{s}(\mathbb{B}^{n})$ with $s=k$ coincides with the classical Sobolev space $H^{k}(\mathbb{B}^{n})$, with equivalent norms. Thus, the spectral fractional Laplacian satisfies
$\widehat{(-\Delta_{\mathbb{B}^{n}})^{s}f}(\beta,\theta)=\bigl(\beta^{2}+\rho^{2}\bigr)^{s}\,\hat{f}(\beta,\theta),\quad\ f\in H^{2s}(\mathbb{B}^{n}).$
2.3 Fractional GJMS Operators and Inequality
For $\gamma\in\bigl(0,\frac{n}{2}\bigr)\setminus\mathbb{N}$, recall the conformal fractional GJMS operator $\mathcal{P}_{\gamma}$ on $\mathbb{H}^{n}$ has the explicit spectral representation (1.2). Equivalently, under the Helgason–Fourier transform, $P_{\gamma}$ acts as a spectral multiplier:
$\widehat{\mathcal{P}_{\gamma}f}(\beta,\theta)=m_{\gamma}(\beta)\,\hat{f}(\beta,\theta),\qquad m_{\gamma}(\beta):=2^{2\gamma}\,\frac{\bigl|\Gamma\!\bigl(\frac{3+2\gamma}{4}+\frac{{\bf i}}{2}\beta\bigr)\bigr|^{2}}{\bigl|\Gamma\!\bigl(\frac{3-2\gamma}{4}+\frac{{\bf i}}{2}\beta\bigr)\bigr|^{2}}.$
However, $P_{\gamma}$ does not satisfy the following intertwining relations (1.3) (1.4). To recover a direct intertwining with the Euclidean fractional Laplacian under the conformal identifications of $\mathbb{H}^{n}$ with $\mathbb{R}^{n}_{+}$ and $\mathbb{B}^{n}$, Lu et al. introduced an auxiliary family of fractional operators $\widetilde{\mathcal{P}}_{\gamma}$ on $\mathbb{H}^{n}$ in [24, Theorem 1.7]. Equivalently, under the Helgason–Fourier transform one has the multiplier representation
$\widehat{\widetilde{\mathcal{P}}_{\gamma}f}(\beta,\theta)=\widetilde{m}_{\gamma}(\beta)\,\hat{f}(\beta,\theta),\qquad\widetilde{m}_{\gamma}(\beta):=\frac{\bigl|\Gamma\!\bigl(\gamma+\frac{1}{2}+{\bf i}\beta\bigr)\bigr|^{2}}{\bigl|\Gamma\!\bigl(\frac{1}{2}+{\bf i}\beta\bigr)\bigr|^{2}}.$ (2.8)
We denote the bottoms of the $L^{2}$ spectra by
$\lambda_{0,s}^{\mathrm{conf}}:=\inf\sigma(\mathcal{P}_{s}),\qquad\widetilde{\lambda}_{0,s}^{\mathrm{conf}}:=\inf\sigma(\widetilde{\mathcal{P}}_{s}).$
Since $A$ has spectrum $\sigma(A)=[0,\infty)$ and both $P_{s}$ and $\widetilde{\mathcal{P}}_{s}$ are spectral multipliers of $A$, we have
$\lambda_{0,s}^{\mathrm{conf}}=\inf_{\beta\geq 0}m_{s}(\beta),\qquad\widetilde{\lambda}_{0,s}^{\mathrm{conf}}=\inf_{\beta\geq 0}\widetilde{m}_{s}(\beta),$ (2.9)
By the classical inequality (see [11, p. 904, Eq. (8.236)]):
$\bigl|\Gamma(a+{\bf i}\lambda)\bigr|^{2}=\bigl|\Gamma(a)\bigr|^{2}\prod_{k=0}^{\infty}\frac{1}{1+\dfrac{\lambda^{2}}{(a+k)^{2}}},\qquad a\in\mathbb{R},\ a\neq 0,-1,-2,\ldots$
we can obtain
$\frac{\bigl|\Gamma(a+{\bf i}\lambda)\bigr|}{\bigl|\Gamma(b+{\bf i}\lambda)\bigr|}\geq\frac{\bigl|\Gamma(a)\bigr|}{\bigl|\Gamma(b)\bigr|},\qquad\text{provided that }\ (a+k)^{2}\geq(b+k)^{2},\ \forall k\in\mathbb{N}.$
Thus,
$\lambda_{0,s}^{\mathrm{conf}}=m_{s}(0)=2^{2s}\,\frac{\Gamma\!\bigl(\frac{3+2s}{4}\bigr)^{2}}{\Gamma\!\bigl(\frac{3-2s}{4}\bigr)^{2}},\qquad\widetilde{\lambda}_{0,s}^{\mathrm{conf}}=\widetilde{m}_{s}(0)=\frac{\Gamma\!\bigl(s+\frac{1}{2}\bigr)^{2}}{\Gamma\!\bigl(\frac{1}{2}\bigr)^{2}}.$
By [24, Lemma 5.2], we have
$\lambda_{0,s}^{\mathrm{conf}}-\widetilde{\lambda}_{0,s}^{\mathrm{conf}}=\frac{\sin(\pi s)}{\pi}\,\Gamma\bigl(s+\tfrac{1}{2}\bigr)^{2}.$ (2.10)
Since the Helgason–Fourier transform diagonalizes $\mathcal{P}_{k}$:
$\mathcal{P}_{k}=\mathcal{P}_{1}\,(\mathcal{P}_{1}+2)\cdots\bigl(\mathcal{P}_{1}+k(k-1)\bigr)=\prod_{j=1}^{k}\bigl(\mathcal{A}^{2}+(j-\frac{1}{2})^{2}\bigr).$ (2.11)
that is, for every $f\in C_{c}^{\infty}(\mathbb{H}^{n})$,
$\widehat{\mathcal{P}_{k}f}(\beta,\theta)=m_{k}(\beta)\,\hat{f}(\beta,\theta),\quad\forall(\beta,\theta)\in\mathbb{R}\times\mathbb{S}^{n-1},$ (2.12)
with radial multiplier
$m_{k}(\beta)=\prod_{j=1}^{k}\bigl(\beta^{2}+(j-\frac{1}{2})^{2}\bigr).$ (2.13)
Using $\beta^{2}+\bigl(j-\tfrac{1}{2}\bigr)^{2}=\bigl(j-\tfrac{1}{2}+{\bf i}\beta\bigr)\bigl(j-\tfrac{1}{2}-{\bf i}\beta\bigr)$ and the identity $\prod_{j=1}^{k}(z+j-\frac{1}{2})=\Gamma(z+k+\frac{1}{2})/\Gamma(z+\frac{1}{2})$, we obtain the equivalent Gamma–function closed form
$m_{k}(\beta)=\frac{\Gamma\!\bigl(k+\frac{1}{2}+{\bf i}\beta\bigr)}{\Gamma\!\bigl(\frac{1}{2}+{\bf i}\beta\bigr)}\frac{\Gamma\!\bigl(k+\frac{1}{2}-{\bf i}\beta\bigr)}{\Gamma\!\bigl(\frac{1}{2}-{\bf i}\beta\bigr)}=\frac{\bigl|\Gamma\!\bigl(k+\frac{1}{2}+{\bf i}\beta\bigr)\bigr|^{2}}{\bigl|\Gamma\!\bigl(\frac{1}{2}+{\bf i}\beta\bigr)\bigr|^{2}}.$ (2.14)
In particular, this coincides with the multiplier $\widetilde{m}_{\gamma}(\beta)$ of $\widetilde{\mathcal{P}}_{\gamma}$ evaluated at $\gamma=k$.
Moreover, when $\gamma=k\in\mathbb{N}$, the bottom of the spectrum satisfies
$\widetilde{\lambda}_{0,k}^{\mathrm{conf}}=\frac{\Gamma\!\bigl(k+\tfrac{1}{2}\bigr)^{2}}{\Gamma\!\bigl(\tfrac{1}{2}\bigr)^{2}}=\prod_{j=1}^{k}\Bigl(j-\tfrac{1}{2}\Bigr)^{\!2}=\prod_{j=1}^{k}\frac{(2j-1)^{2}}{4},$
which equals the bottom of the spectrum of $\mathcal{P}_{k}$.
It is well known that the sharp fractional Sobolev constant in the Euclidean setting is independent of the underlying domain. In particular, applying (1.12) to the unit ball $\mathbb{B}^{n}$, we obtain
$S_{n,s}\Bigl(\int_{\mathbb{B}^{n}}|v|^{2^{*}_{s}}\,dx\Bigr)^{\!\frac{2}{2^{*}_{s}}}\leq\int_{\mathbb{B}^{n}}v\,(-\Delta)^{s}v\,dx,\quad\forall v\in C_{c}^{\infty}(\mathbb{B}^{n}).$ (2.15)
By conformal invariance, the Euclidean inequality (2.15) is equivalent to a hyperbolic inequality involving $\widetilde{\mathcal{P}}_{s}$. For the integer-order cases $s=1$ and, more generally, $s=k\geq 2$, we refer to alternative proofs in [14, 22]. In the fractional setting, Lu and Yang [24] established the result for $n\geq 3$ and $\frac{n-1}{2}\leq s<\tfrac{n}{2}$. Below we present a unified proof valid for all $s\in(0,\tfrac{n}{2})$.
Proposition 2.3.
Let $0<s<\tfrac{n}{2}$. Then the following Sobolev inequality holds:
$S_{n,s}\Bigl(\int_{\mathbb{H}^{n}}|u|^{2^{*}_{s}}\,dV_{\mathbb{H}^{n}}\Bigr)^{\!\frac{2}{2^{*}_{s}}}\;\leq\;\int_{\mathbb{H}^{n}}u\,\widetilde{\mathcal{P}}_{s}u\,dV_{\mathbb{H}^{n}},\qquad u\in C_{c}^{\infty}(\mathbb{H}^{n}),$ (2.16)
where $2_{s}^{*}=\frac{2n}{n-2s}$ and $S_{n,s}$ is defined in (1.12). Moreover, the constant $S_{n,s}$ in (2.16) is optimal.
Proof: Without loss of generality, we work in the Poincaré ball model $(\mathbb{B}^{n},g_{\mathbb{B}^{n}})$ of the hyperbolic space, where
$g_{\mathbb{B}^{n}}=\phi(x)^{2}\,dx^{2},\qquad\phi(x):=\frac{2}{1-|x|^{2}}.$
For $u\in C_{c}^{\infty}(\mathbb{B}^{n})$, we introduce the critical conformal transformation
$v:=\phi^{\frac{n-2s}{2}}u\in C_{c}^{\infty}(\mathbb{B}^{n}).$
By the exact intertwining identity (1.4) for the operator $\widetilde{\mathcal{P}}_{s}$, we have
$\widetilde{\mathcal{P}}_{s}u\;=\;\phi^{-\frac{n+2s}{2}}\,(-\Delta)^{s}\!\bigl(\phi^{\frac{n-2s}{2}}u\bigr)\;=\;\phi^{-\frac{n+2s}{2}}\,(-\Delta)^{s}v,$
and since $dV_{{\mathbb{B}^{n}}}=\phi(x)^{n}\,dx$, we obtain the exact energy identity
$\int_{\mathbb{B}^{n}}u\,\widetilde{\mathcal{P}}_{s}u\,dV_{\mathbb{B}^{n}}\;=\;\int_{\mathbb{B}^{n}}v\,(-\Delta)^{s}v\,dx.$ (2.17)
At the critical exponent $2_{s}^{*}=\frac{2n}{n-2s}$ one also has the conformal invariance of the $L^{2_{s}^{*}}$ norm:
$\int_{\mathbb{B}^{n}}|u|^{2_{s}^{*}}\,dV_{\mathbb{B}^{n}}\;=\;\int_{\mathbb{B}^{n}}|v|^{2_{s}^{*}}\,dx.$ (2.18)
Applying the sharp Euclidean fractional Sobolev inequality to $v$ and then using (2.17)–(2.18) yields
$S_{n,s}\Bigl(\int_{\mathbb{B}^{n}}|u|^{2_{s}^{*}}\,dV_{\mathbb{B}^{n}}\Bigr)^{\!2/2_{s}^{*}}\;\leq\;\int_{\mathbb{B}^{n}}v\,(-\Delta)^{s}v\,dx\;=\;\int_{\mathbb{B}^{n}}u\,\widetilde{\mathcal{P}}_{s}u\,dV_{\mathbb{B}^{n}},$
which is (2.16).
To prove sharpness, let $\{w_{\varepsilon}\}\subset C_{c}^{\infty}(\mathbb{B}^{n})\setminus\left\{0\right\}$ be a sequence satisfying
$\frac{\displaystyle\int_{\mathbb{B}^{n}}w_{\varepsilon}\,(-\Delta)^{s}w_{\varepsilon}\,dx}{\Bigl(\displaystyle\int_{\mathbb{B}^{n}}|w_{\varepsilon}|^{2_{s}^{*}}\,dx\Bigr)^{\!2/2_{s}^{*}}}\longrightarrow S_{n,s}\quad\varepsilon\rightarrow 0.$
Define $u_{\varepsilon}\in C_{c}^{\infty}(\mathbb{B}^{n})$ by $u_{\varepsilon}:=\phi^{-\frac{n-2s}{2}}\,w_{\varepsilon}$, so that $v_{\varepsilon}=\phi^{\frac{n-2s}{2}}u_{\varepsilon}=w_{\varepsilon}$ on $\mathbb{B}^{n}$. Then by (2.17)–(2.18),
$\frac{\displaystyle\int_{\mathbb{B}^{n}}u_{\varepsilon}\,\widetilde{\mathcal{P}}_{s}u_{\varepsilon}\,dV_{\mathbb{B}^{n}}}{\Bigl(\displaystyle\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2_{s}^{*}}\,dV_{\mathbb{B}^{n}}\Bigr)^{\!2/2_{s}^{*}}}=\frac{\displaystyle\int_{\mathbb{B}^{n}}w_{\varepsilon}\,(-\Delta)^{s}w_{\varepsilon}\,dx}{\Bigl(\displaystyle\int_{\mathbb{B}^{n}}|w_{\varepsilon}|^{2_{s}^{*}}\,dx\Bigr)^{\!2/2_{s}^{*}}}\longrightarrow S_{n,s}.$
Hence the best constant on $\mathbb{B}^{n}$ equals $S_{n,s}$, proving optimality. ∎
3  Analysis of Fractional Poincaré–Sobolev Levels
3.1 Asymptotic Analysis
In this subsection, without loss of generality, we continue working in the ball model of $\mathbb{B}^{n}$ and recall
$2_{s}^{*}=\frac{2n}{n-2s},\quad\phi(x)=\frac{1-|x|^{2}}{2},\quad dV_{\mathbb{B}^{n}}=\phi(x)^{-n}\,dx\quad\text{on }\mathbb{B}^{n}.$
Lemma 3.1.
Let $u\in C_{c}^{\infty}(\mathbb{B}^{n})$ and let
$w(x):=\phi(x)^{s-\frac{n}{2}}u(x)\quad\text{for }x\in\mathbb{B}^{n},\qquad w(x)=0\quad\text{for }x\in\mathbb{R}^{n}\setminus\mathbb{B}^{n}.$
Then $w\in C_{c}^{\infty}(\mathbb{R}^{n})$ and the following identities hold:
$\int_{\mathbb{B}^{n}}(\widetilde{\mathcal{P}}_{s}u)\,u\,dV_{{\mathbb{B}}^{n}}=\int_{\mathbb{R}^{n}}w\,(-\Delta)^{s}w\,dx,\quad\int_{\mathbb{B}^{n}}|u|^{2_{s}^{*}}\,dV_{{\mathbb{B}}^{n}}=\int_{\mathbb{R}^{n}}|w|^{2_{s}^{*}}\,dx.$
Proof: The argument is identical to that of Proposition 2.3, hence we omit the details. ∎
Let
$U(x):=(1+|x|^{2})^{-\frac{n-2s}{2}},\qquad U_{\varepsilon}(x):=\varepsilon^{-\frac{n-2s}{2}}U\Bigl(\frac{x}{\varepsilon}\Bigr),$
and fix $0<\delta<\frac{1}{4}$. Choose $\eta\in C_{c}^{\infty}(\mathbb{B}^{n})$ such that $0\leq\eta\leq 1$, $\eta\equiv 1$ on $B_{\delta}(0)$ and $\eta\equiv 0$ on $\mathbb{B}^{n}\setminus B_{2\delta}(0)$. Define
$w_{\varepsilon}:=\eta\,U_{\varepsilon}\in C_{c}^{\infty}(\mathbb{R}^{n}),\qquad u_{\varepsilon}:=\phi^{\frac{n}{2}-s}w_{\varepsilon}\in C_{c}^{\infty}(\mathbb{B}^{n}).$
Proposition 3.2.
Let $0<s<\frac{n}{2}$. Then
$\int_{\mathbb{R}^{n}}|w_{\varepsilon}|^{2_{s}^{*}}\,dx=\int_{\mathbb{R}^{n}}|U|^{2_{s}^{*}}\,dx+O(\varepsilon^{n})\quad{\rm as}\ \,\varepsilon\to 0,$
where the $O(\varepsilon^{n})$ constant depends only on $n,s$.
Proof: Write $p:=2_{s}^{*}=\frac{2n}{n-2s}$. Since $\frac{n-2s}{2}\,p=n$, we have the exact identity
$|U_{\varepsilon}(x)|^{p}=\varepsilon^{-n}\Bigl(1+\frac{|x|^{2}}{\varepsilon^{2}}\Bigr)^{-n}.$
By the change of variables $x=\varepsilon y$,
$\int_{\mathbb{R}^{n}}|U_{\varepsilon}(x)|^{p}\,dx=\int_{\mathbb{R}^{n}}\varepsilon^{-n}(1+|x|^{2}/\varepsilon^{2})^{-n}\,dx=\int_{\mathbb{R}^{n}}(1+|y|^{2})^{-n}\,dy=\int_{\mathbb{R}^{n}}|U(y)|^{p}\,dy.$
Since $w_{\varepsilon}=\eta U_{\varepsilon}$ and $0\leq\eta\leq 1$,
$\int_{\mathbb{R}^{n}}|w_{\varepsilon}|^{p}\,dx=\int_{\mathbb{R}^{n}}\eta(x)^{p}|U_{\varepsilon}(x)|^{p}\,dx=\int_{\mathbb{R}^{n}}|U_{\varepsilon}|^{p}\,dx-\int_{\mathbb{R}^{n}}(1-\eta^{p})|U_{\varepsilon}|^{p}\,dx.$
Moreover, $1-\eta^{p}$ is supported in $\mathbb{R}^{n}\setminus B_{\delta}(0)$, hence
$\int_{\mathbb{R}^{n}}(1-\eta^{p})|U_{\varepsilon}|^{p}\,dx\sim\int_{|x|\geq\delta}|U_{\varepsilon}(x)|^{p}\,dx.$
Using again $x=\varepsilon y$,
$\int_{|x|\geq\delta}|U_{\varepsilon}(x)|^{p}\,dx=\int_{|y|\geq\delta/\varepsilon}(1+|y|^{2})^{-n}\,dy.$
For $|y|\geq 1$ we have $(1+|y|^{2})^{-n}\leq|y|^{-2n}$, and thus by polar coordinates, for $R\geq 1$,
$\int_{|y|\geq R}(1+|y|^{2})^{-n}\,dy\leq|\mathbb{S}^{n-1}|\int_{R}^{\infty}r^{n-1}r^{-2n}\,dr=\frac{|\mathbb{S}^{n-1}|}{n}\,R^{-n}.$
Taking $R=\delta/\varepsilon$ yields
$\int_{|x|\geq\delta}|U_{\varepsilon}(x)|^{p}\,dx=O(\varepsilon^{n}).$
which gives
$\int_{\mathbb{R}^{n}}|w_{\varepsilon}|^{p}\,dx=\int_{\mathbb{R}^{n}}|U_{\varepsilon}|^{p}\,dx+O(\varepsilon^{n})=\int_{\mathbb{R}^{n}}|U|^{p}\,dx+O(\varepsilon^{n}).$
Proposition 3.3.
Let $0<s<\frac{n}{2}$. Then as $\varepsilon\to 0^{+}$
$\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2}\,dV_{{\mathbb{B}}^{n}}=\begin{cases}c_{1,s}\,\varepsilon^{2s}\;+\;O(\varepsilon^{n-2s}),&n>4s,\\[5.69054pt] c_{2,s}\,\varepsilon^{2s}|\log\varepsilon|\;+\;O(\varepsilon^{2s}),&n=4s,\\[5.69054pt] c_{3,s}\,\varepsilon^{n-2s}\;+\;O(\varepsilon^{2s}),&2s<n<4s,\end{cases}$
with $c_{i,s}>0,i=1,2,3$ depending only on $n,s$.
Proof: Using $u_{\varepsilon}=\phi^{\frac{n}{2}-s}w_{\varepsilon}$ and $dV_{{\mathbb{B}}^{n}}=\phi^{-n}dx$,
$\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2}\,dV_{{\mathbb{B}}^{n}}=\int_{\mathbb{R}^{n}}|w_{\varepsilon}(x)|^{2}\,\phi(x)^{-2s}\,dx.$
Since $\operatorname{supp}w_{\varepsilon}\subset B_{2\delta}(0)$ and $\phi$ is smooth and positive there,
$\phi(x)^{-2s}=\phi(0)^{-2s}+O(x^{2})\qquad\text{for }|x|\leq 2\delta.$
Hence
$\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2}\,dV_{{\mathbb{B}}^{n}}=\phi(0)^{-2s}\int_{\mathbb{R}^{n}}|w_{\varepsilon}|^{2}\,dx\;+\;O\ \!\Bigl(\int_{\mathbb{R}^{n}}|x|^{2}|w_{\varepsilon}(x)|^{2}\,dx\Bigr).$ (3.1)
We firstly prove that:
$\int_{\mathbb{R}^{n}}|w_{\varepsilon}|^{2}\,dx=\begin{cases}c_{1,s}\,\varepsilon^{2s}\;+\;O(\varepsilon^{n-2s}),&n>4s,\\[5.69054pt] c_{2,s}\,\varepsilon^{2s}|\log\varepsilon|\;+\;O(\varepsilon^{2s}),&n=4s,\\[5.69054pt] c_{3,s}\,\varepsilon^{n-2s}\;+\;O(\varepsilon^{2s}),&2s<n<4s,\end{cases}$ (3.2)
with $c_{i,s}>0,i=1,2,3$ depending only on $n,s$.
Since $w_{\varepsilon}=\eta U_{\varepsilon}$ and $\eta\equiv 1$ on $B_{\delta}(0)$,
$\int_{\mathbb{R}^{n}}|w_{\varepsilon}|^{2}\,dx=\int_{B_{\delta}(0)}|U_{\varepsilon}|^{2}\,dx+\int_{B_{2\delta}(0)\setminus B_{\delta}(0)}|\eta U_{\varepsilon}|^{2}\,dx.$ (3.3)
The second term is always of order $O(\varepsilon^{\,n-2s})$: indeed, on $B_{2\delta}(0)\setminus B_{\delta}(0)$, we have $|x|\geq\delta$, hence
$|U_{\varepsilon}(x)|=\varepsilon^{-\frac{n-2s}{2}}\Bigl(1+\frac{|x|^{2}}{\varepsilon^{2}}\Bigr)^{-\frac{n-2s}{2}}\sim\,\varepsilon^{-\frac{n-2s}{2}}\Bigl(\frac{|x|}{\varepsilon}\Bigr)^{-(n-2s)}\sim\,\varepsilon^{\frac{n-2s}{2}},$
and therefore
$\int_{B_{2\delta}\setminus B_{\delta}}|\eta U_{\varepsilon}|^{2}\,dx=O(\varepsilon^{\,n-2s}).$ (3.4)
For the main part, use the change of variables $x=\varepsilon y$:
$\int_{B_{\delta}(0)}|U_{\varepsilon}(x)|^{2}\,dx=\varepsilon^{2s}\int_{B_{\delta/\varepsilon}(0)}(1+|y|^{2})^{-(n-2s)}\,dy.$ (3.5)
The asymptotics of the integral as $\varepsilon\to 0$ depend on the integrability at infinity of $(1+|y|^{2})^{-(n-2s)}$:
$(1+|y|^{2})^{-(n-2s)}\sim|y|^{-2(n-2s)}\quad|y|\to\infty.$
Hence:
  • If $n>4s$, so the function is integrable at infinity, and dominated convergence yields
    $\int_{B_{\delta/\varepsilon}}(1+|y|^{2})^{-(n-2s)}dy=\int_{\mathbb{R}^{n}}(1+|y|^{2})^{-(n-2s)}dy+O(\varepsilon^{\,n-4s}).$
    Plugging into (3.5) gives
    $\int_{B_{\delta}}|U_{\varepsilon}|^{2}dx=a_{n,s}\,\varepsilon^{2s}+O(\varepsilon^{\,n-2s}),$
    where
    $a_{n,s}:=\int_{\mathbb{R}^{n}}(1+|y|^{2})^{-(n-2s)}\,dy>0.$
  • If $n=4s$, then 2(n2s)=n2(n-2s)=n, so the tail behaves like $|y|^{-n}$:
    $\int_{B_{\delta/\varepsilon}}(1+|y|^{2})^{-(n-2s)}dy=c_{1}\,|\log\varepsilon|+O(1),$
    for some $c_{1}>0$. Thus
    $\int_{B_{\delta}}|U_{\varepsilon}|^{2}dx=c_{1}\,\varepsilon^{2s}|\log\varepsilon|+O(\varepsilon^{2s}).$
  • If $2s<n<4s$, then 2(n2s)<n2(n-2s)<n, using polar coordinates one finds
    $\int_{B_{\delta/\varepsilon}}(1+|y|^{2})^{-(n-2s)}dy=c_{2}\,(\delta/\varepsilon)^{\,4s-n}+O(1),$
    hence
    $\int_{B_{\delta}}|U_{\varepsilon}|^{2}dx=c_{2}\,\varepsilon^{\,n-2s}+O(\varepsilon^{2s}).$
Combining these with (3.4) and (3.3) yields exactly the three regimes (3.2).
Finally, we show that the error term in (3.1) is negligible. Using again the scaling $x=\varepsilon y$, we obtain
$\int_{\mathbb{R}^{n}}|x|^{2}|w_{\varepsilon}(x)|^{2}\,dx$ $=\int_{B_{2\delta}}|x|^{2}|U_{\varepsilon}(x)|^{2}\,dx=\varepsilon^{2s+2}\int_{B_{2\delta/\varepsilon}}|y|^{2}(1+|y|^{2})^{-(n-2s)}\,dy.$ (3.6)
A similar polar-coordinate computation gives the following growth estimate:
$\int_{B_{2\delta/\varepsilon}}|y|^{2}(1+|y|^{2})^{-(n-2s)}\,dy=\begin{cases}O(1),&n>4s+2,\\[2.84526pt] O\bigl(|\log\varepsilon|\bigr),&n=4s+2,\\[2.84526pt] O\bigl(\varepsilon^{-(4s-n+2)}\bigr),&n<4s+2,\end{cases}\qquad\varepsilon\to 0.$ (3.7)
Combining (3.6) and (3.7) we conclude
$\int_{\mathbb{R}^{n}}|x|^{2}|w_{\varepsilon}(x)|^{2}\,dx=\begin{cases}O(\varepsilon^{2s+2}),&n>4s+2\\[2.84526pt] O\bigl(\varepsilon^{2s+2}|\log\varepsilon|\bigr),&n=4s+2\\[2.84526pt] O(\varepsilon^{n-2s}),&n<4s+2\end{cases}\qquad{\rm as}\ \,\varepsilon\to 0.$
In particular, in all cases the remainder term in (3.1) is of higher order and hence negligible compared with the leading term in (3.2) ∎
Lemma 3.4.
Let $0<s<\frac{n}{2}$ and set $a=\frac{n-2s}{2}$. Define
$U(x)=(1+|x|^{2})^{-a},\qquad U_{\varepsilon}(x)=\varepsilon^{-a}\,U\!\left(\frac{x}{\varepsilon}\right).$
Then for every multi-index $\alpha\in\mathbb{N}_{0}^{n}$ and every $\delta>0$ there exists $C_{\alpha,\delta}>0$ such that for all $\varepsilon\in(0,1)$ and all $x\in\mathbb{R}^{n}$ with $|x|\geq\delta$,
$|\partial^{\alpha}U_{\varepsilon}(x)|\leq C_{\alpha,\delta}\,\varepsilon^{\frac{n-2s}{2}}.$ (3.8)
Proof: By the chain rule, for any multi-index $\alpha$,
$\partial^{\alpha}U_{\varepsilon}(x)=\varepsilon^{-a-|\alpha|}\,(\partial^{\alpha}U)\!\left(\frac{x}{\varepsilon}\right).$ (3.9)
We claim that there exists $C_{\alpha}>0$ such that
$|\partial^{\alpha}U(y)|\leq C_{\alpha}\,(1+|y|)^{-(n-2s+|\alpha|)}\qquad\forall\,y\in\mathbb{R}^{n}.$ (3.10)
Indeed, there exist a polynomial $P_{\alpha}$ on $\mathbb{R}^{n}$ and an integer $m=m(\alpha)\in\mathbb{N}_{0}$ such that
$\partial^{\alpha}(1+|y|^{2})^{-a}=\frac{P_{\alpha}(y)}{(1+|y|^{2})^{a+m}},$
with
$\deg P_{\alpha}\leq|\alpha|\qquad\text{and}\qquad m\geq\Bigl\lceil\frac{|\alpha|}{2}\Bigr\rceil.$
Consequently, for $|y|\geq 1$,
$|\partial^{\alpha}U(y)|\leq C_{\alpha}\,\frac{|y|^{|\alpha|}}{|y|^{2(a+m)}}\leq C_{\alpha}\,|y|^{-2a-2\lceil|\alpha|/2\rceil+|\alpha|}\leq C_{\alpha}\,|y|^{-(2a+|\alpha|)}=C_{\alpha}\,|y|^{-(n-2s+|\alpha|)},$
and (3.10) follows. Combining (3.9) and (3.10) yields
$|\partial^{\alpha}U_{\varepsilon}(x)|\leq\varepsilon^{-a-|\alpha|}\,C_{\alpha}\Bigl(1+\Bigl|\frac{x}{\varepsilon}\Bigr|\Bigr)^{-(n-2s+|\alpha|)}\leq C_{\alpha}\,\varepsilon^{-a-|\alpha|}\Bigl|\frac{x}{\varepsilon}\Bigr|^{-(n-2s+|\alpha|)}=C_{\alpha}\,|x|^{-(n-2s+|\alpha|)}\,\varepsilon^{a}$
for $\varepsilon$ small enough. Since $|x|\geq\delta$, we have $|x|^{-(n-2s+|\alpha|)}\leq\delta^{-(n-2s+|\alpha|)}$, and therefore
$|\partial^{\alpha}U_{\varepsilon}(x)|\leq C_{\alpha,\delta}\,\varepsilon^{a}=C_{\alpha,\delta}\,\varepsilon^{\frac{n-2s}{2}},$
which proves (3.8). ∎
We next establish a key energy asymptotic expansion for the truncated bubbles. Unlike the case $s\in(0,1)$ treated in [27], we cannot rely on a pointwise singular-integral representation of the fractional Laplacian to carry out the estimate, since such a representation is no longer available in the present range of $s$. Instead, we work with the Fourier definition of $(-\Delta)^{s}$ to derive the desired asymptotics.
Proposition 3.5.
Let $0<s<\frac{n}{2}$. Then, as $\varepsilon\to 0$,
$\int_{\mathbb{R}^{n}}w_{\varepsilon}\,(-\Delta)^{s}w_{\varepsilon}\,dx=\bigl\|(-\Delta)^{\frac{s}{2}}U\bigr\|_{L^{2}(\mathbb{R}^{n})}^{2}\;+\;O(\varepsilon^{\,n-2s}).$ (3.11)
In particular,
$\int_{\mathbb{R}^{n}}w_{\varepsilon}\,(-\Delta)^{s}w_{\varepsilon}\,dx=S_{n,s}\Bigl(\int_{\mathbb{R}^{n}}|U|^{2_{s}^{*}}\,dx\Bigr)^{\!\frac{2}{2_{s}^{*}}}\;+\;O(\varepsilon^{\,n-2s}).$ (3.12)
Proof: Throughout the proof, we use the Fourier definition
$\widehat{(-\Delta)^{s}f}(\xi)=|\xi|^{2s}\hat{f}(\xi),\qquad\int_{\mathbb{R}^{n}}f\,(-\Delta)^{s}f\,dx=\int_{\mathbb{R}^{n}}|\xi|^{2s}|\hat{f}(\xi)|^{2}\,d\xi=\bigl\|(-\Delta)^{\frac{s}{2}}f\bigr\|_{L^{2}(\mathbb{R}^{n})}^{2}.$
Set
$z_{\varepsilon}:=w_{\varepsilon}-U_{\varepsilon}=(\eta-1)U_{\varepsilon},$
so that $z_{\varepsilon}\equiv 0$ on $B_{\delta}(0)$ and $\operatorname{supp}(z_{\varepsilon})\subset\mathbb{R}^{n}\setminus B_{\delta}(0)$. Then, since $w_{\varepsilon}\in C_{c}^{\infty}(\mathbb{R}^{n}),$
$\int w_{\varepsilon}(-\Delta)^{s}w_{\varepsilon}=\bigl\|(-\Delta)^{\frac{s}{2}}U_{\varepsilon}\bigr\|_{L^{2}(\mathbb{R}^{n})}^{2}+2\int(-\Delta)^{\frac{s}{2}}z_{\varepsilon}\cdot(-\Delta)^{\frac{s}{2}}U_{\varepsilon}+\bigl\|(-\Delta)^{\frac{s}{2}}z_{\varepsilon}\bigr\|_{L^{2}(\mathbb{R}^{n})}^{2}.$ (3.13)
Step 1: Scaling invariance of the main term. Recall
$U_{\varepsilon}(x)=\varepsilon^{-\frac{n-2s}{2}}\,U\!\left(\frac{x}{\varepsilon}\right).$
We claim that
$\bigl\|(-\Delta)^{\frac{s}{2}}U_{\varepsilon}\bigr\|_{L^{2}(\mathbb{R}^{n})}^{2}=\bigl\|(-\Delta)^{\frac{s}{2}}U\bigr\|_{L^{2}(\mathbb{R}^{n})}^{2}.$ (3.14)
Indeed, using the Fourier definition
$\widehat{(-\Delta)^{\frac{s}{2}}f}(\xi)=|\xi|^{s}\,\hat{f}(\xi),\qquad\|(-\Delta)^{\frac{s}{2}}f\|_{2}^{2}=\int_{\mathbb{R}^{n}}|\xi|^{2s}\,|\hat{f}(\xi)|^{2}\,d\xi,$
it suffices to compute the scaling of $\widehat{U_{\varepsilon}}$. By definition,
$\widehat{U_{\varepsilon}}(\xi)$ $=(2\pi)^{-\frac{n}{2}}\int_{\mathbb{R}^{n}}\varepsilon^{-\frac{n-2s}{2}}\,U\!\left(\frac{x}{\varepsilon}\right)e^{-ix\cdot\xi}\,dx.$
With the change of variables $x=\varepsilon y$,
$\widehat{U_{\varepsilon}}(\xi)$ $=(2\pi)^{-\frac{n}{2}}\varepsilon^{-\frac{n-2s}{2}}\int_{\mathbb{R}^{n}}U(y)\,e^{-i(\varepsilon y)\cdot\xi}\,\varepsilon^{n}\,dy=(2\pi)^{-\frac{n}{2}}\varepsilon^{\frac{n+2s}{2}}\int_{\mathbb{R}^{n}}U(y)\,e^{-iy\cdot(\varepsilon\xi)}\,dy$
$=\varepsilon^{\frac{n+2s}{2}}\,\widehat{U}(\varepsilon\xi).$
Hence
$\bigl\|(-\Delta)^{\frac{s}{2}}U_{\varepsilon}\bigr\|_{2}^{2}$ $=\int_{\mathbb{R}^{n}}|\xi|^{2s}\,|\widehat{U_{\varepsilon}}(\xi)|^{2}\,d\xi=\int_{\mathbb{R}^{n}}|\xi|^{2s}\,\varepsilon^{n+2s}\,|\widehat{U}(\varepsilon\xi)|^{2}\,d\xi.$
Now set $\zeta=\varepsilon\xi$. Then
$\bigl\|(-\Delta)^{\frac{s}{2}}U_{\varepsilon}\bigr\|_{2}^{2}$ $=\int_{\mathbb{R}^{n}}\Bigl|\frac{\zeta}{\varepsilon}\Bigr|^{2s}\,\varepsilon^{n+2s}\,|\widehat{U}(\zeta)|^{2}\,\varepsilon^{-n}\,d\zeta=\int_{\mathbb{R}^{n}}|\zeta|^{2s}\,|\widehat{U}(\zeta)|^{2}\,d\zeta=\bigl\|(-\Delta)^{\frac{s}{2}}U\bigr\|_{2}^{2}.$
Step 2: The cross term is $O(\varepsilon^{n})$. We stress that for $s>1$, it is not convenient to interpret $(-\Delta)^{s}U$ pointwise. Instead we use the Fourier definition. Since $U$ is an extremal, it is a critical point of the Sobolev quotient
$\mathcal{J}(f):=\frac{\|(-\Delta)^{\frac{s}{2}}f\|_{2}^{2}}{\|f\|_{2_{s}^{*}}^{2}},$
hence there exists $\kappa_{n,s}>0$ such that the following Euler–Lagrange equation holds in weak form: for every $\varphi\in H^{s}(\mathbb{R}^{n})$,
$\int_{\mathbb{R}^{n}}(-\Delta)^{\frac{s}{2}}U\;(-\Delta)^{\frac{s}{2}}\varphi\,dx=\kappa_{n,s}\int_{\mathbb{R}^{n}}U^{2_{s}^{*}-1}\,\varphi\,dx.$
By scaling invariance, the same identity holds for $U_{\varepsilon}$:
$\int_{\mathbb{R}^{n}}(-\Delta)^{\frac{s}{2}}U_{\varepsilon}\;(-\Delta)^{\frac{s}{2}}\varphi\,dx=\kappa_{n,s}\int_{\mathbb{R}^{n}}U_{\varepsilon}^{2_{s}^{*}-1}\,\varphi\,dx,\quad\forall\,\varphi\in H^{s}(\mathbb{R}^{n}).$ (3.15)
Now let $z_{\varepsilon}:=w_{\varepsilon}-U_{\varepsilon}=(\eta-1)U_{\varepsilon}$. Applying (3.15) with $\varphi=z_{\varepsilon}$ gives
$\int_{\mathbb{R}^{n}}(-\Delta)^{\frac{s}{2}}U_{\varepsilon}\;(-\Delta)^{\frac{s}{2}}z_{\varepsilon}\,dx=\kappa_{n,s}\int_{\mathbb{R}^{n}}U_{\varepsilon}^{2_{s}^{*}-1}\,z_{\varepsilon}\,dx=\kappa_{n,s}\int_{\mathbb{R}^{n}}(\eta-1)\,U_{\varepsilon}^{2_{s}^{*}}\,dx.$ (3.16)
Since $\eta\equiv 1$ on $B_{\delta}(0)$, the integrand is supported in $\{|x|\geq\delta\}$, we estimate (for $\varepsilon$ small so that $\delta/\varepsilon\geq 1$)
$\left|\int_{\mathbb{R}^{n}}(\eta-1)\,U_{\varepsilon}^{2_{s}^{*}}\,dx\right|$ $\leq\int_{|x|\geq\delta}|U_{\varepsilon}(x)|^{2_{s}^{*}}\,dx=\int_{|y|\geq\delta/\varepsilon}(1+|y|^{2})^{-n}\,dy=O(\varepsilon^{n}).$
Combining this with (3.16) yields
$\left|\int_{\mathbb{R}^{n}}(-\Delta)^{\frac{s}{2}}U_{\varepsilon}\;(-\Delta)^{\frac{s}{2}}z_{\varepsilon}\,dx\right|\leq O(\varepsilon^{n}).$ (3.17)
Step 3: The remainder energy is $O(\varepsilon^{n-2s})$. Choose an integer m>sm>\!s. Since $|\xi|^{2s}\leq(1+|\xi|^{2})^{m}$, we have
$\bigl\|(-\Delta)^{\frac{s}{2}}z_{\varepsilon}\bigr\|_{L^{2}(\mathbb{R}^{n})}^{2}=\int|\xi|^{2s}|\widehat{z}_{\varepsilon}(\xi)|^{2}\,d\xi\leq\int(1+|\xi|^{2})^{m}|\widehat{z}_{\varepsilon}(\xi)|^{2}\,d\xi=\|z_{\varepsilon}\|_{H^{m}(\mathbb{R}^{n})}^{2}.$ (3.18)
On the fixed annulus $B_{2\delta}\setminus B_{\delta}$, the cutoff factors $\eta-1$ and all their derivatives are bounded. Moreover, by Lemma 3.4, for every multi-index $\alpha$ with $|\alpha|\leq m$, for $|x|\geq\delta$, one has the pointwise decay
$|\partial^{\alpha}U_{\varepsilon}(x)|\leq C_{\alpha,\delta}\,\varepsilon^{\frac{n-2s}{2}},$ (3.19)
By Leibniz’ rule,
$\partial^{\alpha}z_{\varepsilon}=\sum_{\beta\leq\alpha}\binom{\alpha}{\beta}\partial^{\alpha-\beta}(\eta-1)\,\partial^{\beta}U_{\varepsilon},$
hence (3.19) implies
$\|\partial^{\alpha}z_{\varepsilon}\|_{L^{2}(\mathbb{R}^{n})}^{2}\leq C_{\alpha,\delta}\,\varepsilon^{\,n-2s},\qquad|\alpha|\leq m.$
Summing over $|\alpha|\leq m$ yields
$\|z_{\varepsilon}\|_{H^{m}(\mathbb{R}^{n})}^{2}\leq C_{\delta,m}\,\varepsilon^{\,n-2s}.$ (3.20)
Combining (3.18) and (3.20) gives
$\bigl\|(-\Delta)^{\frac{s}{2}}z_{\varepsilon}\bigr\|_{L^{2}(\mathbb{R}^{n})}^{2}\leq O(\varepsilon^{\,n-2s}).$ (3.21)
Inserting (3.14), (3.17), and (3.21) into (3.13), and using that $n>2s$ we obtain
$\int_{\mathbb{R}^{n}}w_{\varepsilon}\,(-\Delta)^{s}w_{\varepsilon}\,dx=\bigl\|(-\Delta)^{\frac{s}{2}}U\bigr\|_{L^{2}(\mathbb{R}^{n})}^{2}\;+\;O(\varepsilon^{\,n-2s}),$
which is (3.11). ∎
3.2 Proof of Theorem 1.3, 1.4
Proof of Theorem 1.3: Let
$U(x):=(1+|x|^{2})^{-\frac{n-2s}{2}},\qquad U_{\varepsilon}(x):=\varepsilon^{-\frac{n-2s}{2}}U\Bigl(\frac{x}{\varepsilon}\Bigr),$
and fix $0<\delta<\frac{1}{4}$. Choose $\eta\in C_{c}^{\infty}(\mathbb{B}^{n})$ such that $0\leq\eta\leq 1$, $\eta\equiv 1$ on $B_{\delta}(0)$ and $\eta\equiv 0$ on $\mathbb{B}^{n}\setminus B_{2\delta}(0)$. Define
$w_{\varepsilon}:=\eta\,U_{\varepsilon}\in C_{c}^{\infty}(\mathbb{R}^{n}),\qquad u_{\varepsilon}:=\phi^{\frac{n}{2}-s}w_{\varepsilon}\in C_{c}^{\infty}(\mathbb{B}^{n}).$
By Lemma 3.1,
$\int_{\mathbb{B}^{n}}(\widetilde{\mathcal{P}}_{s}u_{\varepsilon})u_{\varepsilon}\,dV_{{\mathbb{B}}^{n}}=\int_{\mathbb{R}^{n}}w_{\varepsilon}(-\Delta)^{s}w_{\varepsilon}\,dx,\quad\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2_{s}^{*}}\,dV_{{\mathbb{B}}^{n}}=\int_{\mathbb{R}^{n}}|w_{\varepsilon}|^{2_{s}^{*}}\,dx.$ (3.22)
In particular, setting
$B_{\varepsilon}:=\Bigl(\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2_{s}^{*}}\,dV_{{\mathbb{B}}^{n}}\Bigr)^{\!\frac{2}{2_{s}^{*}}}=\Bigl(\int_{\mathbb{R}^{n}}|w_{\varepsilon}|^{2_{s}^{*}}\,dx\Bigr)^{\!\frac{2}{2_{s}^{*}}},$
by Proposition 3.2, we have $B_{\varepsilon}=B_{0}+O(\varepsilon^{n})$ with
$B_{0}:=\Bigl(\int_{\mathbb{R}^{n}}|U|^{2_{s}^{*}}\,dx\Bigr)^{\!\frac{2}{2_{s}^{*}}}>0,$
and by (3.11) and (3.22),
$\frac{\displaystyle\int_{\mathbb{B}^{n}}(\widetilde{\mathcal{P}}_{s}u_{\varepsilon})u_{\varepsilon}\,dV_{{\mathbb{B}}^{n}}}{B_{\varepsilon}}\;=\;S_{n,s}\;+\;O(\varepsilon^{\,n-2s}).$ (3.23)
By the condition that
$S_{n,s}\;\leq\;\frac{\displaystyle\int_{\mathbb{B}^{n}}(\widetilde{\mathcal{P}}_{s}u_{\varepsilon})u_{\varepsilon}\,dV_{{\mathbb{B}}^{n}}}{B_{\varepsilon}}\;+\;\lambda\,\frac{\displaystyle\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2}\,dV_{{\mathbb{B}}^{n}}}{B_{\varepsilon}}.$ (3.24)
By (3.23), the first fraction is $S_{n,s}+O(\varepsilon^{n-2s})$. The second fraction has the same order as $\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2}\,dV_{{\mathbb{B}}^{n}}$.
Case $n>4s$. Then (3.2) gives
$\frac{\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2}\,dV_{{\mathbb{B}}^{n}}}{B_{\varepsilon}}=\widetilde{c}_{n,s}\,\varepsilon^{2s}+O(\varepsilon^{n-2s})\qquad\text{with }\widetilde{c}_{n,s}>0,$
whereas $O(\varepsilon^{n-2s})=o(\varepsilon^{2s})$ because n2s>2sn-2s>2s. Thus (3.24) becomes
$S_{n,s}\leq S_{n,s}+\lambda\,\widetilde{c}_{n,s}\,\varepsilon^{2s}+o(\varepsilon^{2s}).$
If $\lambda<0$, the right-hand side is $S_{n,s}$ for $\varepsilon$ small, a contradiction. Hence $\lambda\geq 0$.
Case $n=4s$. Then
$\frac{\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2}\,dV_{{\mathbb{B}}^{n}}}{B_{\varepsilon}}=\widetilde{c}_{n,s}\,\varepsilon^{2s}|\log\varepsilon|+O(\varepsilon^{2s}),\qquad\widetilde{c}_{n,s}>0,$
while the energy error is $O(\varepsilon^{n-2s})=O(\varepsilon^{2s})$. Therefore the logarithmic factor dominates, and the same contradiction argument shows $\lambda\geq 0$.
(iii) Case $2s<n<4s$. In this regime, both the energy term and the $L^{2}$-term are of order $\varepsilon^{n-2s}$. More precisely, one has
$\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2}\,dV_{\mathbb{B}^{n}}=c_{n,s}\,\varepsilon^{n-2s}+O(\varepsilon^{2s}),\qquad c_{n,s}>0.$
Plugging these expansions into (3.24), we obtain
$S_{n,s}\leq S_{n,s}+O(\varepsilon^{n-2s})+\lambda\,c_{n,s}\,\varepsilon^{n-2s}+O(\varepsilon^{2s}).$
Combining this estimate with [24, Theorem 1.8], we can only conclude that
$\lambda\geq-\widetilde{\lambda}_{0,s}^{\mathrm{conf}}.$
Equivalently, there exists a constant $\widetilde{\lambda}_{s}^{\mathrm{conf}}\in(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}]$ such that
$\widetilde{\Lambda}_{n,s}^{\mathrm{HS}}\;\geq\;-\widetilde{\lambda}_{s}^{\mathrm{conf}}.$
This completes the proof. $\Box$
Proof of Theorem 1.4:
(i) Applying [24, Lemma 5.2] with $\gamma=s$ and $\lambda=0$, we obtain
$2^{2s}\,\frac{\Gamma\bigl(\frac{3+2s}{4}\bigr)^{2}}{\Gamma\bigl(\frac{3-2s}{4}\bigr)^{2}}=\frac{\Gamma\bigl(s+\frac{1}{2}\bigr)^{2}}{\Gamma\bigl(\frac{1}{2}\bigr)^{2}}+\frac{\sin(\pi s)}{\pi}\,\Gamma\bigl(s+\tfrac{1}{2}\bigr)^{2}.$ (3.25)
Since $\Gamma(\frac{1}{2})=\sqrt{\pi}$, identity (3.25) becomes
$\lambda_{0,s}^{\mathrm{conf}}=\frac{1+\sin(\pi s)}{\pi}\,\Gamma\bigl(s+\tfrac{1}{2}\bigr)^{2}\geq 0.$ (3.26)
We now compare $\lambda_{0,s}^{\mathrm{conf}}$ with $b_{s}$.
Case 1: $\sin(\pi s)\leq 0$. Then $b_{s}=0$, hence $b_{s}\leq\lambda_{0,s}^{\mathrm{conf}}$. Moreover, (3.26) yields the second line of (1.21).
Case 2: $\sin(\pi s)>0$. Then $b_{s}=\frac{\sin(\pi s)}{\pi}\Gamma\bigl(s+\tfrac{1}{2}\bigr)^{2}$. Subtracting from (3.26) gives
$\lambda_{0,s}^{\mathrm{conf}}-b_{s}=\Bigl(\frac{1+\sin(\pi s)}{\pi}-\frac{\sin(\pi s)}{\pi}\Bigr)\Gamma\bigl(s+\tfrac{1}{2}\bigr)^{2}=\frac{\Gamma\bigl(s+\tfrac{1}{2}\bigr)^{2}}{\pi}>0,$
which proves both $b_{s}<\lambda_{0,s}^{\mathrm{conf}}$ and (1.21).
(ii) We define the spectral multiplier
$M_{s}(\beta):=\Bigl|\Gamma\Bigl(s+\tfrac{1}{2}+i\beta\Bigr)\Bigr|^{2},\qquad\beta\geq 0.$
Using the classical asymptotic formula
$\lim_{|\lambda|\to\infty}\,\bigl|\Gamma(a+i\lambda)\bigr|\,e^{\frac{\pi}{2}|\lambda|}\,|\lambda|^{\frac{1}{2}-a}=\sqrt{2\pi},$
we know that $M_{s}(\beta)\to 0$ as $\beta\to\infty$. By the fact that $|\Gamma(a+i\lambda)|\leq|\Gamma(a)|$,
$\sup_{\beta\geq 0}M_{s}(\beta)=M_{s}(0)\in(0,\infty).$ (3.27)
Thus, by the spectral calculus,
$\|\,M_{s}(A)\,\|_{L^{2}\to L^{2}}\leq M_{s}(0).$
By identity (1.6),
$\mathcal{P}_{s}=\widetilde{\mathcal{P}}_{s}+\frac{\sin(s\pi)}{\pi}\,M_{s}(A),\ s\in\left(0,\frac{n}{2}\right)\setminus\mathbb{N}$ (3.28)
where $M_{s}(A)$ is defined by functional calculus and is a bounded, self-adjoint, nonnegative operator.
Case 1: $\sin(s\pi)\leq 0$. Then the last term in (3.28) is nonpositive, hence
$\langle\widetilde{\mathcal{P}}_{s}u,u\rangle_{L^{2}}\;\geq\;\langle\mathcal{P}_{s}u,u\rangle_{L^{2}}\qquad\text{for all }u\in C_{c}^{\infty}(\mathbb{B}^{n}).$
Now apply Theorem 1.3 to conclude the corresponding lower bound for $\lambda$.
Case 2: $\sin(s\pi)>0$. Since $M_{s}(A)\leq M_{s}(0)\,\mathrm{Id}$ on $L^{2}$, we have from (3.28):
$\langle\mathcal{P}_{s}u,u\rangle_{L^{2}}\leq\langle\widetilde{\mathcal{P}}_{s}u,u\rangle_{L^{2}}+\frac{\sin(s\pi)}{\pi}\,M_{s}(0)\,\|u\|_{L^{2}}^{2}=\langle\widetilde{\mathcal{P}}_{s}u,u\rangle_{L^{2}}+b_{s}\,\|u\|_{L^{2}}^{2}.$
Equivalently,
$\langle\widetilde{\mathcal{P}}_{s}u,u\rangle_{L^{2}}\geq\langle\mathcal{P}_{s}u,u\rangle_{L^{2}}-b_{s}\,\|u\|_{L^{2}}^{2}.$
Thus Theorem 1.3 applies with the shifted coefficient $\lambda^{\prime}:=\lambda+b_{s}$. Reading off the conclusion gives exactly (b), i.e.
$\lambda\geq-b_{s}\quad\text{when }n\geq 4s,\qquad\lambda\geq-b_{s}-\widetilde{\lambda}_{s}^{\mathrm{conf}}\quad\text{when }2s<n<4s.$
Combining this with Lemma 5.4, we obtain the desired conclusion. $\Box$
4  Fractional Brezis–Nirenberg Problems
4.1 Energy Spaces
Now we can study fractional Brezis–Nirenberg type equations (1.8) and (1.9) on the whole hyperbolic space $\mathbb{H}^{n}$ driven by $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$, namely
[…]
and for exponents
$1<p\leq 2_{s}^{*}-1,\qquad 2_{s}^{*}:=\frac{2n}{n-2s}.$
Since the bottom spectrum is
$\lambda_{0,s}^{\mathrm{conf}}:=\inf\sigma(P_{s})=m_{s}(0)\geq 0,\quad\widetilde{\lambda}_{0,s}^{\mathrm{conf}}:=\inf\sigma(\widetilde{P}_{s})=\widetilde{m}_{s}(0)>0$
Consequently, by [19, Theorem E.8], we obtain the sharp fractional Poincaré inequality
$\lambda_{0,s}^{\mathrm{conf}}\,\|u\|_{L^{2}(\mathbb{H}^{n})}^{2}\leq\langle u,\mathcal{P}_{s}u\rangle_{L^{2}(\mathbb{H}^{n})},\ \ \widetilde{\lambda}_{0,s}^{\mathrm{conf}}\|u\|_{L^{2}(\mathbb{H}^{n})}^{2}\leq\langle u,\widetilde{\mathcal{P}}_{s}u\rangle_{L^{2}(\mathbb{H}^{n})}\qquad\text{for all }u\in C_{c}^{\infty}(\mathbb{H}^{n}),$ (4.1)
and hence quadratic forms
$\mathcal{E}_{\lambda,s}(u):=\langle(\mathcal{P}_{s}-\lambda)u,u\rangle_{L^{2}(\mathbb{H}^{n})}=\langle\mathcal{P}_{s}u,u\rangle_{L^{2}(\mathbb{H}^{n})}-\lambda\|u\|_{2}^{2}$
and
${\widetilde{\mathcal{E}}}_{\lambda,s}(u):=\langle(\widetilde{\mathcal{P}}_{s}-\lambda)u,u\rangle_{L^{2}(\mathbb{H}^{n})}=\langle\widetilde{\mathcal{P}}_{s}u,u\rangle_{L^{2}(\mathbb{H}^{n})}-\lambda\|u\|_{2}^{2}$
are nonnegative whenever $\lambda\leq\lambda_{0,s}^{\mathrm{conf}}$ and $\lambda\leq\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$, respectively.
For $u,v\in C_{c}^{\infty}(\mathbb{H}^{n})$, we set
$\langle u,v\rangle_{\lambda,s}:=\langle(\mathcal{P}_{s}-\lambda)u,v\rangle_{L^{2}(\mathbb{H}^{n})}=\int_{\mathbb{H}^{n}}v\ (\mathcal{P}_{s}-\lambda)u\,dV_{\mathbb{H}^{n}},\qquad\|u\|_{\lambda,s}^{2}:=\langle u,u\rangle_{\lambda,s}=\mathcal{E}_{\lambda,s}(u),$
and
$\langle u,v\rangle_{\lambda,s,\sim}:=\langle(\widetilde{\mathcal{P}}_{s}-\lambda)u,v\rangle_{L^{2}(\mathbb{H}^{n})}=\int_{\mathbb{H}^{n}}v\ (\widetilde{\mathcal{P}}_{s}-\lambda)u,dV_{\mathbb{H}^{n}},\qquad\|u\|_{\lambda,s,\sim}^{2}:=\langle u,u\rangle_{\lambda,s,\sim}={\widetilde{\mathcal{E}}}_{\lambda,s}(u).$
By [24, Theorem 1.3 and Theorem 1.8], for every $\lambda\leq\lambda_{0,s}^{\mathrm{conf}}$, the map $u\mapsto\|u\|_{\lambda,s}$ defines a norm on $C_{c}^{\infty}(\mathbb{H}^{n})$, and for every $\lambda\leq\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$, the map $u\mapsto\|u\|_{\lambda,s,\sim}$ also defines a norm on $C_{c}^{\infty}(\mathbb{H}^{n})$
Lemma 4.1.
Let $s\in(0,\frac{n}{2})$. $\widetilde{m}_{s}(\beta)$ is defined in (2.8). Then there exist constants $c_{1},c_{2}>0$ (depending only on $s$) such that for all $\beta\in\mathbb{R}$,
$c_{1}\,(\beta^{2}+1)^{s}\ \leq\ \widetilde{m}_{s}(\beta)\ \leq\ c_{2}\,(\beta^{2}+1)^{s}.$
Proof: First note that $\widetilde{m}_{s}$ is even and strictly positive on $\mathbb{R}$, since $\Gamma$ has no zeros and $\Gamma(\tfrac{1}{2}+i\beta)\neq 0$ for all $\beta\in\mathbb{R}$. We use the uniform Stirling estimate in vertical strips: for each fixed $\sigma>0$, there exist $T\geq 1$ and constants $C_{\pm}=C_{\pm}(\sigma)>0$ such that for all $\sigma^{\prime}\in[\sigma,\sigma+1]$ and all $|t|\geq T$,
$C_{-}\,|t|^{\sigma^{\prime}-\frac{1}{2}}e^{-\frac{\pi}{2}|t|}\ \leq\ |\Gamma(\sigma^{\prime}+it)|\ \leq\ C_{+}\,|t|^{\sigma^{\prime}-\frac{1}{2}}e^{-\frac{\pi}{2}|t|}.$ (4.2)
Apply (4.2) with $\sigma^{\prime}=s+\frac{1}{2}$ and $\sigma^{\prime}=\frac{1}{2}$. For $|\beta|\geq T$ this gives
$\frac{\bigl|\Gamma(s+\frac{1}{2}+i\beta)\bigr|}{\bigl|\Gamma(\frac{1}{2}+i\beta)\bigr|}\sim|\beta|^{s},$
hence
$\widetilde{m}_{s}(\beta)=\Bigl(\frac{\bigl|\Gamma(s+\frac{1}{2}+i\beta)\bigr|}{\bigl|\Gamma(\frac{1}{2}+i\beta)\bigr|}\Bigr)^{\!2}\sim|\beta|^{2s},\qquad|\beta|\geq T.$
Since $(1+\beta^{2})^{s}\sim|\beta|^{2s}$ for $|\beta|\gg 1$, we obtain constants $C_{1},C_{2}>0$ such that
$C_{1}(1+\beta^{2})^{s}\leq\widetilde{m}_{s}(\beta)\leq C_{2}(1+\beta^{2})^{s},\qquad|\beta|\geq T.$ (4.3)
On the compact set $\{|\beta|\leq T\}$, the function $\widetilde{m}_{s}$ is continuous and strictly positive, and $(1+\beta^{2})^{s}$ is also bounded above and below by positive constants on $\{|\beta|\leq T\}$, combining with (4.3) yields global constants (still denoted $C_{1},C_{2}$) such that
$C_{1}(1+\beta^{2})^{s}\leq\widetilde{m}_{s}(\beta)\leq C_{2}(1+\beta^{2})^{s},\qquad\forall\,\beta\in\mathbb{R},$
which is the desired estimate. ∎
By Lemma 4.1 we have
$\mathcal{D}(\widetilde{\mathcal{P}}_{s})=H^{2s}(\mathbb{H}^{n}).$
where $H^{2s}(\mathbb{H}^{n})$ is given in (2.7). Consequently, for every $\lambda<\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$, the quadratic form
$\|u\|_{\lambda,s,\sim}^{2}:=\bigl\langle(\widetilde{\mathcal{P}}_{s}-\lambda)u,u\bigr\rangle_{L^{2}(\mathbb{H}^{n})}$
defines a norm on $H^{s}(\mathbb{H}^{n})$ which is equivalent to the standard $H^{s}$-norm.
Similarly, using the high–frequency asymptotics of $m_{s}$ (via Stirling’s formula) together with the fact that $m_{s}$ is bounded on compact $\beta$-intervals, one obtains the global two–sided comparison
$1+m_{s}(\beta)^{2}\ \sim\ 1+(\beta^{2}+1)^{2s},\qquad 1+\beta\in\mathbb{R},$ (4.4)
with implicit constants depending only on $s$ and $\rho$. In particular,
$\mathcal{D}(\mathcal{P}_{s})=H^{2s}(\mathbb{H}^{n}).$
We emphasize that, in general, one cannot upgrade (4.4) to $m_{s}(\beta)\sim(\beta^{2}+\rho^{2})^{s}$ for all $\beta\in\mathbb{R}$, for instance, this fails in the exceptional cases $s=\frac{3}{2}+2k$, where $m_{s}(\beta)\sim\beta^{2}$ as $\beta\to 0$.
Lemma 4.2.
Let $s\in(0,\frac{n}{2})\setminus\mathbb{N},$ for every $\lambda<\lambda_{0,s}^{\mathrm{conf}}$, the quadratic form
$\|u\|_{\lambda,s}^{2}:=\langle(\mathcal{P}_{s}-\lambda)u,u\rangle_{L^{2}(\mathbb{H}^{n})},\quad u\in C_{c}^{\infty}(\mathbb{H}^{n})$
extends to a norm on $H^{s}(\mathbb{H}^{n})$, which is equivalent to the standard $H^{s}$-norm.
Proof: By Plancherel formula,
$\|u\|_{\lambda,s}^{2}=\langle(\mathcal{P}_{s}-\lambda)u,u\rangle_{2}=\int_{0}^{\infty}\!\!\int_{\mathbb{S}^{n-1}}\bigl(m_{s}(\beta)-\lambda\bigr)\,|\hat{u}(\beta,\theta)|^{2}\,\frac{d\sigma(\theta)\,d\beta}{|c(\beta)|^{2}}.$ (4.5)
Set $\lambda_{0}:=\lambda_{0,s}^{\mathrm{conf}}$ and $\delta:=\lambda_{0}-\lambda>0$. Since $m_{s}(\beta)\geq\lambda_{0}$ for all $\beta$, we have $m_{s}(\beta)-\lambda\geq\delta$. Moreover,
$m_{s}(\beta)-\lambda\leq m_{s}(\beta)+|\lambda|\leq(1+|\lambda|)\,(1+m_{s}(\beta))\qquad\forall\,\beta\in\mathbb{R}.$
To bound $1+m_{s}(\beta)$ from above by $m_{s}(\beta)-\lambda$, set
$k:=\min\left\{\frac{\lambda_{0}-\lambda}{1+\lambda_{0}},\frac{1}{2}\right\}>0.$
Since $m_{s}(\beta)\geq\lambda_{0}$ for all $\beta$, we have
$m_{s}(\beta)-\lambda-k\bigl(1+m_{s}(\beta)\bigr)\geq 0\Leftrightarrow\lambda_{0}\geq\frac{\lambda+k}{1-k}\Leftrightarrow k\leq\frac{\lambda_{0}-\lambda}{1+\lambda_{0}},$
and hence
$m_{s}(\beta)-\lambda\geq k\bigl(1+m_{s}(\beta)\bigr),\qquad\forall\,\beta\in\mathbb{R}.$
Therefore there exists $C_{\lambda}>0$ such that
$C_{\lambda}^{-1}\,(1+m_{s}(\beta))\leq m_{s}(\beta)-\lambda\leq C_{\lambda}\,(1+m_{s}(\beta)),\qquad\forall\,\beta\in\mathbb{R}.$
Plugging the pointwise comparison into (4.5) and using the definition (2.7) of $H^{s}(\mathbb{H}^{n})$ gives
$\|u\|_{\lambda,s}^{2}\ \sim\ \|u\|_{H^{s}(\mathbb{H}^{n})}^{2},\qquad u\in C_{c}^{\infty}(\mathbb{H}^{n}).$
In particular, $\|\cdot\|_{\lambda,s}$ is positive definite on $C_{c}^{\infty}$. Since $C_{c}^{\infty}(\mathbb{H}^{n})$ is dense in $H^{s}(\mathbb{H}^{n})$, the norm extends uniquely to $H^{s}(\mathbb{H}^{n})$ and remains equivalent to the standard $H^{s}$-norm. ∎
Hence, throughout the paper, we will simply write $\|\cdot\|_{\lambda,s}$ to denote either of the equivalent norms associated with $\mathcal{P}_{s}$ or $\widetilde{\mathcal{P}}_{s}$, whenever the distinction is immaterial.
Definition 4.3.
We define the energy space as the completion of $C_{c}^{\infty}(\mathbb{H}^{n})$ under the norm $\|\cdot\|_{\lambda,s}$, namely
$\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n}):=\overline{C_{c}^{\infty}(\mathbb{H}^{n})}^{\ \|\cdot\|_{\lambda,s}},\quad\lambda<\lambda_{0,s}^{\mathrm{conf}}.$
We study the attainability of fractional Sobolev levels associated with $P_{s}-\lambda$ and $\widetilde{P}_{s}-\lambda$
$H_{n,s,p}(\lambda):=\inf_{u\in\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})\setminus\{0\}}\frac{\mathcal{E}_{\lambda,s}(u)}{\|u\|_{L^{p+1}(\mathbb{H}^{n})}^{2}}\in(0,\infty),\quad\lambda<\lambda_{0,s}^{\mathrm{conf}},$ (4.6)
and
$\widetilde{H}_{n,s,p}(\lambda):=\inf_{u\in\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})\setminus\{0\}}\frac{{\widetilde{\mathcal{E}}}_{\lambda,s}(u)}{\|u\|_{L^{p+1}(\mathbb{H}^{n})}^{2}}\in(0,\infty),\quad\lambda<\widetilde{\lambda}_{0,s}^{\mathrm{conf}}.$
For convenience, we set
$I_{\lambda,s}(u):=\frac{\mathcal{E}_{\lambda,s}(u)}{\|u\|_{L^{p+1}(\mathbb{H}^{n})}^{2}},\quad\widetilde{I}_{\lambda,s}(u):=\frac{{\widetilde{\mathcal{E}}}_{\lambda,s}(u)}{\|u\|_{L^{p+1}(\mathbb{H}^{n})}^{2}}\qquad u\in C_{c}^{\infty}(\mathbb{H}^{n})\setminus\{0\}.$
Throughout, we write $\|u\|_{p+1}:=\|u\|_{L^{p+1}(\mathbb{H}^{n})}$ for brevity. It is standard to work on the Nehari manifold
$\mathcal{N}_{\lambda,s}:=\Bigl\{u\in C_{c}^{\infty}(\mathbb{H}^{n})\setminus\{0\}:\mathcal{E}_{\lambda,s}(u)=\|u\|_{p+1}^{p+1}\Bigr\},$
and
$\widetilde{\mathcal{N}}_{\lambda,s}:=\Bigl\{u\in C_{c}^{\infty}(\mathbb{H}^{n})\setminus\{0\}:\widetilde{\mathcal{E}}_{\lambda,s}(u)=\|u\|_{p+1}^{p+1}\Bigr\}.$
Note that for every $u\neq 0$, there exists a unique $t(u)>0$ such that $t(u)u\in\mathcal{N}_{\lambda,s}$, namely
$t(u)^{p-1}=\mathcal{E}_{\lambda,s}(u)/\|u\|_{p+1}^{p+1}.$
Moreover, for $u\in\mathcal{N}_{\lambda,s}$ we have
$I_{\lambda,s}(u)=\|u\|_{p+1}^{p-1}=\mathcal{E}_{\lambda,s}(u)^{\frac{p-1}{p+1}}.$ (4.7)
Hence
$H_{n,s,p}(\lambda)=\inf_{u\in\mathcal{N}_{\lambda,s}}I_{\lambda,s}(u)=\inf_{u\in\mathcal{N}_{\lambda,s}}\|u\|_{p+1}^{p-1}.$
Lemma 4.4.
Let $\mathcal{T}$ is an isometry on $\mathbb{H}^{n}$ and define
$(U_{\mathcal{T}}u)(x):=u(\mathcal{T}x).$
Then, for all $u\in C_{c}^{\infty}(\mathbb{H}^{n})$,
$\mathcal{P}_{s}(U_{\mathcal{T}}u)=U_{\mathcal{T}}(\mathcal{P}_{s}u),\quad\widetilde{\mathcal{P}}_{s}(U_{\mathcal{T}}u)=U_{\mathcal{T}}(\widetilde{\mathcal{P}}_{s}u)$
Consequently,
$\langle U_{\mathcal{T}}u,\mathcal{P}_{s}(U_{\mathcal{T}}u)\rangle_{2}=\langle u,\mathcal{P}_{s}u\rangle_{2},\quad\langle U_{\mathcal{T}}u,\widetilde{\mathcal{P}}_{s}(U_{\mathcal{T}}u)\rangle_{2}=\langle u,\widetilde{\mathcal{P}}_{s}u\rangle_{2}$
Proof: Since $\mathcal{T}$ is an isometry, and $dV_{\mathbb{H}^{n}}$ is invariant under $\mathcal{T}$, hence $U_{\mathcal{T}}$ is unitary on $L^{2}(\mathbb{H}^{n})$ and preserves $L^{p}$ norms by change of variables. Moreover, the Laplace–Beltrami operator is invariant under isometries, i.e. $\Delta_{\mathbb{H}^{n}}(u\circ\mathcal{T})=(\Delta_{\mathbb{H}^{n}}u)\circ\mathcal{T}$. Therefore $U_{\mathcal{T}}^{-1}(-\Delta_{\mathbb{H}^{n}})U_{\mathcal{T}}=-\Delta_{\mathbb{H}^{n}}$. Since on $\mathbb{H}^{n}$, $P_{s}$ and $\widetilde{P}_{s}$ are given by spectral calculus as a Borel function of
$A:=-\Delta_{\mathbb{H}^{n}}-\rho^{2},$
the functional calculus implies $U_{\mathcal{T}}^{-1}\mathcal{P}_{s}U_{\mathcal{T}}=\mathcal{P}_{s}$ and $U_{\mathcal{T}}^{-1}\widetilde{\mathcal{P}}_{s}U_{\mathcal{T}}=\widetilde{\mathcal{P}}_{s}$, which is equivalent to
$\mathcal{P}_{s}(u\circ\mathcal{T})=(\mathcal{P}_{s}u)\circ\mathcal{T},\quad\widetilde{\mathcal{P}}_{s}(u\circ\mathcal{T})=(\widetilde{\mathcal{P}}_{s}u)\circ\mathcal{T}$
Finally, the energy identity follows from unitarity:
$\langle U_{\mathcal{T}}u,\mathcal{P}_{s}(U_{\mathcal{T}}u)\rangle_{2}=\langle U_{\mathcal{T}}u,U_{\mathcal{T}}(P_{s}u)\rangle_{2}=\langle u,\mathcal{P}_{s}u\rangle_{2}$
and
$\langle U_{\mathcal{T}}u,\widetilde{\mathcal{P}}_{s}(U_{\mathcal{T}}u)\rangle_{2}=\langle U_{\mathcal{T}}u,U_{\mathcal{T}}(\widetilde{P}_{s}u)\rangle_{2}=\langle u,\widetilde{\mathcal{P}}_{s}u\rangle_{2}.$
4.2 Pseudodifferential Operators and Energy Asymptotic
Fix $s\in\mathbb{R}$ and $\lambda>4\rho^{2}$ sufficiently large. Define the fractional Bessel potential space by
$\tilde{H}^{s}(\mathbb{H}^{n}):=(\lambda I-\Delta_{\mathbb{H}^{n}})^{-s/2}L^{2}(\mathbb{H}^{n}),\qquad\|u\|_{\tilde{H}^{s}(\mathbb{H}^{n})}:=\bigl\|(\lambda I-\Delta_{\mathbb{H}^{n}})^{s/2}u\bigr\|_{L^{2}(\mathbb{H}^{n})}.$
Endow $\tilde{H}^{s}(\mathbb{H}^{n})$ with the inner product
$\langle u,v\rangle_{\tilde{H}^{s}}:=\Big\langle(\lambda I-\Delta_{\mathbb{H}^{n}})^{s/2}u,\,(\lambda I-\Delta_{\mathbb{H}^{n}})^{s/2}v\Big\rangle_{L^{2}(\mathbb{H}^{n})}.$
Equivalently, by Plancherel and the spectral resolution of $\Delta_{\mathbb{H}^{n}}$,
$\langle u,v\rangle_{\tilde{H}^{s}}=\int_{\mathbb{R}}\!\int_{\mathbb{S}^{n-1}}(\lambda+\beta^{2}+\rho^{2})^{s}\,\hat{u}(\beta,\theta)\,\overline{\hat{v}(\beta,\theta)}\,\frac{d\theta\,d\beta}{|c(\beta)|^{2}}.$
Thus, by the definition of (2.7),
$\tilde{H}^{s}(\mathbb{H}^{n})={H}^{s}(\mathbb{H}^{n}),\quad s\geq 0.$
We recall that $P_{s}$ with $s\in(0,\frac{n}{2})\setminus{\mathbb{N}}$ on $\mathbb{H}^{n}$ defined by spectral calculus as
$\mathcal{P}_{s}=m_{s}\!\bigl(\sqrt{-\mathcal{L}}\bigr),\qquad\mathcal{L}:=\Delta_{\mathbb{H}^{n}}+\rho^{2},$ (4.8)
where $\operatorname{Spec}(-\mathcal{L})=[0,\infty)$ and, in the Helgason–Fourier representation,
$m_{s}(\beta)=2^{2s}\,\frac{\bigl|\Gamma\bigl(\frac{3+2s}{4}+\frac{{\bf i}}{2}\beta\bigr)\bigr|^{2}}{\bigl|\Gamma\bigl(\frac{3-2s}{4}+\frac{{\bf i}}{2}\beta\bigr)\bigr|^{2}},\qquad\beta\in\mathbb{R}.$ (4.9)
Lemma 4.5.
Let $\gamma>0$ and set
$a:=\frac{3+2\gamma}{4}>0,\qquad b:=\frac{3-2\gamma}{4}.$
Define, for $\zeta\in\mathbb{C}$,
$m_{\gamma}(\zeta):=2^{2\gamma}\,\frac{\Gamma\!\bigl(a+\frac{i}{2}\zeta\bigr)\,\Gamma\!\bigl(a-\frac{i}{2}\zeta\bigr)}{\Gamma\!\bigl(b+\frac{i}{2}\zeta\bigr)\,\Gamma\!\bigl(b-\frac{i}{2}\zeta\bigr)}.$ (4.10)
Then $m_{\gamma}$ is holomorphic in every strip $|\Im\zeta|<W$ with $W<\epsilon$, where one may take
$\epsilon:=2a=\frac{3+2\gamma}{2}.$
Moreover, for each $W<\epsilon$ and each $k\in\mathbb{N}_{0}$ there exists a constant $C_{k,\gamma,W}>0$ such that
$|\partial_{\zeta}^{k}m_{\gamma}(\zeta)|\leq C_{k,\gamma,W}\,(1+|\zeta|)^{2\gamma-k},\qquad|\Im\zeta|<W.$ (4.11)
Proof: The only possible singularities of (4.10) come from the poles of $\Gamma\!\bigl(a\pm\frac{i}{2}\zeta\bigr)$, which occur precisely when $a\pm\frac{i}{2}\zeta\in\{0,-1,-2,\dots\}$, i.e.
$\zeta=\pm 2i(a+k),\qquad k\in\mathbb{N}_{0}.$
Hence $m_{\gamma}$ is holomorphic in the strip $|\Im\zeta|<2a=:\epsilon$. Poles of $\Gamma\!\bigl(b\pm\frac{i}{2}\zeta\bigr)$, if any, produce zeros of $m_{\gamma}$ and therefore do not affect holomorphy.
Fix $W<\epsilon$ and write
$m_{\gamma}(\zeta)=2^{2\gamma}\,R_{+}(\zeta)\,R_{-}(\zeta),\qquad R_{\pm}(\zeta):=\frac{\Gamma\!\bigl(a\pm\frac{i}{2}\zeta\bigr)}{\Gamma\!\bigl(b\pm\frac{i}{2}\zeta\bigr)}.$
Note that $a=b+\gamma$. For $|\Im\zeta|<W$ and $\pm$ fixed, set
$z_{\pm}:=b\pm\frac{i}{2}\zeta,\qquad\text{so that}\qquad R_{\pm}(\zeta)=\frac{\Gamma(z_{\pm}+\gamma)}{\Gamma(z_{\pm})}.$
Uniform Stirling estimates in vertical strips yield
$|R_{\pm}(\zeta)|\leq C_{\gamma,W}\,(1+|\zeta|)^{\gamma},\qquad|\Im\zeta|<W,$
and therefore
$|m_{\gamma}(\zeta)|\leq C_{\gamma,W}\,(1+|\zeta|)^{2\gamma},\qquad|\Im\zeta|<W.$
We next estimate derivatives for $|\zeta|\to\infty$ inside the strip. Choose $R_{0}\geq 1$ so large that for all $\zeta$ with $|\Im\zeta|<W$ and $|\zeta|\geq R_{0}$, the points $z_{\pm}$ stay at a positive distance from the poles $\{0,-1,-2,\dots\}$ of the digamma and polygamma functions. On the region
$\Omega_{R_{0},W}:=\{\zeta\in\mathbb{C}:\ |\Im\zeta|<W,\ |\zeta|\geq R_{0}\},$
we may choose a holomorphic branch of $\log R_{\pm}$ and use $\frac{d}{dz}\log\Gamma(z)=\psi(z)$ to compute
$\partial_{\zeta}\log R_{\pm}(\zeta)=\pm\frac{i}{2}\Bigl(\psi(z_{\pm}+\gamma)-\psi(z_{\pm})\Bigr).$
By the uniform asymptotic expansions in vertical strips,
$\psi(z)=\log z+O\!\left(\frac{1}{z}\right),\qquad\psi^{(j)}(z)=O\!\left(\frac{1}{|z|^{j}}\right)\quad(j\geq 1),\qquad|z|\to\infty,$
and $\log(z+\gamma)-\log z=O(1/z)$, we obtain
$\psi(z_{\pm}+\gamma)-\psi(z_{\pm})=O\!\left(\frac{1}{|z_{\pm}|}\right)=O\!\left(\frac{1}{|\zeta|}\right),\qquad|\zeta|\to\infty,\quad\zeta\in\Omega_{R_{0},W}.$
Moreover, for every $j\geq 1$, the polygamma bounds yield
$\partial_{\zeta}^{\,j}\Bigl(\psi(z_{\pm}+\gamma)-\psi(z_{\pm})\Bigr)=O\!\left(\frac{1}{|\zeta|^{j+1}}\right),\qquad|\zeta|\to\infty,\quad\zeta\in\Omega_{R_{0},W}.$
Consequently, for all $j\in\mathbb{N}_{0}$,
$\Bigl|\partial_{\zeta}^{\,j}\bigl(\partial_{\zeta}\log R_{\pm}(\zeta)\bigr)\Bigr|\leq C_{j,\gamma,W}\,(1+|\zeta|)^{-j-1},\qquad\zeta\in\Omega_{R_{0},W}.$
Since $R_{\pm}=\exp(\log R_{\pm})$, Faà di Bruno’s formula (Bell polynomials) implies that for every $k\in\mathbb{N}_{0}$,
$|\partial_{\zeta}^{\,k}R_{\pm}(\zeta)|\leq C_{k,\gamma,W}\,|R_{\pm}(\zeta)|\,(1+|\zeta|)^{-k}\leq C_{k,\gamma,W}\,(1+|\zeta|)^{\gamma-k},\qquad\zeta\in\Omega_{R_{0},W}.$
Consider the compact set
$K_{R_{0},W}:=\{\zeta\in\mathbb{C}:\ |\Im\zeta|\leq W,\ |\zeta|\leq R_{0}\}.$
Since $m_{\gamma}$ (hence $R_{\pm}$) is holomorphic in a neighbourhood of $K_{R_{0},W}$, each derivative $\partial_{\zeta}^{\,k}R_{\pm}$ is continuous and bounded on $K_{R_{0},W}$. Thus there exists $C_{k,\gamma,W}>0$ such that
$|\partial_{\zeta}^{\,k}R_{\pm}(\zeta)|\leq C_{k,\gamma,W}\,(1+|\zeta|)^{\gamma-k},\qquad\zeta\in K_{R_{0},W}.$
Combining this with the estimate on $\Omega_{R_{0},W}$ yields, for all $\zeta$ with $|\Im\zeta|<W$,
$|\partial_{\zeta}^{\,k}R_{\pm}(\zeta)|\leq C_{k,\gamma,W}\,(1+|\zeta|)^{\gamma-k}.$
Finally, by Leibniz’ rule and the previous bounds,
$|\partial_{\zeta}^{\,k}m_{\gamma}(\zeta)|\leq C_{k,\gamma,W}\sum_{j=0}^{k}(1+|\zeta|)^{\gamma-j}(1+|\zeta|)^{\gamma-(k-j)}\leq C_{k,\gamma,W}\,(1+|\zeta|)^{2\gamma-k},\qquad|\Im\zeta|<W,$
which proves (4.11). ∎
Remark 4.1.
The above conclusion also holds for the multiplier $\widetilde{m}_{\gamma}$ (see 2.8), the proof is entirely analogous and will be omitted.
Proposition 4.6.
Let $\gamma\in\bigl(0,\frac{n}{2}\bigr)$. Then both operators $P_{\gamma}$ and $\widetilde{P}_{\gamma}$ extend by continuity to bounded linear maps
$\widetilde{H}^{\,s}(\mathbb{H}^{n})\longrightarrow\widetilde{H}^{\,s-2\gamma}(\mathbb{H}^{n}),\qquad\forall\,s\in\mathbb{R}.$ (4.12)
Proof: By definition of $\tilde{H}^{s}$, (4.12) is equivalent to the boundedness on $L^{2}(\mathbb{H}^{n})$ of the conjugated operator
$T_{\lambda}:=(\lambda I-\Delta)^{\frac{s-2\gamma}{2}}\,\mathcal{P}_{\gamma}\,(\lambda I-\Delta)^{-\frac{s}{2}}:L^{2}(\mathbb{H}^{n})\to L^{2}(\mathbb{H}^{n}),$ (4.13)
where $\Delta=\Delta_{\mathbb{H}^{n}}$. Indeed, if $u=(\lambda I-\Delta)^{-s/2}f$ with $f\in L^{2}$, then
$\|\mathcal{P}_{\gamma}u\|_{\tilde{H}^{s-2\gamma}}=\bigl\|(\lambda I-\Delta)^{\frac{s-2\gamma}{2}}P_{\gamma}(\lambda I-\Delta)^{-\frac{s}{2}}f\bigr\|_{L^{2}}=\|T_{\lambda}f\|_{L^{2}}.$
Recall $\mathcal{L}=\Delta+\rho^{2}$, hence
$\lambda I-\Delta=(\lambda+\rho^{2})I-\mathcal{L},$ (4.14)
see [30, Scetion 5]. Using (4.8) and (4.14), we may view every factor in (4.13) as a function of the same self-adjoint operator $\sqrt{-\mathcal{L}}$. In spectral variable $\zeta\geq 0$ (corresponding to $\sqrt{-\mathcal{L}}$), we obtain
$T_{\lambda}=\Psi_{\lambda}\!\bigl(\sqrt{-\mathcal{L}}\bigr),\qquad\Psi_{\lambda}(\zeta):=(\lambda+\rho^{2}+\zeta^{2})^{-\gamma}\,m_{\gamma}(\zeta).$
Let $\mathcal{S}_{W}^{m}$ denote the standard strip-holomorphic symbol class appearing in the $L^{p}$ multiplier theorem on noncompact manifolds with bounded geometry: namely $\Phi\in\mathcal{S}_{W}^{m}$ if $\Phi$ is an even function and extends holomorphically to $\{\zeta\in\mathbb{C}:|\Im\zeta|<W\}$ and satisfies $|\partial_{\zeta}^{k}\Phi(\zeta)|\leq C_{k}(1+|\zeta|)^{m-k}$ there. By [30, Chapter IV, Proposition 2.4], it is enough to show that there exists $W>0$ such that
$\Psi_{\lambda}\in\mathcal{S}_{W}^{0}.$ (4.15)
Indeed, taking sufficiently small $\epsilon>0,$ for $W=\frac{\epsilon}{2}<\epsilon$, the functions $\zeta\mapsto(\lambda+\rho^{2}+\zeta^{2})^{-\gamma}$ are holomorphic in $|\Im\zeta|<W$, and a direct differentiation shows
$\bigl|\partial_{\zeta}^{k}(\lambda+\rho^{2}+\zeta^{2})^{-\gamma}\bigr|\leq C_{k,\gamma,\lambda,\rho,W}\,(1+|\zeta|)^{-2\gamma-k},\qquad|\Im\zeta|<W,$
so $(\lambda+\rho^{2}+\zeta^{2})^{-\gamma}\in\mathcal{S}_{\frac{\epsilon}{2}}^{-2\gamma}$.
By Lemma 4.5, the multiplier $m_{\gamma}$ in (4.9) is even and extends holomorphically to every strip $|\Im\zeta|<W<\epsilon$ and satisfies
$|\partial_{\zeta}^{k}m_{\gamma}(\zeta)|\leq C_{k,\gamma,\rho,W}\,(1+|\zeta|)^{2\gamma-k},\qquad|\Im\zeta|<W.$
Hence $m_{\gamma}\in S_{\frac{\epsilon}{2}}^{2\gamma}$. Multiplying the two symbols yields (4.15), see [16, Chapter 18]
$\Psi_{\lambda}(\zeta)=(\lambda+\rho^{2}+\zeta^{2})^{-\gamma}m_{\gamma}(\zeta)\in S_{\frac{\epsilon}{4}}^{0}.$
This completes the proof for $P_{\gamma}$. The argument for $\widetilde{P}_{\gamma}$ is entirely analogous. ∎
Proposition 4.7.
Let $s\in\bigl(0,\frac{n}{2}\bigr)$ and let $\phi\in C_{c}^{\infty}(\mathbb{H}^{n})$. Define the commutator
$\mathcal{P}_{s},\phi]:=\mathcal{P}_{s}\circ M_{\phi}-M_{\phi}\circ\mathcal{P}_{s},\qquad M_{\phi}u:=\phi u.$
Then $\mathcal{P}_{s},\phi]$ extends by continuity to a bounded operator
$\mathcal{P}_{s},\phi]:\widetilde{H}^{\,s-1}(\mathbb{H}^{n})\longrightarrow\widetilde{H}^{-s}(\mathbb{H}^{n}),\qquad\|[\mathcal{P}_{s},\phi]f\|_{\widetilde{H}^{-s}}\leq C_{\phi}\,\|f\|_{\widetilde{H}^{\,s-1}},$ (4.16)
for some constant $C_{\phi}>0$. Moreover, the same conclusion holds with $\mathcal{P}_{s}$ replaced by $\widetilde{\mathcal{P}}_{s}$.
Proof: From the proof of Proposition 4.6, for any $W>0$ small enough one has
$m_{s}\in S_{W}^{2s}.$
By [30, Chapter IV, Proposition 1.2], it is equivalent to $\mathcal{P}_{s}\in\Psi_{W}^{2s}(\mathbb{H}^{n}).$
Multiplication by $\phi$ is a zero-order operator, $M_{\phi}\in\Psi_{W}^{0}(\mathbb{H}^{n})$. By the standard symbolic calculus for pseudodifferential operators, the commutator satisfies
$\mathcal{P}_{s},M_{\phi}]\in\Psi_{W}^{2s-1}(\mathbb{H}^{n}).$
Indeed, if $p_{s}(x,\xi)$ denotes the full symbol of $\mathcal{P}_{s}$, then the symbol of the commutator has an asymptotic expansion beginning with
$\sigma([\mathcal{P}_{s},M_{\phi}])(x,\xi)\sim\sum_{|\alpha|\geq 1}\frac{1}{\alpha!}\,\partial_{\xi}^{\alpha}p_{s}(x,\xi)\,D_{x}^{\alpha}\phi(x),$
so the principal term is of order 2s12s-1.
Finally, we use the Sobolev mapping property for pseudodifferential operators ([30, Proposition 4.5]): for $T\in\Psi_{W}^{m}(\mathbb{H}^{n})$ and any $t\in\mathbb{R}$,
$T:\tilde{H}^{t}(\mathbb{H}^{n})\longrightarrow\tilde{H}^{t-m}(\mathbb{H}^{n})\quad\text{bounded}.$ (4.17)
Taking $T=[\mathcal{P}_{s},\phi]$, m=2s1m=2s-1 and t=s1t=s-1 in (4.17), we obtain
$\mathcal{P}_{s},\phi]:\tilde{H}^{s-1}(\mathbb{H}^{n})\to\tilde{H}^{-s}(\mathbb{H}^{n}),$
together with the estimate (4.16). The constant $C_{\phi}$ depends only on $\phi$. ∎
Lemma 4.8.
Let $\{v_{j}\}\subset C_{c}^{\infty}(\mathbb{H}^{n})$ be a bounded sequence and assume that $v_{j}\rightharpoonup 0$ weakly in $\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$. Then, for every $\phi\in C_{c}^{\infty}(\mathbb{H}^{n})$, the following hold as $j\to\infty$:
  1. (i)
    Let $s\in\bigl(0,\frac{n}{2}\bigr)\setminus\mathbb{N}$ and $\lambda<\lambda_{0,s}^{\mathrm{conf}}$. Then
    $\langle v_{j},\phi^{2}v_{j}\rangle_{\lambda,s}=\mathcal{E}_{\lambda,s}(\phi v_{j})+o(1).$
  2. (ii)
    Let $s\in\bigl(0,\frac{n}{2}\bigr)$ and $\lambda<\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$. Then
    $\langle v_{j},\phi^{2}v_{j}\rangle_{\lambda,s,\sim}=\widetilde{\mathcal{E}}_{\lambda,s}(\phi v_{j})+o(1).$
Proof: (i) Expanding the difference,
$\mathcal{E}_{\lambda,s}(\phi v_{j})-\langle v_{j},\phi^{2}v_{j}\rangle_{\lambda,s}$ $=\langle\phi v_{j},\mathcal{P}_{s}(\phi v_{j})\rangle_{2}-\langle v_{j},\mathcal{P}_{s}(\phi^{2}v_{j})\rangle_{2}$
$=\langle v_{j},\phi\mathcal{P}_{s}(\phi v_{j})-P_{s}(\phi^{2}v_{j})\rangle_{2}=-\langle v_{j},[\mathcal{P}_{s},\phi](\phi v_{j})\rangle_{2}.$
We estimate this duality pairing in $\tilde{H}^{s}\times\tilde{H}^{-s}$:
$|\langle v_{j},[\mathcal{P}_{s},\phi](\phi v_{j})\rangle_{2}|\leq\|v_{j}\|_{H^{s}}\,\|[\mathcal{P}_{s},\phi](\phi v_{j})\|_{\tilde{H}^{-s}}\leq\|v_{j}\|_{\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})}\,\|[\mathcal{P}_{s},\phi](\phi v_{j})\|_{\tilde{H}^{-s}}.$
By Lemma 4.7 with $f=\phi v_{j}$,
$\|[\mathcal{P}_{s},\phi](\phi v_{j})\|_{\tilde{H}^{-s}}\leq C_{\phi}\,\|\phi v_{j}\|_{\tilde{H}^{s-1}}.$
Let $A:=(\lambda I-\Delta_{\mathbb{H}^{n}})^{1/2}$, so that $\|u\|_{\widetilde{H}^{t}}=\|A^{t}u\|_{L^{2}}$ for all $t\in\mathbb{R}$. Let $\{v_{j}\}\subset\widetilde{H}^{s}(\mathbb{H}^{n})$ be bounded and set $f_{j}:=A^{s}v_{j}\in L^{2}(\mathbb{H}^{n})$. Then
$\|\phi v_{j}\|_{\widetilde{H}^{s-1}}=\|A^{s-1}(\phi v_{j})\|_{L^{2}}=\|K_{\phi}f_{j}\|_{L^{2}},$
where
$K_{\phi}:=A^{\,s-1}M_{\phi}A^{-s}=A^{\,s-1}M_{\phi}A^{-(s-1)}\,A^{-1}:L^{2}(\mathbb{H}^{n})\to L^{2}(\mathbb{H}^{n}).$
We claim that $K_{\phi}$ is compact on $L^{2}(\mathbb{H}^{n})$. Indeed, $A^{-1}=(\lambda I-\Delta_{\mathbb{H}^{n}})^{-1/2}$ is an elliptic pseudodifferential operator of order 1-1, hence
$A^{-1}:L^{2}(\mathbb{H}^{n})\longrightarrow H^{1}(\mathbb{H}^{n})\quad\text{bounded}.$
Set
$B_{\phi}:=A^{\,s-1}M_{\phi}A^{-(s-1)}\in\Psi^{0}(\mathbb{H}^{n}),\qquad\text{so that}\qquad K_{\phi}=B_{\phi}\,A^{-1}.$
Choose $\chi\in C_{c}^{\infty}(\mathbb{H}^{n})$ such that $\chi\equiv 1$ on a neighbourhood of $\operatorname{supp}\phi$, and set $K:=\operatorname{supp}\chi\Subset\mathbb{H}^{n}$. Since $M_{\phi}=\chi M_{\phi}$, we have
$B_{\phi}=A^{\,s-1}\chi M_{\phi}A^{-(s-1)}=(\chi A^{\,s-1}\chi)\,M_{\phi}\,A^{-(s-1)}.$
In particular, for every $f\in L^{2}(\mathbb{H}^{n})$,
$K_{\phi}f=B_{\phi}(A^{-1}f)=(\chi A^{\,s-1}\chi)\,M_{\phi}\,A^{-s}f,$
and the operator $\chi A^{\,s-1}\chi$ is properly supported with Schwartz kernel supported in $K\times K$. Therefore,
$\operatorname{supp}(K_{\phi}f)\subset K,\qquad\forall\,f\in L^{2}(\mathbb{H}^{n}).$
Combining the order 0 boundedness of $B_{\phi}$ with the regularizing property of $A^{-1}$, we obtain a bounded map
$K_{\phi}:\ L^{2}(\mathbb{H}^{n})\longrightarrow H^{1}_{\mathrm{loc}}(\mathbb{H}^{n}).$
Together with $\operatorname{supp}(K_{\phi}f)\subset K$, this yields
$K_{\phi}:\ L^{2}(\mathbb{H}^{n})\longrightarrow H^{1}(K)\quad\text{bounded}.$
Since the Rellich–Kondrachov theorem gives the compact embedding $H^{1}(K)\hookrightarrow\hookrightarrow L^{2}(K)$, we obtain
$K_{\phi}:\ L^{2}(\mathbb{H}^{n})\longrightarrow L^{2}(\mathbb{H}^{n})$
is compact. Since $v_{j}\rightharpoonup 0$ weakly in $\tilde{H}^{s}(\mathbb{H}^{n})=\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$, $f_{j}\rightharpoonup 0$ weakly in $L^{2}(\mathbb{H}^{n})$, it follows that $K_{\phi}f_{j}\to 0$ strongly in $L^{2}$. Hence
$\|\phi v_{j}\|_{\tilde{H}^{s-1}(\mathbb{H}^{n})}\longrightarrow 0,\qquad j\to\infty.$
Since $\{v_{j}\}$ is bounded in $\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$, it follows that $\langle v_{j},[\mathcal{P}_{s},\phi](\phi v_{j})\rangle_{2}\to 0$, proving the claim.
(ii) The proof is entirely analogous to that of (i), and we omit the details. ∎
4.3 Proof of Theorem 1.7, 1.8
Proof of Theorem 1.8. Since $\lambda<\lambda_{0,s}^{\mathrm{conf}}$, we have $H_{n,s,p}(\lambda)>0$. Therefore, it suffices to establish the existence of minimizers for $H_{n,s,p}(\lambda)$, see (4.6).
Let $\{u_{j}\}\subset\mathcal{N}_{\lambda,s}$ be a minimizing sequence such that
$I_{\lambda,s}(u_{j})\to H_{n,s}(\lambda).$
Then by (4.7),
$\|u_{j}\|_{p+1}^{p+1}=\mathcal{E}_{\lambda,s}(u_{j})=I_{\lambda,s}(u_{j})^{\frac{p+1}{p-1}}\to H_{n,s}(\lambda)^{\frac{p+1}{p-1}}.$ (4.18)
In particular, $\{u_{j}\}$ is bounded in $\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$.
For $R>0$, define the concentration function
$Q_{j}(R):=\sup_{z_{0}\in\mathbb{R}^{n-1}}\int_{B(z_{0},R)}|u_{j}|^{p+1}\,dV_{\mathbb{H}^{n}},\qquad B(z_{0},R):=\{(r,z)\in\mathbb{H}^{n}:\ r^{2}+|z-z_{0}|^{2}<R^{2}\}.$
Fix any number
$0<\delta<H_{n,s}(\lambda)^{\frac{p+1}{p-1}}.$
By continuity of $R\mapsto Q_{j}(R)$ and (4.18), for each $j$, we can choose $z_{j}\in\mathbb{R}^{n-1}$ and $R_{j}>0$ such that
$\delta=\int_{B(z_{j},R_{j})}|u_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}=\sup_{z_{0}\in\mathbb{R}^{n-1}}\int_{B(z_{0},R_{j})}|u_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}.$ (4.19)
Let $\mathcal{T}_{j}$ be a hyperbolic isometry sending $B(0,1)$ onto $B(z_{j},R_{j})$:
$\mathcal{T}_{j}(r,z)=(R_{j}r,z_{j}+R_{j}z)$
and define $v_{j}:=u_{j}\circ\mathcal{T}_{j}$. By Lemma 4.4, $\{v_{j}\}\subset\mathcal{N}_{\lambda,s}$ is still minimizing and satisfies
$\delta=\int_{B(0,1)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}=\sup_{z_{0}\in\mathbb{R}^{n-1}}\int_{B(z_{0},1)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}.$ (4.20)
By Ekeland’s variational principle, we may assume that $\{v_{j}\}$ is a Palais–Smale sequence for $I_{\lambda,s}$ on $\mathcal{N}_{\lambda,s}$. Equivalently,
$\langle v_{j},\varphi\rangle_{\lambda,s}=\int_{\mathbb{H}^{n}}|v_{j}|^{p-1}v_{j}\,\varphi\,dV_{\mathbb{H}^{n}}+o(1)\qquad\text{for all }\varphi\in\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n}),$ (4.21)
where $o(1)\to 0$ as $j\to\infty$ uniformly for $\varphi$ in bounded sets. Since $\{v_{j}\}$ is bounded in $\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n})$, up to a subsequence,
$v_{j}\rightharpoonup v\ \text{in }\mathcal{H}^{s}_{\lambda}(\mathbb{H}^{n}),\qquad v_{j}\to v\ \text{a.e. in }\mathbb{H}^{n},$
and $v_{j}\to v$ in $L^{q}_{\mathrm{loc}}(\mathbb{H}^{n})$ for every $q<2_{s}^{*}$.
We first claim that $v\not\equiv 0$. Assume by contradiction that $v\equiv 0$. We claim that, for every $z_{0}\in\mathbb{R}^{n-1}$ and every $\phi\in C_{c}^{\infty}(B(z_{0},1))$ with $0\leq\phi\leq 1$, one has
$\int_{\mathbb{H}^{n}}|\phi v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\longrightarrow 0\qquad\text{as }j\to\infty.$ (4.22)
Testing the Palais–Smale identity (4.21) with $\varphi=\phi^{2}v_{j}$ yields
$\langle v_{j},\phi^{2}v_{j}\rangle_{\lambda,s}=\int_{\mathbb{H}^{n}}|v_{j}|^{p-1}(\phi v_{j})^{2}\,dV_{\mathbb{H}^{n}}+o(1).$ (4.23)
By Lemma 4.8,
$\langle v_{j},\phi^{2}v_{j}\rangle_{\lambda,s}=\mathcal{E}_{\lambda,s}(\phi v_{j})+o(1).$ (4.24)
Combining (4.23)–(4.24) we obtain
$\mathcal{E}_{\lambda,s}(\phi v_{j})=\int_{\mathbb{H}^{n}}|v_{j}|^{p-1}(\phi v_{j})^{2}\,dV_{\mathbb{H}^{n}}+o(1).$ (4.25)
By the definition of $H_{n,s,p}(\lambda)$, see (4.6)
$H_{n,s,p}(\lambda)\,\|\phi v_{j}\|_{p+1}^{2}\leq\mathcal{E}_{\lambda,s}(\phi v_{j}).$
Using (4.25) and Hölder’s inequality, we get
$H_{n,s,p}(\lambda)\,\|\phi v_{j}\|_{p+1}^{2}$ $\leq\int_{\mathbb{H}^{n}}|v_{j}|^{p-1}(\phi v_{j})^{2}\,dV_{\mathbb{H}^{n}}+o(1)$
$\leq\|\phi v_{j}\|_{p+1}^{2}\left(\int_{B(z_{0},1)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\right)^{\!\frac{p-1}{p+1}}+o(1).$
If $\|\phi v_{j}\|_{p+1}\not\to 0$, dividing by $\|\phi v_{j}\|_{p+1}^{2}$ yields
$H_{n,s,p}(\lambda)\leq\left(\int_{B(z_{0},1)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\right)^{\!\frac{p-1}{p+1}}+o(1),$
hence
$\liminf_{j\to\infty}\int_{B(z_{0},1)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\geq H_{n,s}(\lambda)^{\frac{p+1}{p-1}}.$ (4.26)
Recall that in (4.20) we fixed $0<\delta<H_{n,s}(\lambda)^{\frac{p+1}{p-1}}$, this contradicts (4.26). Therefore necessarily $\|\phi v_{j}\|_{p+1}\to 0$, which is exactly (4.22).
We now deduce a nontrivial concentration property near $(1,0)\in\mathbb{R}\times\mathbb{R}^{n-1}$. For every $R>0$,
$\liminf_{j\to\infty}\int_{B((1,0),R)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}>0,$ (4.27)
where $B((1,0),R)$ is the Euclidean ball of radius $R$ centered at $(1,0)$.
Fix $R>0$. If $R\geq 1$, then by the normalization (4.20) and (4.22) we obtain (4.27). Assume now $0<R<1$. Suppose by contradiction that
$\int_{B((1,0),R)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\longrightarrow 0.$ (4.28)
Next consider the annulus
$A:=\overline{B((1,0),1)\setminus B((1,0),R)}.$
Since $A$ is compact in the Euclidean topology, it can be covered by finitely many unit balls $B(z_{\ell},1)$, $\ell=1,\dots,L$, with $z_{\ell}\in\mathbb{R}^{n-1}$. For each $\ell$ choose $\phi_{\ell}\in C_{c}^{\infty}(B(z_{\ell},1))$ such that $0\leq\phi_{\ell}\leq 1$ and $\phi_{\ell}\equiv 1$ on a slightly smaller ball. Applying (4.22) to each $\phi_{\ell}$, we obtain
$\int_{B(z_{\ell},1)}|\phi_{\ell}v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\to 0\qquad\text{for every }\ell=1,\dots,L.$
Since $\{\phi_{\ell}\}_{\ell=1}^{L}$ dominates $A$, we deduce
$\int_{A}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\leq\sum_{\ell=1}^{L}\int_{B(z_{\ell},1)}|\phi_{\ell}v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\xrightarrow[]{j\to\infty}0.$
Combining this with (4.28) yields
$\int_{B((1,0),1)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}=\int_{B((1,0),R)}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}+\int_{A}|v_{j}|^{p+1}\,dV_{\mathbb{H}^{n}}\longrightarrow 0,$
which contradicts the normalization (4.20). Hence (4.28) is impossible, and (4.27) follows. This is impossible if $p+1<2_{s}^{*}$, because we assumed $v=0.$ Next we prove the case of $p+1=2_{s}^{*}.$
Fix $R_{0}\in(0,1)$ and $\psi\in C_{c}^{\infty}(B((1,0),R_{0}))$ with $\psi\equiv 1$ on $B((1,0),R_{0}/2)$ and $0\leq\psi\leq 1$. Similarly, we have
$\liminf_{j\to\infty}\int_{B((1,0),R_{0})}|v_{j}|^{2_{s}^{*}}\,dV_{\mathbb{H}^{n}}\geq H_{n,s}(\lambda)^{\frac{p+1}{p-1}},$ (4.29)
and hence
$\int_{\mathbb{H}^{n}\setminus B((1,0),R_{0})}|v_{j}|^{2_{s}^{*}}\,dV_{\mathbb{H}^{n}}\rightarrow 0.$ (4.30)
Proceeding as in (4.23)–(4.24) (with $\phi$ replaced by $\psi$), we obtain
$\mathcal{E}_{\lambda,s}(\psi v_{j})=\int_{\mathbb{H}^{n}}|v_{j}|^{2_{s}^{*}-2}(\psi v_{j})^{2}\,dV_{\mathbb{H}^{n}}+o(1)=\int_{\mathbb{H}^{n}}|\psi v_{j}|^{2_{s}^{*}}\,dV_{\mathbb{H}^{n}}+o(1),$ (4.31)
where in the last equality we used $\operatorname{supp}\psi\subset B((1,0),R_{0})$ together.
Consequently,
$I_{\lambda,s}(\psi v_{j})=\frac{\mathcal{E}_{\lambda,s}(\psi v_{j})}{\|\psi v_{j}\|_{2_{s}^{*}}^{2}}\longrightarrow H_{n,s}(\lambda).$
Moreover, since $\|\psi v_{j}\|_{2}\to 0$,
$H_{n,s}(\lambda)=\lim_{j\to\infty}\frac{\int_{\mathbb{H}^{n}}\psi v_{j}\mathcal{P}_{s}(\psi v_{j})dV_{\mathbb{H}^{n}}}{\|\psi v_{j}\|_{p}^{2}}\geq H_{n,s}(0),$
which contradict with $H_{n,s}(\lambda)<H_{n,s}(0),\lambda\in\mathcal{G}_{n,s}\!\big[H_{n,s}\big]$. Therefore $v\not\equiv 0$, we complete the proof. $\Box$
Proof of Theorem 1.7. The proof is the same as that of Theorem 1.8, except for the localization step. More precisely, the test-function identity (4.23) remains valid when $\mathcal{P}_{s}$ is replaced by $\widetilde{\mathcal{P}}_{s}$, as established in Lemma 4.8. With this modification, the rest of the argument carries over verbatim from the $\mathcal{P}_{s}$–case, and we omit the details. ∎
4.4 Proof of Theorem 1.5, 1.6
Define the operator
$A_{s}:\ C_{c}^{\infty}(\mathbb{B}^{n})\longrightarrow\mathcal{D}^{\prime}(\mathbb{B}^{n})$
by the duality pairing
$\langle A_{s}f,\varphi\rangle:=\int_{\mathbb{B}^{n}}(\widetilde{\mathcal{P}}_{s}f)(x)\,\varphi(x)\,dV_{\mathbb{B}^{n}}(x),\qquad f,\varphi\in C_{c}^{\infty}(\mathbb{B}^{n}).$
In particular, $A_{s}$ is linear by construction.
Lemma 4.9.
AA is continuous: if $f_{j}\to f$ in $C_{c}^{\infty}(\mathbb{B}^{n})$, then for every $\varphi\in C_{c}^{\infty}(\mathbb{B}^{n})$,
$\langle Af_{j},\varphi\rangle\longrightarrow\langle Af,\varphi\rangle.$
Proof: Let $f_{j}\to f$ in $C_{c}^{\infty}(\mathbb{B}^{n})$. Then there exists a compact set $K\Subset\mathbb{B}^{n}$ such that $\operatorname{supp}f_{j}\subset K$ for all $j$, and for every integer $\ell\geq 0$,
$\max_{|\alpha|\leq\ell}\ \sup_{x\in K}\bigl|\partial^{\alpha}(f_{j}-f)(x)\bigr|\longrightarrow 0.$
In particular, for fixed integer $m>2s$ we have
$\|f_{j}-f\|_{H^{m}(\mathbb{B}^{n})}\longrightarrow 0.$
By Lemma 4.6,
$\|\widetilde{\mathcal{P}}_{s}(f_{j}-f)\|_{H^{m-2s}(\mathbb{B}^{n})}\leq C_{m}\,\|f_{j}-f\|_{H^{m}(\mathbb{B}^{n})}\longrightarrow 0.$
Since m2s>0m-2s>0, we have the continuous embedding $H^{m-2s}(\mathbb{B}^{n})\hookrightarrow L^{2}(\mathbb{B}^{n})$, hence
$\|\widetilde{\mathcal{P}}_{s}(f_{j}-f)\|_{L^{2}(\mathbb{B}^{n})}\longrightarrow 0.$
Fix any $\varphi\in C_{c}^{\infty}(\mathbb{B}^{n})\subset L^{2}(\mathbb{B}^{n})$. Then by Cauchy–Schwarz,
$\bigl|\langle Af_{j}-Af,\varphi\rangle\bigr|=\left|\int_{\mathbb{B}^{n}}\bigl(\widetilde{\mathcal{P}}_{s}(f_{j}-f)\bigr)(x)\,\varphi(x)\,dV_{\mathbb{B}^{n}}(x)\right|\leq\|\widetilde{\mathcal{P}}_{s}(f_{j}-f)\|_{L^{2}}\,\|\varphi\|_{L^{2}}\longrightarrow 0.$
This proves $\langle Af_{j},\varphi\rangle\to\langle Af,\varphi\rangle$ for every $\varphi\in C_{c}^{\infty}(\mathbb{B}^{n})$, i.e. $Af_{j}\rightharpoonup Af$ in $\mathcal{D}^{\prime}(\mathbb{B}^{n})$. ∎
Therefore all assumptions of the Schwartz kernel theorem corresponding to smooth manifolds are satisfied, and there exists a unique distribution $K_{s}\in\mathcal{D}^{\prime}(\mathbb{B}^{n}\times\mathbb{B}^{n})$ such that
$\langle\widetilde{\mathcal{P}}_{s}f,\varphi\rangle=\langle K_{s},\,f\otimes\varphi\rangle,\qquad f,\varphi\in C_{c}^{\infty}(\mathbb{B}^{n}),$
where $(f\otimes\varphi)(x,y):=\varphi(x)f(y).$ Equivalently,
$\int_{\mathbb{B}^{n}}(\widetilde{\mathcal{P}}_{s}f)(x)\,\varphi(x)\,dV_{\mathbb{B}^{n}}(x)=\iint_{\mathbb{B}^{n}\times\mathbb{B}^{n}}K_{s}(x,y)\,\varphi(x)\,f(y)\,dV_{\mathbb{B}^{n}}(x)\,dV_{\mathbb{B}^{n}}(y),$
with the right-hand side interpreted as the pairing of $K_{s}$ with the test function $(x,y)\mapsto\varphi(x)f(y)$.
Let $G=\mathrm{Isom}(\mathbb{B}^{n})$ act on functions by
$(L_{g}f)(x):=f(g^{-1}x),\qquad g\in G.$
By Lemma 4.4,
$\widetilde{\mathcal{P}}_{s}L_{g}=L_{g}\widetilde{\mathcal{P}}_{s}\qquad\text{for all }g\in G.$
The commutation relation implies that $K_{s}$ is $G$-invariant in the distributional sense:
[…]
i.e. $\langle K_{s},\Phi\rangle=\langle K_{s},\Phi\circ(g\times g)\rangle$ for all $\Phi\in C_{c}^{\infty}(\mathbb{B}^{n}\times\mathbb{B}^{n})$. Consequently, $K_{s}$ is radial: it depends only on the geodesic distance $d(x,y)$. In particular, there exists a radial distribution $k_{s}$ on $0,\infty)$ such that, for every $f,\varphi\in C_{c}^{\infty}(\mathbb{B}^{n})$,
$\int_{\mathbb{B}^{n}}(\widetilde{\mathcal{P}}_{s}f)(x)\,\varphi(x)\,dV_{\mathbb{B}^{n}}(x)=\iint_{\mathbb{B}^{n}\times\mathbb{B}^{n}}k_{s}\!\bigl(d(x,y)\bigr)\,\varphi(x)\,f(y)\,dV_{\mathbb{B}^{n}}(x)\,dV_{\mathbb{B}^{n}}(y).$
Recall that $T_{x}$ is an isometry of $(\mathbb{B}^{n},g_{\mathbb{B}})$ and satisfies $T_{x}(x)=0$, see subsection 2.1. Hence,
$d(x,y)=d\bigl(T_{x}(x),T_{x}(y)\bigr)=d\bigl(0,T_{x}(y)\bigr).$
Define the radial function on $\mathbb{B}^{n}$ by
$\kappa_{s}(z):=k_{s}\!\bigl(d(0,z)\bigr),\qquad z\in\mathbb{B}^{n}.$
Then, using the hyperbolic convolution (2.1), we obtain that for every $\varphi\in C_{c}^{\infty}(\mathbb{B}^{n})$
$\int_{\mathbb{B}^{n}}(\widetilde{\mathcal{P}}_{s}f)(x)\,\varphi(x)\,dV_{\mathbb{B}^{n}}(x)=\iint_{\mathbb{B}^{n}\times\mathbb{B}^{n}}k_{s}\!\bigl(T_{x}(y))\bigr)\,\varphi(x)\,f(y)\,dV_{\mathbb{B}^{n}}(x)\,dV_{\mathbb{B}^{n}}(y).$
By the Helgason fourier transform, we can get
$\widehat{\kappa_{s}}(\beta)=\widetilde{m}_{s}(\beta)\qquad\beta\in\mathbb{R}.$ (4.32)
in the following sense: for every test function $\psi\in C_{c}^{\infty}(\mathbb{B}^{n})$,
$\langle\kappa_{s},\psi\rangle=\int_{\mathbb{R}}\widetilde{m}_{s}(\beta)\,\Bigl(\int_{\mathbb{B}^{n}}\psi(x)\,\Phi_{\beta}(x)\,dV_{\mathbb{B}^{n}}(x)\Bigr)\,\frac{d\beta}{|c(\beta)|^{2}}.$ (4.33)
Define spherical function (see [21, 24])
$\Phi_{\beta}(x):=\int_{\mathbb{S}^{n-1}}h_{-\beta,\theta}(x)d\sigma(\theta).$
Then $\Phi_{\beta}(0)=1,$ $\Phi_{\beta}(x)$ is ridial and has the explicit Legendre representation
$\Phi_{\beta}(x):=2^{\frac{n-2}{2}}\,\Gamma\!\Bigl(\frac{n}{2}\Bigr)\,(\sinh d(x,0))^{\frac{2-n}{2}}\,P_{-\frac{1}{2}+i\beta}^{\,\frac{2-n}{2}}(\cosh d(x,0)),$
where $P_{\nu}^{\mu}$ is the associated Legendre function of the first kind. It is known that $P_{\nu}^{\mu}$ can be defined via hypergeometric function:
$P_{\nu}^{\mu}(z)=\frac{1}{\Gamma(1-\mu)}\left(\frac{z+1}{z-1}\right)^{\frac{\mu}{2}}\,{}_{2}F_{1}\!\left(-\nu,\nu+1;1-\mu;\frac{1-z}{2}\right),\quad\mu\neq 1,2,3\cdots$
where $_{2}F_{1}(a,b;c;z)$ is hypergeometric function
$_{2}F_{1}(a,b;c;z)=\sum_{k=0}^{\infty}\frac{(a)_{k}(b)_{k}}{(c)_{k}}\,\frac{z^{k}}{k!},\qquad c\neq 0,-1,-2,\dots$
Proposition 4.10.
Let $n\geq 2,s\in\bigl(0,\frac{n}{2}\bigr)$ and $\rho=\frac{n-1}{2}$. Then for all $f,g\in C_{c}^{\infty}(\mathbb{B}^{n})$ with
$\operatorname{dist}(\operatorname{supp}f,\operatorname{supp}g)\geq R>1,$
there exists a constant $C=C(s,n,f,g)>0$ such that
$\bigl|\langle\widetilde{\mathcal{P}}_{s}f,\,g\rangle_{L^{2}(\mathbb{B}^{n})}\bigr|\leq C\,e^{-\rho R},\quad\bigl|\langle{\mathcal{P}}_{s}f,\,g\rangle_{L^{2}(\mathbb{B}^{n})}\bigr|\leq C\,e^{-\rho R}.$
Proof: When $s=k\in\mathbb{N}$, since $\widetilde{\mathcal{P}}_{k}$ is local differential operator, $\operatorname{supp}(\widetilde{\mathcal{P}}_{k}f)\subset\operatorname{supp}f$. The assumption $\operatorname{dist}(\operatorname{supp}f,\operatorname{supp}g)\geq R>0$ implies $\operatorname{supp}(\widetilde{\mathcal{P}}_{k}f)\cap\operatorname{supp}g=\varnothing$, hence $\langle\widetilde{\mathcal{P}}_{k}f,\,g\rangle=0$.
When $s\in(0,\frac{n}{2})\setminus\mathbb{N}$, by Fubini’s theorem,
$\langle\widetilde{\mathcal{P}}_{s}f,g\rangle_{L^{2}(\mathbb{B}^{n})}$ $=\int_{\mathbb{B}^{n}}\!\!\int_{\mathbb{B}^{n}}\kappa_{s}\!\bigl(d(x,y)\bigr)\,f(y)\,g(x)\,dV_{\mathbb{B}^{n}}(x)dV_{\mathbb{B}^{n}}(y).$
By Fubini’s theorem and the change of variables $x=T_{y}(z)$, this becomes
$\langle\widetilde{\mathcal{P}}_{s}f,g\rangle=\int_{\mathbb{B}^{n}}f(y)\,\Big\langle\kappa_{s},\psi_{y}\Big\rangle\,dV_{\mathbb{B}^{n}}(y),\qquad\psi_{y}(z):=g(T_{y}(z))\in C_{c}^{\infty}(\mathbb{B}^{n}).$
Applying the distributional inversion formula for $\kappa_{s}$ (4.33) yields
$\Big\langle\kappa_{s},\psi_{y}\Big\rangle=\int_{\mathbb{R}}\widetilde{m}_{s}(\beta)\Bigl(\int_{\mathbb{B}^{n}}\psi_{y}(z)\Phi_{\beta}(z)\,dV_{\mathbb{B}^{n}}(z)\Bigr)\frac{d\beta}{|c(\beta)|^{2}}.$
Changing variables $z=T_{y}(x)$ in the inner integral, we obtain
$\int_{\mathbb{B}^{n}}\psi_{y}(z)\Phi_{\beta}(z)\,dV_{\mathbb{B}^{n}}(z)=\int_{\mathbb{B}^{n}}g(x)\,\Phi_{\beta}(T_{y}(x))\,dV_{\mathbb{B}^{n}}(x).$
Substituting back and exchanging the order of integration gives
$\langle\widetilde{\mathcal{P}}_{s}f,g\rangle_{L^{2}(\mathbb{B}^{n})}=\int_{\mathbb{R}}\widetilde{m}_{s}(\beta)\,\Bigl(\iint_{\mathbb{B}^{n}\times\mathbb{B}^{n}}f(y)\,g(x)\,\Phi_{\beta}\!\bigl(T_{y}(x)\bigr)\,dV_{\mathbb{B}^{n}}(x)\,dV_{\mathbb{B}^{n}}(y)\Bigr)\,\frac{d\beta}{|c(\beta)|^{2}}.$ (4.34)
Let $K_{f}:=\operatorname{supp}f$ and $K_{g}:=\operatorname{supp}g$. By assumption, $r:=d(x,y)\geq R$ for all $(x,y)\in K_{g}\times K_{f}$. Since $\Phi_{\beta}$ is radial and $T_{y}$ is an isometry sending $y$ to 0, we have $\Phi_{\beta}\bigl(T_{y}(x)\bigr)=\Phi_{\beta}\bigl(r\bigr),$ where
$\Phi_{\beta}(r)=2^{\frac{n-2}{2}}\,\Gamma\!\Bigl(\frac{n}{2}\Bigr)\,(\sinh r)^{\frac{2-n}{2}}\,P_{-\frac{1}{2}+i\beta}^{\,\frac{2-n}{2}}(\cosh r).$
Hence, using (4.34) and the Fubini Theorem, we can write
$\big\langle\widetilde{\mathcal{P}}_{s}f,g\big\rangle=\iint_{K_{g}\times K_{f}}f(y)g(x)\,k_{s}(r(x,y))\,dV_{\mathbb{B}^{n}}(x)\,dV_{\mathbb{B}^{n}}(y),\qquad k_{s}(r):=\int_{\mathbb{R}}\widetilde{m}_{s}(\beta)\Phi_{\beta}(r)\frac{d\beta}{|c(\beta)|^{2}}.$
Therefore,
$\Big|\big\langle\widetilde{\mathcal{P}}_{s}f,g\big\rangle\Big|\leq\|f\|_{L^{1}}\|g\|_{L^{1}}\,\sup_{r\geq R}|k_{s}(r)|.$ (4.35)
In particular, one should not take absolute values inside the $\beta$–integral, since this would destroy the oscillatory cancellation in $\beta$ coming from the large–$r$ asymptotics of $\Phi_{\beta}(r)$.
We use the Harish–Chandra expansion from [28, Theorem 3.1] (for rank one) which, in our notation, reads
$\Phi_{\beta}(r)=c(\beta)e^{(i\beta-\rho)r}\phi_{\beta}(r)+c(-\beta)e^{(-i\beta-\rho)r}\phi_{-\beta}(r),\qquad\phi_{\beta}(r)=\sum_{k=0}^{\infty}\Gamma_{k}(\beta)e^{-2kr},$ (4.36)
with $\Gamma_{0}(\beta)\equiv 1$ and $\rho=\frac{n-1}{2}$. We estimate $k_{s}(r)$ for $r\geq 1$. Using (4.36),
$k_{s}(r)=e^{-\rho r}\Big(J_{+}(r)+J_{-}(r)\Big),$
where
$J_{+}(r):=\int_{\mathbb{R}}e^{i\beta r}\,\widetilde{m}_{s}(\beta)\,\frac{\phi_{\beta}(r)}{\overline{c(\beta)}}\,d\beta,\qquad J_{-}(r):=\int_{\mathbb{R}}e^{-i\beta r}\,\widetilde{m}_{s}(\beta)\,\frac{\phi_{-\beta}(r)}{\overline{c(-\beta)}}\,d\beta.$
We treat $J_{+}$ (the other term is analogous). Split $\phi_{\beta}(r)=1+\big(\phi_{\beta}(r)-1\big)$. Accordingly,
$J_{+}(r)=J_{+}^{(0)}(r)+J_{+}^{(1)}(r),\qquad J_{+}^{(0)}(r):=\int_{\mathbb{R}}e^{i\beta r}\,\widetilde{m}_{s}(\beta)\,\overline{c(\beta)}^{-1}\,d\beta.$
Let $a(\beta):=\widetilde{m}_{s}(\beta)\,\overline{c(\beta)}^{-1}$. By Stirling estimates for Gamma ratios and their derivatives together with Lemma 4.2 of [28], we have for every integer $k\geq 0$ the symbol estimate
$|a^{(k)}(\beta)|\leq C_{k}\,(1+|\beta|)^{2s+\frac{n-1}{2}-k},\qquad\beta\in\mathbb{R}.$ (4.37)
Choose $N\in\mathbb{N}$ so large that $N>2s+\frac{n+1}{2}$. Then $a^{(N)}\in L^{1}(\mathbb{R}\setminus(-1,1))$. Let $\chi\in C_{c}^{\infty}(\mathbb{R})$ satisfy $\chi\equiv 1$ on [1,1][-1,1] and $\chi\equiv 0$ on $|\beta|\geq 2$. Write
$J_{+}^{(0)}(r)=\int\chi(\beta)a(\beta)e^{i\beta r}\,d\beta+\int(1-\chi(\beta))a(\beta)e^{i\beta r}\,d\beta=:J_{+,\mathrm{low}}^{(0)}(r)+J_{+,\mathrm{high}}^{(0)}(r).$
Since $\chi a$ is compactly supported and smooth, repeated integration by parts gives $|J_{+,\mathrm{low}}^{(0)}(r)|\leq C_{N}r^{-N}$. For the high–frequency part, integrating by parts $N$ times yields
$J_{+,\mathrm{high}}^{(0)}(r)=\frac{1}{(ir)^{N}}\int\partial_{\beta}^{N}\!\big((1-\chi(\beta))a(\beta)\big)\,e^{i\beta r}\,d\beta,$
hence by (4.37),
$|J_{+,\mathrm{high}}^{(0)}(r)|\leq r^{-N}\int_{\mathbb{R}}\Big|\partial_{\beta}^{N}\!\big((1-\chi)a\big)\Big|\,d\beta\leq C_{N}r^{-N}.$
Therefore,
$|J_{+}^{(0)}(r)|\leq C_{N}r^{-N},\qquad r\geq 1.$ (4.38)
By (4.36),
$\phi_{\beta}(r)-1=\sum_{k\geq 1}\Gamma_{k}(\beta)e^{-2kr}=e^{-2r}\sum_{k\geq 1}\Gamma_{k}(\beta)e^{-2(k-1)r}.$
For each fixed $r\geq 1$, the series converges uniformly in $\beta$ on compact sets [28, Remark 1 after Theorem 3.1]. Moreover, the coefficients $\Gamma_{k}(\beta)$ have at most polynomial growth in $\beta$ (see the discussion following Theorem 3.1 and the estimates on $\Gamma_{k}$ in Section 3 of [28]), hence the factor $e^{-2r}$ implies that $J_{+}^{(1)}(r)$ enjoys the same oscillatory integration–by–parts bound as $J_{+}^{(0)}(r)$, up to an additional $e^{-2r}$ factor. Concretely, repeating the above argument with the amplitude $a(\beta)\big(\phi_{\beta}(r)-1\big)$ yields
$|J_{+}^{(1)}(r)|\leq C_{N}e^{-2r}r^{-N},\qquad r\geq 1.$ (4.39)
Combining (4.38)–(4.39) (and the analogous bounds for $J_{-}$) gives
$|k_{s}(r)|\leq C_{N}e^{-\rho r}r^{-N}\lesssim e^{-\rho r},\qquad r\geq 1.$
Thus $\sup_{r\geq R}|k_{s}(r)|\leq C_{N}e^{-\rho R}$ for $R\geq 1$. Plugging this into (4.35) yields the desired estimate.
The corresponding estimate for $\mathcal{P}_{s}$ can be obtained in the same way. ∎
Proposition 4.11.
Let $n\geq 2$ and $s\in\bigl(0,\frac{n}{2}\bigr)$. Then,
(i) for $\lambda>\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$, $\widetilde{H}_{n,s}(\lambda)=-\infty;$
(ii) for $\lambda>\lambda_{0,s}^{\mathrm{conf}}$, $H_{n,s}(\lambda)=-\infty.$
Proof: (i) Fix $\lambda>\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$. By the definition of $\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$, there exists $\phi\in C_{c}^{\infty}(\mathbb{H}^{n})\setminus\{0\}$ such that
$Q_{\lambda}(\phi):=\int_{\mathbb{H}^{n}}(\widetilde{\mathcal{P}}_{s}\phi)\,\phi\,dV_{\mathbb{H}^{n}}-\lambda\int_{\mathbb{H}^{n}}|\phi|^{2}\,dV_{\mathbb{H}^{n}}<0.$
Set $q:=-Q_{\lambda}(\phi)>0$ and $p:=2_{s}^{*}$.
Step 1: many far-apart copies. Let $K:=\operatorname{supp}\phi$, which is compact. For each $N\in\mathbb{N}$ we choose isometries $\{\tau_{j}\}_{j=1}^{N}\subset\mathrm{Isom}(\mathbb{H}^{n})$ such that the sets
$K_{j}:=\tau_{j}(K)\quad\text{are pairwise disjoint and satisfy}\quad\operatorname{dist}(K_{i},K_{j})\geq R_{N}\ \ \text{for all }i\neq j,$
where $R_{N}>0$ will be chosen later. Define $\phi_{j}:=\phi\circ\tau_{j}^{-1}$ and
$u_{N}:=\sum_{j=1}^{N}\phi_{j}\in C_{c}^{\infty}(\mathbb{H}^{n}).$
Since isometries preserve $dV_{\mathbb{H}^{n}}$ and the supports are disjoint, we have
$\int_{\mathbb{H}^{n}}|u_{N}|^{2}\,dV_{\mathbb{H}^{n}}=\sum_{j=1}^{N}\int_{\mathbb{H}^{n}}|\phi_{j}|^{2}\,dV_{\mathbb{H}^{n}}=N\int_{\mathbb{H}^{n}}|\phi|^{2}\,dV_{\mathbb{H}^{n}},$ (4.40)
and
$\int_{\mathbb{H}^{n}}|u_{N}|^{p}\,dV_{\mathbb{H}^{n}}=\sum_{j=1}^{N}\int_{\mathbb{H}^{n}}|\phi_{j}|^{p}\,dV_{\mathbb{H}^{n}}=N\int_{\mathbb{H}^{n}}|\phi|^{p}\,dV_{\mathbb{H}^{n}}.$ (4.41)
Hence the denominator satisfies
$\Bigl(\int_{\mathbb{H}^{n}}|u_{N}|^{p}\,dV_{\mathbb{H}^{n}}\Bigr)^{\!2/p}=N^{2/p}\Bigl(\int_{\mathbb{H}^{n}}|\phi|^{p}\,dV_{\mathbb{H}^{n}}\Bigr)^{\!2/p}=N^{\frac{n-2s}{n}}\Bigl(\int_{\mathbb{H}^{n}}|\phi|^{p}\,dV_{\mathbb{H}^{n}}\Bigr)^{\!2/p}.$ (4.42)
Step 2: control of the quadratic form. Write $u_{N}=\sum_{j=1}^{N}\phi_{j}$. Expanding the quadratic form gives
$Q_{\lambda}(u_{N})$ $=\sum_{j=1}^{N}Q_{\lambda}(\phi_{j})\;+\;2\sum_{1\leq i<j\leq N}\int_{\mathbb{H}^{n}}(\widetilde{\mathcal{P}}_{s}\phi_{i})\,\phi_{j}\,dV_{\mathbb{H}^{n}}.$
By Lemma 4.4 and $dV_{\mathbb{H}^{n}}$, $Q_{\lambda}(\phi_{j})=Q_{\lambda}(\phi)=-q$, thus $\sum_{j=1}^{N}Q_{\lambda}(\phi_{j})=-Nq$.
By Proposition 4.10, there exist constants $\alpha>0$ and $C>0$ (depending on $n,s$ and $\phi$) such that whenever $\operatorname{dist}(\operatorname{supp}f,\operatorname{supp}g)\geq R$,
$\biggl|\int_{\mathbb{H}^{n}}(\widetilde{\mathcal{P}}_{s}f)\,g\,dV_{\mathbb{H}^{n}}\biggr|\leq C\,e^{-\alpha R}.$ (4.43)
Applying (4.43) with $f=\phi_{i}$, $g=\phi_{j}$ and $R=R_{N}$ yields
$\biggl|\sum_{1\leq i<j\leq N}\int_{\mathbb{H}^{n}}(\mathcal{P}_{s}\phi_{i})\,\phi_{j}\,dV_{\mathbb{H}^{n}}\biggr|\leq C\binom{N}{2}e^{-\alpha R_{N}}\leq CN^{2}e^{-\alpha R_{N}}.$
Therefore
$Q_{\lambda}(u_{N})\leq-Nq+2CN^{2}e^{-\alpha R_{N}}.$ (4.44)
Choose $R_{N}:=\frac{2}{\alpha}\log N+R_{0}$ with $R_{0}>0$ so large that $2Ce^{-\alpha R_{0}}\leq\frac{q}{4}$. Then $2CN^{2}e^{-\alpha R_{N}}\leq\frac{q}{4}\,N$, and (4.44) becomes
$Q_{\lambda}(u_{N})\leq-\frac{3q}{4}\,N\leq-\frac{q}{2}\,N\qquad\text{for all sufficiently large }N.$ (4.45)
Combining (4.42) and (4.45), for large $N$ we obtain
$\frac{Q_{\lambda}(u_{N})}{\bigl(\int_{\mathbb{H}^{n}}|u_{N}|^{p}\,dV_{\mathbb{H}^{n}}\bigr)^{2/p}}\leq-\frac{q}{2}\,\frac{N}{N^{2/p}}\,\Bigl(\int_{\mathbb{H}^{n}}|\phi|^{p}\,dV_{\mathbb{H}^{n}}\Bigr)^{-2/p}\\ =-kN^{1-\frac{2}{p}}=-kN^{\frac{2s}{n}}\longrightarrow-\infty,$
where $k:=\dfrac{q}{2}\Bigl(\int_{\mathbb{H}^{n}}|\phi|^{p}\,dV_{\mathbb{H}^{n}}\Bigr)^{-2/p}>0$. Taking the infimum over $u$ gives $H_{n,s}(\lambda)=$-$\infty$.
(ii) The same argument applies to $\mathcal{P}_{s}$ as well, and we omit the proof. ∎
Proof of Theorem 1.5: (i) By Theorem 1.3 (i) and Proposition 5.3, we obtain $\mathcal{G}_{n,s}[\widetilde{H}_{n,s}]=(0,+\infty)$. By the proof of Theorem 1.7, $\widetilde{H}_{n,s}(\lambda)$ is achieved for every $\lambda\in(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}})$, thus by Lemma 5.6 (ii), $\widetilde{H}_{n,s}(\lambda)$ is strictly decreasing in $(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}})$. The remaining conclusions follow from Proposition 4.11, Proposition 5.3 and Lemma 5.5.
(ii) By Theorem 1.3 (ii), there exists $\widetilde{\lambda}_{s}^{\mathrm{conf}}\in(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}]$ such that $\bigl(\widetilde{\lambda}_{s}^{\mathrm{conf}},\infty\bigr)\subset\mathcal{G}_{n,s}\bigl[\widetilde{H}_{n,s}\bigr]$, the remaining conclusions follow from Proposition 4.11, Proposition 5.3 and Lemma 5.5.
(iii) By [24, Theorem 1.9], we obtain $\widetilde{H}_{n,s}\!\bigl(\widetilde{\lambda}_{0,s}^{\mathrm{conf}}\bigr)=S_{n,s}.$ By Proposition 5.3, $\widetilde{H}_{n,s}\!\bigl(0\bigr)=S_{n,s},$ thus we complete the proof.
Proof of Theorem 1.6: (i) For $s\in\left(0,\frac{n}{4}\right]\cap\mathcal{B}_{0}$, we have $b_{s}=0.$ By Theorem 1.4 (i), we obtain $H_{n,s}(\lambda)<S_{n,s}\ \,{\rm for}\ \lambda>0,$ the remaining conclusions follow from Proposition 5.3.
(ii) For $s\in\left(0,\frac{n}{4}\right]\cap\mathcal{B}_{+}$, we have $\lambda_{0,s}^{\mathrm{conf}}>b_{s}>0$. By Proposition 5.3 and (1.6) we know $S_{n,s}=\widetilde{H}_{n,s}(0)\leq H_{n,s}(0)\leq S_{n,s},$ thus $H_{n,s}(0)=S_{n,s}.$ Then by Theorem 1.4 (i), we have $(b_{s},\infty)\subset\mathcal{G}_{n,s}\bigl[H_{n,s}\bigr].$ By the proof of Theorem 1.8 and Lemma 5.6 (ii), $H_{n,s}(\lambda)$ is achieved and strictly decreasing for every $\lambda\in(b_{s},\lambda_{0,s}^{\mathrm{conf}})$.
(iii) follows immediately by combining Theorem 1.4 (b) with Proposition 5.3.
(iv) By [24, Theorem 1.4] and Proposition 5.3, for $\lambda\leq{\lambda}_{0,s}^{\mathrm{conf}}$, we obtain $H_{n,s}\!\bigl({\lambda}_{0,s}^{\mathrm{conf}}\bigr)=S_{n,s}=H_{n,s}(\lambda).$ The remaining conclusions follow from Proposition 4.11.
5  Appendix
5.1 Stability and Attainability of Sobolev Levels
Proposition 5.1.
Let $\Omega\subset\mathbb{R}^{n}$ be a nonempty open set, $n\geq 3$ and $2^{*}=\frac{2n}{n-2}$. For $p\in(0,\infty)$ and $\mu>0$ define
$S^{(p)}_{\Omega}:=\inf_{v\in C_{c}^{\infty}(\Omega)\setminus\{0\}}\frac{\displaystyle\int_{\mathbb{R}^{n}}\bigl(|\nabla v|^{2}+\mu|v|^{p}\bigr)\,dx}{\Bigl(\displaystyle\int_{\mathbb{R}^{n}}|v|^{2^{*}}\,dx\Bigr)^{2/2^{*}}},\quad S_{\Omega}:=\inf_{v\in C_{c}^{\infty}(\Omega)\setminus\{0\}}\frac{\displaystyle\int_{\mathbb{R}^{n}}|\nabla v|^{2}\,dx}{\Bigl(\displaystyle\int_{\mathbb{R}^{n}}|v|^{2^{*}}\,dx\Bigr)^{2/2^{*}}}.$
Then
$S^{(p)}_{\Omega}=S_{\Omega}=S_{n,1}.$
where $S_{n,1}$ is defined in (1.12).
Proof: For $v\in C_{c}^{\infty}(\Omega)\setminus\{0\},$ set
$S_{\Omega}(v):=\frac{\displaystyle\int_{\mathbb{R}^{n}}|\nabla v|^{2}\,dx}{\Bigl(\displaystyle\int_{\mathbb{R}^{n}}|v|^{2^{*}}\,dx\Bigr)^{2/2^{*}}},\qquad S^{(p)}_{\Omega}(v):=\frac{\displaystyle\int_{\mathbb{R}^{n}}\bigl(|\nabla v|^{2}+\mu|v|^{p}\bigr)\,dx}{\Bigl(\displaystyle\int_{\mathbb{R}^{n}}|v|^{2^{*}}\,dx\Bigr)^{2/2^{*}}}.$
It is obvious that
$S^{(p)}_{\Omega}=\inf S^{(p)}_{\Omega}(v)\;\geq\;\inf S_{\Omega}(v)=S_{\Omega}.$
Next, we show that $S^{(p)}_{\Omega}\leq S_{\Omega}$ We split the argument into the cases $p\neq 2$ and $p=2$.
Case A: $p\neq 2$. Fix $u\in C_{c}^{\infty}(\Omega)\setminus\{0\}$ and set
$A:=\int_{\mathbb{R}^{n}}|\nabla u|^{2}\,dx,\quad B:=\int_{\mathbb{R}^{n}}|u|^{p}\,dx,\quad C:=\Bigl(\int_{\mathbb{R}^{n}}|u|^{2^{*}}\,dx\Bigr)^{2/2^{*}}>0.$
For $t>0$ consider the amplitude scaling $u_{t}:=t\,u$. Then $u_{t}\in C_{c}^{\infty}(\Omega)$ and
$\int_{\mathbb{R}^{n}}|\nabla u_{t}|^{2}\,dx=t^{2}A,\quad\int_{\mathbb{R}^{n}}|u_{t}|^{p}\,dx=t^{p}B,\quad\Bigl(\int_{\mathbb{R}^{n}}|u_{t}|^{2^{*}}\,dx\Bigr)^{2/2^{*}}=t^{2}C.$
Therefore
$S^{(p)}_{\Omega}(u_{t})=\frac{t^{2}A+\mu t^{p}B}{t^{2}C}=S_{\Omega}(u)+\mu\,t^{p-2}\,\frac{B}{C}.$
If $p>2$, let $t\to 0^{+}$; then $t^{p-2}\to 0$ and
$\lim_{t\to 0^{+}}S^{(p)}_{\Omega}(u_{t})=S_{\Omega}(u).$
If $0<p<2$, let $t\to+\infty$; then $t^{p-2}\to 0$ and the same limit holds. In both cases we obtain
$S^{(p)}_{\Omega}\leq\inf_{t>0}S^{(p)}_{\Omega}(u_{t})\leq S_{\Omega}(u).$
Since $u$ is arbitrary, taking the infimum over $u$ yields $S^{(p)}_{\Omega}\leq S_{\Omega}.$
Case B: $p=2$. It is well known that $S_{\Omega}$ coincides with the Sobolev constant on any ball contained in $\Omega$. Choose $x_{0}\in\Omega$ and $r>0$ such that $B_{r}(x_{0})\subset\Omega$, and let
$S_{\Omega}=S_{B_{r}(x_{0})}=\inf_{v\in C_{c}^{\infty}(B_{r}(x_{0}))\setminus\{0\}}S_{B_{r}(x_{0})}(v).$
Hence, for any given $\varepsilon>0$, there exists $u\in C_{c}^{\infty}(B_{r}(x_{0}))\setminus\{0\}$ such that
$S_{B_{r}(x_{0})}(u)\leq S_{\Omega}+\varepsilon.$
For $\lambda\geq 1$, define the critical Sobolev scaling around $x_{0}$,
$u_{\lambda}(x):=\lambda^{\frac{N-2}{2}}u\bigl(x_{0}+\lambda(x-x_{0})\bigr).$
Then $u_{\lambda}\in C_{c}^{\infty}(B_{r/\lambda}(x_{0}))\subset B_{r}(x_{0})\subset\Omega$, and a direct change-of-variables computation gives
$\int_{\mathbb{R}^{n}}|\nabla u_{\lambda}|^{2}\,dx=\int_{\mathbb{R}^{n}}|\nabla u|^{2}\,dx,\quad\int_{\mathbb{R}^{n}}|u_{\lambda}|^{2^{*}}\,dx=\int_{\mathbb{R}^{n}}|u|^{2^{*}}\,dx$
and
$\int_{\mathbb{R}^{n}}|u_{\lambda}|^{2}\,dx=\lambda^{-2}\int_{\mathbb{R}^{n}}|u|^{2}\,dx.$
Therefore,
$S^{(2)}_{\Omega}(u_{\lambda})=\frac{\displaystyle\int_{\mathbb{R}^{n}}|\nabla u|^{2}\,dx+\mu\lambda^{-2}\displaystyle\int_{\mathbb{R}^{n}}|u|^{2}\,dx}{\Bigl(\displaystyle\int_{\mathbb{R}^{n}}|u|^{2^{*}}\,dx\Bigr)^{2/2^{*}}}=S_{B_{r}(x_{0})}(u)+\mu\lambda^{-2}\,\frac{\displaystyle\int_{\mathbb{R}^{n}}|u|^{2}\,dx}{\Bigl(\displaystyle\int_{\mathbb{R}^{n}}|u|^{2^{*}}\,dx\Bigr)^{2/2^{*}}}.$
Letting $\lambda\to\infty$ we obtain
$\lim_{\lambda\to\infty}S^{(2)}_{\Omega}(u_{\lambda})=S_{B_{r}(x_{0})}(u)\leq S_{\Omega}+\varepsilon.$
Hence, $S^{(2)}_{\Omega}\leq S_{\Omega}+\varepsilon$ for every $\varepsilon>0$, which implies $S^{(2)}_{\Omega}\leq S_{\Omega}$. ∎
Proposition 5.2.
Let $\Omega\subset\mathbb{R}^{n}$ be a nonempty open set, $n>2s,s\in(0,1)$ and $2_{s}^{*}=\frac{2n}{n-2s}$. For $p\in(0,\infty)$ and $\mu>0$ define
$S^{(p)}_{\Omega,s}:=\inf_{u\in C_{c}^{\infty}(\Omega)\setminus\{0\}}\frac{\displaystyle\int_{\Omega}u\,(-\Delta)^{s}u\,dx+\mu\displaystyle\int_{\mathbb{R}^{n}}|u|^{p}\,dx}{\bigl(\displaystyle\int_{\mathbb{R}^{n}}|u|^{2_{s}^{*}}\,dx\bigr)^{2/2_{s}^{*}}},\quad S_{\Omega,s}:=\inf_{u\in C_{c}^{\infty}(\Omega)\setminus\{0\}}\frac{\displaystyle\int_{\Omega}u\,(-\Delta)^{s}u\,dx}{\bigl(\displaystyle\int_{\mathbb{R}^{n}}|u|^{2_{s}^{*}}\,dx\bigr)^{2/2_{s}^{*}}}.$
Then
$S^{(p)}_{\Omega,s}=S_{\Omega,s}=S_{n,s}\qquad\text{for all }p>0.$
where $S_{n,s}$ is defined in (1.12).
Proof: The argument is entirely analogous to the local case in Proposition 5.1: one employs the critical dilation
$u_{t}(x):=t^{\frac{n-2s}{2}}\,u\bigl(x_{0}+t(x-x_{0})\bigr),\qquad t\geq 1,$
which preserves both the quadratic energy and the critical normalization,
$\int_{\mathbb{R}^{n}}u_{t}\,(-\Delta)^{s}u_{t}\,dx=\int_{\mathbb{R}^{n}}u\,(-\Delta)^{s}u\,dx,\qquad\Bigl(\int_{\mathbb{R}^{n}}|u_{t}|^{2_{s}^{*}}\,dx\Bigr)^{\!\frac{2}{2_{s}^{*}}}=\Bigl(\int_{\mathbb{R}^{n}}|u|^{2_{s}^{*}}\,dx\Bigr)^{\!\frac{2}{2_{s}^{*}}}.$
Thus, we omit the details. ∎
Proposition 5.3.
Let $n\geq 2$ and $s\in(0,\tfrac{n}{2})$. Then for all $\lambda\leq 0$,
$H_{n,s}(0)\leq H_{n,s}(\lambda)\leq S_{n,s};\quad\widetilde{H}_{n,s}(\lambda)=\widetilde{H}_{n,s}(0)=S_{n,s}$
where $S_{n,s},H_{n,s}(\lambda),\widetilde{H}_{n,s}(\lambda)$ are defined in (1.12) (1.1)(1.7).
Proof: It is sufficient to show that $H_{n,s}(\lambda)\leq S_{n,s}$ and $\widetilde{H}_{n,s}(\lambda)\leq S_{n,s}$. Fix $p\in\mathbb{H}^{n}$ and choose $0\not\equiv\phi\in C_{c}^{\infty}(B_{1}(0))$. For $\varepsilon\in(0,1)$, in geodesic normal coordinates at $p$, define
$u_{\varepsilon}(x):=\varepsilon^{-\frac{n-2s}{2}}\,\phi\Bigl(\frac{\exp_{p}^{-1}(x)}{\varepsilon}\Bigr),\qquad x\in\mathbb{H}^{n}.$
Then $u_{\varepsilon}\in C_{c}^{\infty}(\mathbb{H}^{n})$ and $\operatorname{supp}u_{\varepsilon}\subset B_{\mathbb{H}^{n}}(p,\varepsilon)$.
Write $x=\exp_{p}(\varepsilon z)$, so $y=\exp_{p}^{-1}(x)=\varepsilon z$ and $u_{\varepsilon}(\exp_{p}(\varepsilon z))=\varepsilon^{-\frac{n-2s}{2}}\phi(z)$. In normal coordinates,
$g_{ij}(y)=\delta_{ij}+O(|y|^{2}),\qquad dV_{\mathbb{H}^{n}}(y)=\bigl(1+O(|y|^{2})\bigr)\,dy,$
hence under $y=\varepsilon z$,
$dV_{\mathbb{H}^{n}}(x)=\bigl(1+O(\varepsilon^{2}|z|^{2})\bigr)\,\varepsilon^{n}\,dz,$
uniformly on the support of $\phi$.
[…]
$\int_{\mathbb{H}^{n}}|u_{\varepsilon}|^{2_{s}^{*}}\,dV_{\mathbb{H}^{n}}=\int_{\mathbb{R}^{n}}|\phi(z)|^{2_{s}^{*}}\,\bigl(1+O(\varepsilon^{2}|z|^{2})\bigr)\,dz=\int_{\mathbb{R}^{n}}|\phi|^{2_{s}^{*}}\,dz+O(\varepsilon^{2}).$
Similarly,
$\int_{\mathbb{H}^{n}}|u_{\varepsilon}|^{2}\,dV_{\mathbb{H}^{n}}=\int_{\mathbb{R}^{n}}\varepsilon^{-(n-2s)}|\phi(z)|^{2}\,\bigl(1+O(\varepsilon^{2}|z|^{2})\bigr)\,\varepsilon^{n}\,dz=\varepsilon^{2s}\int_{\mathbb{R}^{n}}|\phi|^{2}\,dz+o(\varepsilon^{2s}).$
Moreover, using the decomposition (1.6), we write
$\int_{\mathbb{H}^{n}}u_{\varepsilon}\,\mathcal{P}_{s}u_{\varepsilon}\,dV_{\mathbb{H}^{n}}=\int_{\mathbb{H}^{n}}u_{\varepsilon}\,\widetilde{\mathcal{P}}_{s}u_{\varepsilon}\,dV_{\mathbb{H}^{n}}+\int_{\mathbb{H}^{n}}u_{\varepsilon}\,B_{s}(\mathcal{A})\,u_{\varepsilon}\,dV_{\mathbb{H}^{n}},$
where $B_{s}(\mathcal{A}):=\frac{\sin(\pi s)}{\pi}\,|\Gamma(s+\tfrac{1}{2}+i\mathcal{A})|^{2}$ is a bounded spectral multiplier, see (3.27). Hence
$\Bigl|\int_{\mathbb{H}^{n}}u_{\varepsilon}\,B_{s}(\mathcal{A})\,u_{\varepsilon}\,dV_{\mathbb{H}^{n}}\Bigr|\;\lesssim\;\|u_{\varepsilon}\|_{L^{2}(\mathbb{H}^{n})}^{2}\;=\;\varepsilon^{2s}\int_{\mathbb{R}^{n}}|\phi|^{2}\,dz+o(\varepsilon^{2s}).$
For the main part, working in the ball model and using the intertwining identity
$\Bigl(\frac{1-|x|^{2}}{2}\Bigr)^{\!s+\frac{n}{2}}(-\Delta)^{s}\!\Bigl[\Bigl(\frac{1-|x|^{2}}{2}\Bigr)^{\!s-\frac{n}{2}}u\Bigr]=\widetilde{\mathcal{P}}_{s}u\qquad\text{in }(\mathbb{H}^{n},g_{\mathbb{H}^{n}}),$
we set
$v_{\varepsilon}(x):=\Bigl(\frac{1-|x|^{2}}{2}\Bigr)^{\!s-\frac{n}{2}}u_{\varepsilon}(x).$
Since $\operatorname{supp}u_{\varepsilon}\subset B_{\mathbb{H}^{n}}(p,\varepsilon)$, in normal coordinates one has $\bigl(\tfrac{1-|x|^{2}}{2}\bigr)^{s-\frac{n}{2}}=1+O(\varepsilon^{2})$ and $dV_{\mathbb{H}^{n}}(x)=(1+O(\varepsilon^{2}|z|^{2}))\,\varepsilon^{n}\,dz$ on the support. Using the critical scaling of $u_{\varepsilon}$ and Plancherel,
$\int_{\mathbb{H}^{n}}u_{\varepsilon}\,\widetilde{\mathcal{P}}_{s}u_{\varepsilon}\,dV_{\mathbb{H}^{n}}=\int_{\mathbb{R}^{n}}v_{\varepsilon}\,(-\Delta)^{s}v_{\varepsilon}\,dx=\int_{\mathbb{R}^{n}}\phi\,(-\Delta)^{s}\phi\,dy\;+\;O(\varepsilon^{2}).$
Combining the two pieces yields
$\int_{\mathbb{H}^{n}}u_{\varepsilon}\,\mathcal{P}_{s}u_{\varepsilon}\,dV_{\mathbb{H}^{n}}=\int_{\mathbb{R}^{n}}\phi\,(-\Delta)^{s}\phi\,dy\;+\;O(\varepsilon^{2})\;+\;O(\varepsilon^{2s})=\int_{\mathbb{R}^{n}}\phi\,(-\Delta)^{s}\phi\,dy\;+\;O\!\bigl(\varepsilon^{2\min\{1,s\}}\bigr).$
Thus, for $\lambda\leq 0$, we have
$\frac{\displaystyle\int_{\mathbb{H}^{n}}(\mathcal{P}_{s}u_{\varepsilon})u_{\varepsilon}\,dV_{\mathbb{H}^{n}}-\lambda\int_{\mathbb{H}^{n}}|u_{\varepsilon}|^{2}\,dV_{\mathbb{H}^{n}}}{\Bigl(\displaystyle\int_{\mathbb{H}^{n}}|u_{\varepsilon}|^{2_{s}^{*}}\,dV_{\mathbb{H}^{n}}\Bigr)^{\!2/2_{s}^{*}}}$ $=\frac{\displaystyle\Bigl[\int_{\mathbb{R}^{n}}\phi(-\Delta)^{s}\phi\,dy\Bigr]+O(\varepsilon^{2\min\{1,s\}})\;+\;(-\lambda)\,\varepsilon^{2s}\!\int_{\mathbb{R}^{n}}|\phi|^{2}\,dy+o(\varepsilon^{2s})}{\Bigl(\displaystyle\int_{\mathbb{R}^{n}}|\phi|^{2_{s}^{*}}\,dy+O(\varepsilon^{2})\Bigr)^{\!2/2_{s}^{*}}}$
$\leq\frac{\displaystyle\int_{\mathbb{R}^{n}}\phi(-\Delta)^{s}\phi\,dy}{\Bigl(\displaystyle\int_{\mathbb{R}^{n}}|\phi|^{2_{s}^{*}}\,dy\Bigr)^{\!2/2_{s}^{*}}}\;+\;C\Bigl(\varepsilon^{2\min\{1,s\}}+|\lambda|\,\varepsilon^{2s}+\varepsilon^{2}\Bigr),$
where $C>0$ depends only on $n,s$ and $\phi$. Letting $\varepsilon\downarrow 0$ and then infimizing over $\phi\in C_{c}^{\infty}(\mathbb{R}^{n})\setminus\{0\}$ gives $H_{n,s}(\lambda)\leq S_{n,s}$. Similarly, we obtain $\widetilde{H}_{n,s}(\lambda)\leq S_{n,s}$. By Proposition 2.3, we have
$\widetilde{H}_{n,s}(\lambda)=\widetilde{H}_{n,s}(0)=S_{n,s}.$
Since $\lambda_{0,s}^{\mathrm{conf}}$ and $\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$ are the bottom of the spectrum of $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s},$ respectively, we immediately obtain the following two results.
Lemma 5.4.
Let $n\geq 2$ and $s\in\bigl(0,\frac{n}{2}\bigr)$. Then, for $\lambda\in\mathbb{R}$,
$\int_{\mathbb{H}^{n}}(\mathcal{P}_{s}u)\,u\,dV_{\mathbb{H}^{n}}-\lambda\int_{\mathbb{H}^{n}}|u|^{2}\,dV_{\mathbb{H}^{n}}\ \geq 0\qquad\forall\,u\in C_{c}^{\infty}(\mathbb{H}^{n})$
holds if and only if $\lambda\leq\lambda_{0,s}^{\mathrm{conf}}$. In particular, $H_{n,s}(\lambda)<0$ iff $\lambda>\lambda_{0,s}^{\mathrm{conf}}.$
Lemma 5.5.
Let $n\geq 2$ and $s\in\bigl(0,\frac{n}{2}\bigr)$. Then, for $\lambda\in\mathbb{R}$,
$\int_{\mathbb{H}^{n}}(\widetilde{\mathcal{P}}_{s}u)\,u\,dV_{\mathbb{H}^{n}}-\lambda\int_{\mathbb{H}^{n}}|u|^{2}\,dV_{\mathbb{H}^{n}}\ \geq 0\qquad\forall\,u\in C_{c}^{\infty}(\mathbb{H}^{n})$
holds if and only if $\lambda\leq\widetilde{\lambda}_{0,s}^{\mathrm{conf}}$. In particular, $\widetilde{H}_{n,s}(\lambda)<0$ iff $\lambda>\widetilde{\lambda}_{0,s}^{\mathrm{conf}}.$
Lemma 5.6.
Let $X$ be a nontrivial function space and fix a parameter $\lambda$. Let
$f_{\lambda},\,g_{\lambda}:X\setminus\{0\}\to\mathbb{R}$
be two functionals such that
$f_{\lambda}(u)<g_{\lambda}(u)\qquad\text{for all }u\in X\setminus\{0\}.$ (5.1)
Define
$f(\lambda):=\inf_{u\in X\setminus\{0\}}f_{\lambda}(u),\qquad g(\lambda):=\inf_{u\in X\setminus\{0\}}g_{\lambda}(u).$
Then the following properties hold:
(i) If $f(\lambda)=g(\lambda)$, then $g(\lambda)$ is not attained.
(ii) Conversely, if $g(\lambda)$ is attained at some $u_{0}\in X\setminus\{0\}$, then $f(\lambda)<g(\lambda).$
Proof: Since $f_{\lambda}(u)<g_{\lambda}(u)$ for all $u\in X\setminus\{0\}$, we clearly have
$f(\lambda)=\inf_{u\in X\setminus\{0\}}f_{\lambda}(u)\;\leq\;\inf_{u\in X\setminus\{0\}}g_{\lambda}(u)=g(\lambda).$
(i) Assume by contradiction that $f(\lambda)=g(\lambda)$ and that $g(\lambda)$ is attained, i.e., there exists $u_{0}\in X\setminus\{0\}$ such that $g_{\lambda}(u_{0})=g(\lambda).$ Then, by the strict pointwise inequality (5.1),
$f_{\lambda}(u_{0})<g_{\lambda}(u_{0})=g(\lambda)=f(\lambda),$
which contradicts the definition of $f(\lambda)$ as the infimum of $f_{\lambda}$. Hence $g(\lambda)$ cannot be attained in $X\setminus\{0\}$.
(ii) Conversely, assume that $g(\lambda)$ is attained at some $u_{0}\in X\setminus\{0\}$, so that $g_{\lambda}(u_{0})=g(\lambda).$ By (5.1) we have
$f_{\lambda}(u_{0})<g_{\lambda}(u_{0})=g(\lambda).$
Using the definition of $f(\lambda)$,
$f(\lambda)=\inf_{u\in X\setminus\{0\}}f_{\lambda}(u)\;\leq\;f_{\lambda}(u_{0})<g(\lambda).$
Therefore $f(\lambda)<g(\lambda)$, which proves the second claim. ∎
5.2 Proof of Propositions 1.1 and 1.2
Proof of Proposition 1.1.
Part 1: $s=1$.
(i) By Proposition 5.1 we have, for all $\lambda\leq 0$, $S_{n,1,\Omega}(\lambda)=S_{n,1}.$ Next, we show that $S_{n,1,\Omega}(\lambda_{1,1}\left(\Omega\right))=0.$ Let $\phi_{1}$ be the first Dirichlet eigenfunction, $\Delta\phi_{1}=\lambda_{1,1}(\Omega)\phi_{1}$ in $\Omega$, $\phi_{1}\in H_{0}^{1}(\Omega)\setminus\{0\}$. Then
$0\leq S_{n,1,\Omega}(\lambda)\;\leq\;\frac{\displaystyle\int_{\Omega}|\nabla\phi_{1}|^{2}\,dx-\lambda\int_{\Omega}\phi_{1}^{2}\,dx}{\|\phi_{1}\|_{L^{2^{*}}(\Omega)}^{2}}=(\lambda_{1}-\lambda)\,\frac{\displaystyle\int_{\Omega}\phi_{1}^{2}\,dx}{\|\phi_{1}\|_{L^{2^{*}}(\Omega)}^{2}},$
thus, $S_{n,1,\Omega}(\lambda_{1,1}\left(\Omega\right))=0.$ Fix $v\in C_{c}^{\infty}(\Omega)\setminus\{0\}$, by Hölder’s inequality, we get
$\int_{\Omega}|v|^{2}\,dx\leq|\Omega|^{2/n}\Bigl(\int_{\Omega}|v|^{2^{*}}\,dx\Bigr)^{\frac{2}{2^{*}}}.$
Therefore,
$\frac{\displaystyle\int_{\Omega}|\nabla v|^{2}\,dx-\lambda\int_{\Omega}v^{2}\,dx}{\Bigl(\displaystyle\int_{\Omega}|v|^{2^{*}}\,dx\Bigr)^{2/2^{*}}}\geq-\mu\,|\Omega|^{2/n}\quad\text{for}\quad\mu>\lambda_{1,1}(\Omega).$
Taking the infimum over all $v\in C_{c}^{\infty}(\Omega)\setminus\{0\}$ gives $S_{n,1,\Omega}(\lambda)\geq-\mu|\Omega|^{2/n}>$-$\infty$.
By [4, Lemma 1.1 and Lemma 1.2] together with Lemma 5.6 (ii), then for any
$0<\mu_{1}<\mu_{2}<\lambda_{1,1}(\Omega)<\mu_{3}$
one has the strict chain
$S_{n,1,\Omega}(\mu_{3})\;<\;0\;=\;S_{n,1,\Omega}(\lambda_{1,1}\left(\Omega\right))\;<\;S_{n,1,\Omega}(\mu_{2})\;<\;S_{n,1,\Omega}(\mu_{1})\;<\;S_{n,1}.$
Combining with $S_{n,1,\Omega}(\lambda)=S_{n,1}$ for $\lambda\leq 0$ yields $\mathcal{G}_{n,1,\Omega}[S_{n,1,\Omega}]=(0,\infty).$
It remains to characterize attainability. If $\lambda>0$, by [4, Lemma 1.1 and Lemma 1.2], the infimum is attained. For $\lambda=0$, it is well known that $S_{n,1,\Omega}(0)=S_{n,1}$ is not attained on bounded domains.
If $\lambda<0$, suppose by contradiction that $S_{n,1,\Omega}(\lambda)$ is attained. Then Lemma 5.6 (ii) implies the strict monotonicity $S_{n,1,\Omega}(\lambda)>S_{n,1,\Omega}(\lambda/2),$ contradicting Proposition 5.1, which gives $S_{n,1,\Omega}(\lambda)=S_{n,1,\Omega}(\lambda/2)=S_{n,1}$ for all $\lambda\leq 0$. Therefore $S_{n,1,\Omega}(\lambda)$ is attained if and only if $\lambda\in\mathcal{G}_{n,1,\Omega}[S_{n,1,\Omega}]=(0,\infty)$.
(iii) By [4, Lemmas 1.2–1.3], Lemma 5.6(ii), and Proposition 5.1, for any choice of parameters
$\lambda\leq 0<\frac{\lambda_{1,1}(B_{1})}{4}<\mu_{1}<\mu_{2}<\lambda_{1,1}(B_{1})<\mu_{3},$
we have the chain of strict inequalities
$S_{3,1,B_{1}}(\mu_{3})<0=S_{3,1,B_{1}}\bigl(\lambda_{1,1}(B_{1})\bigr)<S_{3,1,B_{1}}(\mu_{2})<S_{3,1,B_{1}}(\mu_{1})\leq S_{3,1,B_{1}}\!\Bigl(\frac{\lambda_{1,1}(B_{1})}{4}\Bigr)\leq S_{3,1}=S_{3,1,B_{1}}(\lambda).$
We claim that
$S_{3,1,B_{1}}\!\Bigl(\frac{\lambda_{1,1}(B_{1})}{4}\Bigr)=S_{3,1}.$
Indeed, if this were false, then [4, Lemma 1.2] would imply that $S_{3,1,B_{1}}\!\bigl(\frac{\lambda_{1,1}(B_{1})}{4}\bigr)$ is attained, contradicting [4, Lemma 1.4]. Consequently, [4, Lemma 1.3] yields
$S_{3,1,B_{1}}(\mu_{1})\;<\;S_{3,1,B_{1}}\!\Bigl(\frac{\lambda_{1,1}(B_{1})}{4}\Bigr).$
By an analogous argument in (i), we conclude that $S_{3,1,B_{1}}(\lambda)$ is attained if and only if $\lambda\in\mathcal{G}_{3,1,B_{1}}[S_{3,1,B_{1}}]$.
Part 2: $s\in(0,1)$:
(i) Using [3, Claim 14.1], Proposition 5.2, Lemma 5.6, the proof is completely analogous to the case $s=1$, and we therefore omit the details.
(ii) By [3, Proposition 16.4], we obtain $(\lambda_{s}^{*},\infty)\subset\mathcal{G}_{n,s}[S_{n,s}]$ and thus, $S_{n,s}(\lambda)$ is attained if $\lambda\in(\lambda_{s}^{*},\infty)$. By Proposition 5.2, Lemma 5.6, for any $\lambda\leq 0<\lambda_{s}^{*}<\mu_{1}<\mu_{2}$,
$S_{n,s,\Omega}(\mu_{2})\;<\;S_{n,s,\Omega}(\mu_{1})\;\leq\;S_{n,s,\Omega}(\lambda_{s}^{*})\;\leq\;S_{n,s,\Omega}(0)\;=\;S_{n,s,\Omega}(\lambda).$
The proof ends. $\Box$
Proof of Proposition 1.2. When $k=1$, the proof is the following.
(i) By Proposition 5.3 and Proposition 2.3, we know that for any $\lambda\leq 0,$
$H_{n,1}(\lambda)=H_{n,1}(0)=S_{n,s}.$
By the proof of [25, Theorem 1.5, Theorem 1.6], we know that $H_{n,1}(\lambda)$ is achieved if and only if $\lambda\in(0,\lambda_{0,1}^{\mathrm{conf}}].$ Thus, by Lemma 5.6(ii), for any $\lambda\leq 0<\mu_{1}<\mu_{2}<\frac{1}{4},$ we have
$H_{n,1}(\frac{1}{4})<H_{n,2}(\mu_{1})<H_{n,1}(\mu_{1})<H_{n,1}(\lambda)=H_{n,1}(0)=S_{n,s}.$
By the Poincaré–Sobolev inequality (see [25]) and Lemma 5.4, we obtain
$H_{n,1}(\frac{1}{4})>0>H_{n,1}(\mu_{3}).$
Thus, by Lemma 5.6 (ii) we obtain the strict decreasing property and $\mathcal{G}_{n,1}[H_{n,1}]=(0,\infty).$ By Proposition 4.11, we obtain $H_{n,1}(\mu)=$-$\infty$ for $\mu>\lambda_{0,1}^{\mathrm{conf}}.$
(ii) For dimension $n=3$, by [1, Theorem 1.1], we have $H_{3,1}(\frac{1}{4})=S_{3,1}.$ Thus, by Proposition 5.3, we obtain for any $\mu_{1}\leq\frac{1}{4},$
$H_{3,1}(\mu_{1})=H_{3,1}(\frac{1}{4})=S_{3,1}.$
Again by the Poincaré–Sobolev inequality and Lemma 5.4, we obtain for any $\mu_{2}>\frac{1}{4},$
$H_{3,1}(\frac{1}{4})>0>H_{n,1}(\mu_{2}).$
Thus, we prove the inequality and $\mathcal{G}_{3,1}[H_{3,1}]=(\frac{1}{4},\infty).$ By [25, Theorem 1.7], we obtain $H_{3,1}(\lambda)$ is never achieved for any $\lambda\in\mathbb{R}$.
Now we deal with the case $k\geq 2$.
(i) By [20, Theorem 1.14] and [23, Theorem 1.7], we know that $H_{n,k}(\lambda)<S_{n,k}$ and $H_{n,k}(\lambda)$ is achieved whenever $\lambda\in(0,\lambda_{0,k}^{\mathrm{conf}})$ when $n\geq 4k$. By arguments completely analogous to those used in the preceding propositions in high dimensions, we obtain the desired conclusion here, and therefore omit the details.
(ii) By [20, Theorem 1.14] and [23, Theorem 1.7], there exists $\lambda_{k}^{\mathrm{conf}}\in(0,\lambda_{0,k}^{\mathrm{conf}})$ such that $H_{n,k}(\lambda)<S_{n,k}$ when $\lambda>\lambda_{k}^{\mathrm{conf}}$ and $H_{n,k}(\lambda)$ is achieved whenever $\lambda\in(\lambda_{k}^{\mathrm{conf}},\lambda_{0,k}^{\mathrm{conf}})$. The remainder of the proof follows by an entirely analogous argument, combined with Lemma 5.6 (ii) and Proposition 4.11, and we therefore omit the details.
(iii) When $n=2k+1$, by [24, Theorem 1.6] or [23, Theorem 1.2], we obtain $H_{n,k}(\lambda_{0,k}^{\mathrm{conf}})=S_{n,k}.$ The rest of the proof is completely analogous to (ii), and we omit it. $\Box$